### v1

# Sharp Asymptotic Minimaxity for One-Group Priors in Sparse Normal Means Problem

Abstract
In this paper, we consider the asymptotic properties of the Bayesian multiple testing rules when the mean parameter of the sparse normal means problem is modeled by a broad class of global-local priors, expressed as a scale mixture of normals. We are interested in studying the least possible risk, i.e., the minimax risk for two frequentist losses-one being the usual misclassification (or Hamming) loss, and the other one, measured as the sum of FDR and FNR. Under the beta-min separation condition used by Abraham et al. (2024) in the same context, at first, assuming the level of sparsity to be known, we propose a condition on the global parameter of our chosen class of priors, such that the resultant decision rule attains the minimax risk for both of the losses mentioned above. When the level of sparsity is unknown, we either use an estimate of the global parameter obtained from the data, the empirical Bayes approach, or propose an absolutely continuous prior on it, a full Bayes one. For both of the procedures, under some assumption on the unknown level of sparsity, we show that the decision rules also attain the minimax risk, again for both of the losses. Our results also provide a guideline regarding the selection of priors, in the sense that beyond a subclass(horseshoe-type priors) of our chosen class of priors, the minimax risk is not achievable with respect to any one of the two loss functions considered in this article. However, the subclass, horseshoe-type priors, is such a large subclass that it contains Horseshoe, Strawderman–Berger, standard double Pareto, inverse–gamma priors with $\alpha=0.5$, just to name a few. In this way, along with the most popular BH procedure and $\ell$-value based approach using spike and slab prior, a multiple testing rule based on one-group priors also achieves the optimal boundary. To the best of our knowledge, these are the first results in the literature of global-local priors which ensure the optimal minimax risk can be achieved exactly.
Keywords— Multiple testing, False Discovery Rate, global-local priors, Classification Loss.
1 Introduction
Multiple hypothesis testing is a prominent area of research in statistics due to its wide range of applications related to biology, image processing, astronomy, just to name a few. The primary objective in a multiple testing problem is to form a decision rule which controls the overall type I error. The most popular one in this context is the False Discovery Rate (FDR, defined below) introduced by Benjamini and Hochberg (1995). Suppose we have an observation vector $\mathbf{X}\in\mathbb{R}^{n}$, $\mathbf{X}=(X_{1},X_{2},\cdots,X_{n})$, where each observation $X_{i}$ is expressed as,
$X_{i}=\theta_{i}+\epsilon_{i},i=1,2,\cdots,n,$ (1)
where the parameters $\theta_{1},\theta_{2},\cdots,\theta_{n}$ are unknown mean parameters and $\epsilon_{i}$’s are the random errors present in the model. Generally, $\epsilon_{i}$’s are assumed to be independent $\mathcal{N}(0,\sigma^{2})$ random variables. In the literature of multiple testing, usually $\sigma^{2}$ is assumed to be known and hence without loss of generality, a typical choice of $\sigma^{2}$ is $\sigma^{2}=1$. This is the famous normal means model or Gaussian sequence model. In the normal means problem as given in (1), one is interested in testing simultaneously $H_{0i}:\theta_{i}=0$ vs $H_{1i}:\theta_{i}\neq 0$ for $i=1,\ldots,n$. A multiple testing procedure, defined in terms of a measurable function of the data is as follows: Let $\psi$ be a measurable function of the data, i.e., $\psi:x\in\mathbb{R}\mapsto(\psi_{i}(x))_{i=1,2,\cdots,n}\in\{0,1\}^{n}$ where $\psi_{i}\equiv\psi_{i}(X)=1$ implies rejecting the $i^{th}$ null hypothesis, $H_{0i}$. Given $\boldsymbol{\theta}=(\theta_{1},\ldots,\theta_{n})\in\mathbb{R}^{n}$ and any testing procedure $\psi$, the FDR of $\psi$ at $\boldsymbol{\theta}$ for the sparse normal model is defined as
$FDR(\boldsymbol{\theta},\psi)$ $=\mathbb{E}_{\boldsymbol{\theta}}\bigg{[}\frac{\sum_{i:\theta_{i}=0}\psi_{i}}{max(\sum_{i=1}^{n}\psi_{i},1)}\bigg{]}.$ (2)
Benjamini and Hochberg (1995) established that their proposed procedure based on p-values controls the FDR at level $\alpha$ under the assumption that the test statistics are independent. This work was later extended by Benjamini and Yekutieli (2001), when they proved the same even under positive dependence of the test statistics. After this, several authors including Sarkar (2002), Chen and Sarkar (2004), Sarkar (2007), Storey (2007), Sarkar (2008), Sarkar and Guo (2010) have investigated the problem of proposing procedures for controlling FDR in the context of the famous sparse normal means model. Given that a testing rule controls the FDR at some desired level, the next obvious question is whether the same rule can provide any control over the type II error, which results in good behaviour of power corresponding to the procedure. This probability of error is measured by the False Negative Rate (FNR). For a testing rule $\psi$, FNR at $\boldsymbol{\theta}$ is of the form
$FNR(\boldsymbol{\theta},\psi)$ $=\mathbb{E}_{\boldsymbol{\theta}}\bigg{[}\frac{\sum_{i:\theta_{i}\neq 0}(1-\psi_{i})}{max(\sum_{i=1}^{n}1_{\theta_{i}\neq 0},1)}\bigg{]}.$ (3)
In this context, a question in which researchers are often interested is the optimality of the testing procedure, i.e., the least possible value of the sum of type I and type II errors that is attainable by any multiple testing rule. In case of a single testing problem, this question was addressed in depth for a long time by Giné and Nickl (2021) and Ingster and Suslina (2012) from a non-parametric viewpoint and by Baraud (2002), Ingster et al. (2010) and Nickl and Van De Geer (2013) for different loss functions in high-dimensional problems. However, the same question was studied very recently in the context of multiple hypothesis testing. Fromont et al. (2016) studied the problem in terms of the familywise error rate. On the other hand, when the signals are non-random, this problem was initially investigated by Arias-Castro and Chen (2017), and subsequent works have been done by Belitser and Nurushev (2020) and Rabinovich et al. (2020). The multiple testing risk at $\boldsymbol{\theta}$ for the testing procedure $\psi$ is of the form
$R(\boldsymbol{\theta},\psi)$ $=FDR(\boldsymbol{\theta},\psi)+FNR(\boldsymbol{\theta},\psi).$ (4)
Given $FDR(\boldsymbol{\theta},\psi)$ and $FNR(\boldsymbol{\theta},\psi)$, for obtaining the optimal risk, the benchmark is to consider the minimax risk, which is defined as
$R(\boldsymbol{\Theta})$ $=\inf_{\psi}\sup_{\boldsymbol{\theta}\in\Theta}R(\boldsymbol{\theta},\psi),$ (5)
where the infimum is taken over all multiple testing rules and the parameter set $\Theta$ is some subset of $l_{0}[q_{n}]$ discussed before.
Note that the most naive testing rule $\psi_{i}=0$ for all $i$ implies the corresponding minimax risk is exactly $1$. Hence, it is of practitioners’ importance to consider a non-trivial testing procedure such that the resultant $R(\boldsymbol{\Theta})$ is strictly less than $1$. Intuition suggests that it is necessary to assume some conditions on the least possible value of signals for the separation of those from the noises present in the data. The most natural condition in this context is the beta-min condition (stated later formally), which ensures the non-null $\theta_{i}$’s should exceed a certain threshold. As already stated in Chapter 1, it was Arias-Castro and Chen (2017) who first studied the problem from a frequentist’s perspective. Under the beta-min condition, assuming that all signals having the same absolute value, say $M$, at first, they established that if $M>a\sqrt{2\log(\frac{n}{q_{n}})}$ for some $a<1$, there does not exist any multiple testing rule at least based on thresholding, which yields a non-trivial amount of minimax risk uniformly over $\boldsymbol{\Theta}(M)$. On the other hand, if $M>a\sqrt{2\log(\frac{n}{q_{n}})}$, for some $a$ being greater than $1$, the parameter $\alpha$ corresponding to the Benjamini-Hochberg (BH) procedure can be chosen appropriately, such that the minimax risk based on BH procedure converges to 00. As mentioned by Abraham et al. (2024), these two results act as a stepping stone for the confirmation of the importance of the beta-min condition in order to obtain a non-trivial amount of minimax risk. In other words, these ensure that for a multiple testing rule, based on thresholding, the detection boundary for signals should be close to the universal threshold, $\sqrt{2\log(\frac{n}{q_{n}})}$. Rabinovich et al. (2020) considered the same problem and derived non-asymptotic lower and upper bounds of $R(\boldsymbol{\Theta})$ when $a=a_{n}>1$ is known. They also established that these proposed bounds are of the same order as the BH procedure. However, they did not study the situation when their testing procedures result in some non-vanishing minimax risk, i.e, the precise asymptotic boundary on the signal strength $M$, for which $R(\boldsymbol{\Theta})$ goes from 00 to $1$ was missing in these existing works.
The rest of the work is organized as follows. In Section 2, we provide an extensive literature review of the works done in the context of finding minimax risk for the Gaussian location model and its extension. The results obtained based on one-group priors are stated in Section 3. It consists of two subsections, Subsection 3.1 is related to the risk obtained for the classification loss, whereas results related to the loss, defined as the sum of FDR and FNR, are discussed in Subsection 3.2. Section 4 contains an overall discussion of our work along with some future problems related to it. Proofs corresponding to our results are provided in Section 5.
1.1 Notations
For any two sequences $\{a_{n}\}$ and $\{b_{n}\}$ with $\{b_{n}\}\neq 0$, we write $a_{n}\asymp b_{n}$ to denote that there exist two constants $c_{1}$ and $c_{2}$ such that $0<c_{1}\leq a_{n}/b_{n}\leq c_{2}<\infty$ for sufficiently large $n$. Similarly, $a_{n}\lesssim b_{n}$ denotes that for sufficiently large $n$, there exists a constant $c>0$ such that $a_{n}\leq cb_{n}$. Similarly, $a_{n}\propto b_{n}$ implies there exists some constant $C>0$ such that $a_{n}=Cb_{n}$ for all $n$. Finally, $a_{n}=o(b_{n})$ denotes $\lim_{n\to\infty}\frac{a_{n}}{b_{n}}=0$. $\phi(\cdot)$ denotes the density of a standard normal distribution.
2 Existing results on minimax risk
[…]
$\boldsymbol{\Theta}=\boldsymbol{\Theta}(\tilde{a},q_{n})=\{\boldsymbol{\theta}\in l_{0}[q_{n}]:|\theta_{i}|\geq\tilde{a}\textrm{ for }i\in S_{\boldsymbol{\theta}},|S_{\boldsymbol{\theta}}|=q_{n}\},$ (6)
where $S_{\boldsymbol{\theta}}=\{i:\theta_{i}\neq 0\}$. Motivated by the previous works of Arias-Castro and Chen (2017) and Rabinovich et al. (2020), they chose $\tilde{a}$ defined in (6) to be close to the universal threshold, $\sqrt{2\log(\frac{n}{q_{n}})}$, and considered the set:
$\boldsymbol{\Theta}_{b}=\boldsymbol{\Theta}(\tilde{a}_{b},q_{n}),\tilde{a}_{b}=\sqrt{2\log(\frac{n}{q_{n}})}+b,$ (7)
with $\boldsymbol{\Theta}(\tilde{a},q_{n})$ is as same as (6). They established that for any $b\in\mathbb{R}$, the minimax risk over $\boldsymbol{\Theta}_{b}$ is $1-\Phi(b)+o(1)$, i.e.
$\inf_{\psi}\sup_{\boldsymbol{\theta}\in\boldsymbol{\Theta}_{b}}R(\boldsymbol{\theta},\psi)=1-\Phi(b)+o(1).$ (8)
This result continues to hold even if $b=b_{n}\to\infty$ and $b=b_{n}\to-\infty$. As a byproduct of this result, they obtained that the oracle thresholding rule, $\psi_{i}=\mathbf{1}\{|X_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}\},i=1,\ldots,n$ is asymptotically minimax. This result also helps them to identify a sharp boundary for nontrivial testing, which is of the form $\tilde{a}_{b}=\sqrt{2\log(\frac{n}{q_{n}})}+b$ with $b\to-\infty$. This boundary corresponds to a regime where there does not exist any multiple testing procedure which has a lower minimax risk than the trivial ones ($\psi_{i}=0$ for all $i$ and $\psi_{i}=1$ for all $i$). However, restricting $b$ to be finite (even if negative) yields some non-trivial value of the minimax risk.
Since the oracle thresholding rule depends on the number of non-zero means, $q_{n}$, which is unknown quite often, an obvious question is whether there exists any multiple testing rule which can attain the minimax risk of (8), even if the level of sparsity is unknown. Assuming that $q_{n}\lesssim n^{c}$, for some $10<c<$ 1, they proved that both the BH procedure of Benjamini and Hochberg (1995) with appropriately chosen $\alpha=\alpha_{n}\to 0$ as $n\to\infty$ and l-value based procedure using the two-groups spike-and-slab priors with a proper choice of slab distribution results in providing a positive answer to the question with $b$ defined in (8) either being finite or $b=b_{n}$ tending to $\infty$. Next, we describe the l-value based procedure briefly. For a detailed discussion, see Section S-8 in the supplementary material of Abraham et al. (2024) and the references therein. Consider the normal means model given in (1), where each $\theta_{i}$ is modeled by a two-groups spike-and-slab prior given by,
$\theta_{i}\stackrel{{\scriptstyle iid}}{{\sim}}{}(1-p)\delta_{0}+pF,i=1,2,\cdots,n,$ (9)
where $\delta_{0}$ denotes Dirac measure which is degenerate at 00 and $F$ denotes a non-degenerate absolutely continuous distribution over $\mathbb{R}$. The distribution $F$ in (9) is chosen in such a way that it can be used to model the intermediate and large values of $\theta_{i}$. In a sparse situation, $p$ is assumed to be small and denotes the probability that the true value of $\theta_{i}$ is non-zero. This modelling is widely studied in the literature, see for e.g., Mitchell and Beauchamp (1988), Johnstone and Silverman (2004), Efron (2004). As a consequence, the marginal distribution of $X_{i}$ is of the form,
[…]
In this problem, Abraham et al. (2024) assumed the slab density to be quasi-Cauchy, such that the convolution of $\phi*f$ ($f$ being the density of $F$ given in (9)), say $\tilde{f}$ is such that
$\tilde{f}(x)=\frac{1}{\sqrt{2\pi}}x^{-2}\bigg{(}1-\exp(-\frac{x^{2}}{2})\bigg{)},x\in\mathbb{R}.$
Now, given any $t\in(0,1)$, they considered a multiple testing rule which depends on the posterior inclusion probability of $\theta_{i}$, and is of the form,
$\psi^{\ell}(\mathbf{X})={(\mathbf{1}\{\ell_{i,p}(\mathbf{X})<t\})_{i=1,2,\cdots,n}},$
where
$\ell_{i,p}(\mathbf{X})=\Pi_{p}(\theta_{i}=0|\mathbf{X})=\frac{(1-p)\phi(X_{i})}{(1-p)\phi(X_{i})+p\tilde{f}(X_{i})}.$
Since $p$ is unknown, they used the Marginal Maximum Likelihood Estimate (MMLE) of $p$, which is defined as the maximizer of the likelihood (or equivalently log-likelihood) function of $\mathbf{X}$ given $p$ and is of the form $\hat{p}=argmax_{p\in[\frac{1}{n},1]}L(p)$, $L(p)$ being the log-likelihood function. Hence, Abraham et al. (2024) considered the decision rule of the form $\psi^{\hat{l}}(\mathbf{X})$, where $\psi^{\hat{\ell}}(\mathbf{X})=\psi^{{\ell}}(\mathbf{X})$ evaluated at $p=\hat{p}$.
Next, they generalized the problem of obtaining the minimax risk by relaxing assumptions both based on signal strength and the noise. For weakening the assumption based on noise, they assumed that the observations are generated from a continuous distribution on $\mathbb{R}$ such that the CDF of the observations is $\tilde{L}-$Lipschitz for some constant $\tilde{L}$. Some assumptions on the expected number of nulls were also required. On the other hand, in order to study more general situations based on the signal strength, they considered a more general set of signals containing different signal strengths. For any given $\tilde{\mathbf{a}}=(\tilde{a}_{1},\tilde{a}_{2},\ldots,\tilde{a}_{q_{n}})$, with $\tilde{a}_{j}>0$ for all $j=1,2,\ldots,q_{n}$, let us consider the set
$\boldsymbol{\Theta}(\tilde{\mathbf{a}},q_{n})=\{\boldsymbol{\theta}\in l_{0}[q_{n}]:\exists i_{1},i_{2},\ldots,i_{q_{n}}\textrm{ all distinct},|\theta_{i_{j}}|\geq\tilde{a}_{j}\textrm{ for }j=1,2,\ldots,q_{n}\},$ (11)
To study the risk, next they introduced $\Lambda_{n}(\tilde{\mathbf{a}})=\frac{1}{q_{n}}\sum_{j=1}^{q_{n}}F_{\tilde{a}_{j}}(a^{*}_{n})$ where $a^{*}_{n}=(\zeta\log(\frac{n}{q_{n}}))^{\frac{1}{\zeta}}$ for some $\zeta>1$. Under the assumptions mentioned above, for location models, the minimax risk is obtained as
$\inf_{\psi}\sup_{\boldsymbol{\theta}\in\boldsymbol{\Theta}(\tilde{\mathbf{a}},q_{n})}R(\boldsymbol{\theta},\psi)=\Lambda_{n}(\tilde{\mathbf{a}})+o(1),$ (12)
and for scale models, the minimax risk is of the form
$\inf_{\psi}\sup_{\boldsymbol{\theta}\in\boldsymbol{\Theta}(\tilde{\mathbf{a}},q_{n})}R(\boldsymbol{\theta},\psi)=2\Lambda_{n}(\tilde{\mathbf{a}})-1+o(1),$ (13)
Note that for the normal means model with varying signal strength assumption as given in (11), we have $\Lambda_{n}(\tilde{\mathbf{a}})=\frac{1}{q_{n}}\sum_{j=1}^{q_{n}}[1-\Phi(\tilde{a}_{j}-a^{*}_{n})]$, where $a^{*}_{n}=\sqrt{2\log(\frac{n}{q_{n}})}$. Similar to the signal strength as given in (7), in this case, too, they were able to establish that both the BH procedure (with a proper choice of $\alpha$ converging to zero) and the $\ell$-value based procedure achieve the risk bound, i.e.
$\sup_{\boldsymbol{\theta}\in\boldsymbol{\Theta}(\tilde{\mathbf{a}},q_{n})}R(\boldsymbol{\theta},\psi)=\Lambda_{n}(\tilde{\mathbf{a}})+o(1),$ (14)
even if the level of sparsity is unknown.
In the latter half of the paper, Abraham et al. (2024) studied the minimax risk corresponding to the classification (or Hamming) loss $L(\boldsymbol{\theta},\psi)$, which is defined as
$L(\boldsymbol{\theta},\psi)=\sum_{i=1}^{n}[\mathbf{1}\{\theta_{i}=0\}\mathbf{1}\{\psi_{i}\neq 0\}+\mathbf{1}\{\theta_{i}\neq 0\}\mathbf{1}\{\psi_{i}=0\}].$ (15)
Earlier, Butucea et al. (2018) and Butucea et al. (2023) studied the variable selection problem based on the Hamming loss. They introduced the notion of almost full recovery of a testing procedure, which is defined as follows: given a testing procedure $\psi$, almost full recovery is achieved by it with respect to the classification loss for a subset of $l_{0}[q_{n}]$, say $\boldsymbol{\Theta}_{q_{n}}$, if, as $n\to\infty$,
$\frac{1}{q_{n}}\sup_{\boldsymbol{\theta}\in\boldsymbol{\Theta}_{q_{n}}}\mathbb{E}L(\boldsymbol{\theta},\psi)=o(1).$ (16)
Abraham et al. (2024) for the Gaussian location problem obtained the minimax risk corresponding to the classification loss, which is of the form
$\frac{1}{q_{n}}\inf_{\psi}\sup_{\boldsymbol{\theta}\in\boldsymbol{\Theta}_{b}}\mathbb{E}L(\boldsymbol{\theta},\psi)=1-\Phi(b)+o(1),$ (17)
where $\boldsymbol{\Theta}_{b}$ is already defined in (7) and $b$ may be either finite or $b=b_{n}\to\pm\infty$. This result also holds for the BH and $l-$value-based procedures under the assumption of polynomial sparsity stated before.
To the best of our knowledge, the only work based on one-group priors related to minimax risk in the normal means model is due to Salomond (2017). Under the assumption that observations are generated from a normal distribution as given in (1), they considered modeling each $\theta_{i}$ by a continuous prior instead of a two-groups prior with the hierarchical formulation as
$\theta_{i}|\sigma^{2}_{i}\stackrel{{\scriptstyle ind}}{{\sim}}\mathcal{N}(0,\sigma^{2}_{i}),\sigma^{2}_{i}\stackrel{{\scriptstyle ind}}{{\sim}}\pi(\sigma^{2}_{i}),$ (18)
where $\pi(\cdot)$ denotes a density on the positive real numbers. Here $\sigma^{2}_{i}$ plays the dual role of capturing the overall sparsity of the model along with detecting the real signals. It is clear from (18) that all one-group global-local priors stated earlier can be included in this rich class of priors. This class of priors was previously studied by van der Pas et al. (2016), where they were interested in obtaining the optimal posterior contraction rate of the parameter vector around the truth. In this work, they were interested in testing $H_{0i}:\theta_{i}=0$ vs $H_{1i}:\theta_{i}\neq 0$ simultaneously. Motivated by the existing works of Carvalho et al. (2009), Datta and Ghosh (2013), Ghosh et al. (2016) and Ghosh and Chakrabarti (2017), they proposed a multiple testing rule $\boldsymbol{\psi}=(\psi_{1},\ldots,\psi_{n})$ , with $\psi_{i}=\mathbf{1}(\mathbb{E}(\varphi_{i}|\mathbf{X}=\mathbf{x})>\alpha)$ for some fixed threshold $\alpha$, where $\varphi_{i}=\frac{\sigma^{2}_{i}}{1+\sigma^{2}_{i}}$. They remarked that for $\alpha=\frac{1}{2}$, their proposed testing rule becomes equivalent to that of Ghosh and Chakrabarti (2017). Motivated by works of Arias-Castro and Chen (2017) and Rabinovich et al. (2020), they also considered the problem of studying the minimax bound of a risk which is defined as the sum of FDR and FNR. Based on the conditions proposed by van der Pas et al. (2016) along with some additional assumptions on the level of sparsity, they obtained an upper bound of the minimax risk. This risk converges to zero as the number of hypotheses $n$ diverges to $\infty$. However, there were some gaps in their work. First, their chosen boundary for the signal strengths exceeded the universal threshold. Hence, it remained uncertain whether the same upper bound (on the minimax risk) holds for the case when non-zero $\theta_{i}$’s exceed the universal threshold, i.e., the parameter space stated in (7). Second, they also did not study the situation when there are multiple levels of signals. Hence, there remain some questions still unanswered regarding the behavior of their proposed rule corresponding to that problem. Third, there is no mention of minimaxity due to classification loss in their work. All these imply that whether a one-group prior can attain the minimax risk for the Gaussian location model was not answered in depth. In a nutshell, the study of minimax risk based on one-group priors needs to be revisited once again. In the next Section, we try to provide positive answers to the questions raised in the last part of this Section based on one-group priors.
3 Results based on one-group priors
Inspired by the works of Abraham et al. (2024) and Salomond (2017), we were interested in studying the multiple testing problem $H_{1i}:\theta_{i}\neq 0$ simultaneously, when each $\theta_{i}$ is modeled as a scale mixture of normals as,
$\theta_{i}|\lambda_{i},\tau$ $\stackrel{{\scriptstyle ind}}{{\sim}}\mathcal{N}(0,\lambda^{2}_{i}\sigma^{2}\tau^{2})$
$\lambda^{2}_{i}$ $\stackrel{{\scriptstyle ind}}{{\sim}}\pi_{1}(\lambda^{2}_{i})$ (19)
$\tau$ $\sim\pi_{2}(\tau),$
for some choices of $\pi_{1}$ and $\pi_{2}$. Note that, in the hierarchical formulation of (3), the posterior mean of $\theta_{i}$ is obtained as
$\mathbb{E}(\theta_{i}|\mathbf{X},\tau,\sigma)=(1-\mathbb{E}(\kappa_{i}|X_{i},\tau,\sigma))X_{i},$ (20)
where $\kappa_{i}=1/(1+\lambda^{2}_{i}\tau^{2})$. As a result, in the literature of one-group priors, $\mathbb{E}(\kappa_{i}|X_{i},\tau,\sigma)$ is interpreted as a shrinkage coefficient and hence such priors are also known as one-group shrinkage priors. Out of the two sets of hyperparameters ($\lambda_{i}$ and $\tau$) present in the hierarchy (3), $\lambda_{i}$’s are named as local shrinkage parameters and $\tau$ as global shrinkage parameter. Hence, alternatively, one-group priors are also termed as global-local shrinkage priors. Different choices of $\pi_{1}(\cdot)$ proposed by several authors have enriched the literature over the years. Some of the examples of such priors are due to t-prior (Tipping (2001)), the Laplace prior (Park and Casella (2008)), the normal-exponential-gamma prior (Griffin and Brown (2005)), horseshoe prior (Carvalho et al. (2009), Carvalho et al. (2010), Polson and Scott (2010), Polson and Scott (2012)), three parameter beta normal priors (Armagan et al. (2011)), generalized double pareto priors (Armagan et al. (2013)), the Dirichlet-laplace prior (Bhattacharya et al. (2015)), horseshoe+ prior (Bhadra et al. (2017)) etc.
In order to mimic the optimal properties (either in terms of estimation or testing) of its two-groups counterparts, this class of priors should have the dual nature of squelching the noises towards zero along with keeping intact the true signals. In this hierarchical formulation, the global parameter $\tau$ plays the level of sparsity and hence a smaller value of it ensures a large mass around the origin; on the other hand, $\lambda_{i}$’s are used to detect the signals present in the model. This results in noise observations being shrunk towards zero while the real signals are left unshrunk. This is known as tail robustness and hence the priors are named as tail robust priors.
Motivated by the works of Polson and Scott (2010), Ghosh et al. (2016) considered a general class of priors of the form
$\pi_{1}(\lambda^{2}_{i})=K(\lambda^{2}_{i})^{-a-1}L(\lambda^{2}_{i}),$ (21)
where $K>0$ is the constant of proportionality and $a$ is any positive real number and L is the slowly varying function defined earlier. They established that for different values of $a$ and different functions $L(\cdot)$, the class of priors they consider contain three parameter beta normal priors introduced by Armagan et al. (2011), generalized double pareto prior due to Armagan et al. (2013), just to name a few. Three parameter beta normal is a rich class containing horseshoe, Strawderman-Berger prior, Normal-exponential-gamma priors.
Using a simulation study, Carvalho et al. (2009) at first provides an indication that $\mathbb{E}(1-\kappa_{i}|X_{i},\tau)$ with $\kappa_{i}=1/(1+\lambda^{2}_{i}\tau^{2})$ mimics the role of the posterior inclusion probability ($\Pi(\theta_{i}\neq 0|\mathbf{X})$) corresponding to the spike and slab prior. They proposed the following decision rule intuitively,
$\text{Reject}\hskip 1.42271ptH_{0i}\hskip 1.42271pt\text{if}\hskip 14.22636pt\mathbb{E}(1-\kappa_{i}|X_{i},\tau)>\frac{1}{2}\hskip 1.42271pt,i=1,2,\cdots,n.$ (22)
Later Datta and Ghosh (2013) and Ghosh and Chakrabarti (2017) established optimality of the decision rule (22) in the context of minimizing the Bayes risk under $0-1$ loss function. Motivated by these works, in order to study the minimax risk of different loss functions based on one-group priors for the normal means problem, we also consider the decision rule $\psi=(\psi_{1},\psi_{2},\cdots,\psi_{n})$, where for $i=1,2,\cdots,n$,
$\psi_{i}\equiv\psi_{i}(X_{i})=\mathbf{1}\{\mathbb{E}(1-\kappa_{i}|X_{i},\tau)>\frac{1}{2}\}.$ (23)
Based on the decision rule (23), we want to investigate whether any global-local prior of the form (3) satisfying (21) can attain the minimax risk either based on Hamming loss or based on the sum of FDR and FNR, irrespective of the level of sparsity being known or not.
For the theoretical study of the decision rule (23), we assume the slowly varying function $L(\cdot)$ defined in (21) satisfies Assumption 1. For $a\geq\frac{1}{2}$:
(a) There exists some $c_{0}(>0)$ such that $L(t)\geq c_{0}\;\forall t\geq t_{0}$, for some $t_{0}>0$, which depends on both $L$ and $c_{0}$.
(b) There exists some $M\in(0,\infty)$ such that $\sup_{t\in(0,\infty)}L(t)\leq M$.
It is evident from earlier results of Datta and Ghosh (2013), van der Pas et al. (2014), Ghosh et al. (2016), Ghosh and Chakrabarti (2017) that the global shrinkage parameter $\tau$ plays an important role in capturing the underlying sparsity of the model. However, the sparsity level is usually unknown. Since the decision rule (23) also contains $\tau$, it is of equal importance to propose a modified version of it when the level of sparsity is unknown. Using the empirical Bayes estimate of $\tau$, proposed by van der Pas et al. (2014), the decision rule modifies to, for $i=1,2,\cdots,n$,
$\psi_{i}\equiv\psi^{\text{EB}}_{i}(\mathbf{X})=\mathbf{1}\{\mathbb{E}(1-\kappa_{i}|X_{i},\widehat{\tau})>\frac{1}{2}\},$ (24)
where
$\widehat{\tau}=\max\{\frac{1}{n},\frac{1}{c_{2}n}\sum_{i=1}^{n}\mathbf{1}(|X_{i}|>\sqrt{c_{1}\log n})\},$ (25)
with $c_{1}\geq 2$ and $c_{2}\geq 1$. An alternative approach is to model $\tau$ by an absolutely continuous prior $\pi(\tau)$ on it, known as the full Bayes approach. Motivated by the previous work of Paul and Chakrabarti (2022), here, we consider a general class of priors $\pi(\cdot)$ satisfying the following condition:
[…]
In this procedure, the decision rule is modified as, for $i=1,2,\cdots,n$,
$\psi_{i}\equiv\psi^{\text{FB}}_{i}(\mathbf{X})=\mathbf{1}\{\mathbb{E}(1-\kappa_{i}|\mathbf{X})>\frac{1}{2}\},$ (26)
where the full Bayes posterior mean of $1-\kappa_{i}$ is defined as,
[…]
We will study the optimality of the decision rules (23), (24), and (26) depending on the level of sparsity to be known or not, for both classification loss and the loss measured as the sum of FDR and FNR. Results of the minimax risk related to the Hamming or classification loss using one-group priors are discussed in Subsection 3.1. On the other hand, Subsection 3.2 contains results based on minimax risk corresponding to the loss defined as the sum of FDR and FNR using one-group shrinkage priors.
We end this Section with a Proposition based on a lower bound on the probability of type II error corresponding to the $i^{th}$ hypothesis. This shows that why $a>\frac{1}{2}$ in (21) can not be used as a good choice for the broad class of one-group priors considered here.
Proposition 1.
Suppose $X_{i}\stackrel{{\scriptstyle ind}}{{\sim}}\mathcal{N}(\theta_{i},1),i=1,2,\cdots,n$. We want to test hypotheses $H_{0i}:\theta_{i}=0$ vs $H_{1i}:\theta_{i}\neq 0$ simultaneously based on the decision rule (23) using one-group priors. Assume that $\tau\to 0$ as $n\to\infty$ such that $\frac{n\tau}{q_{n}}\to C\in(0,\infty)$. Also assume that slowly varying function $L(\cdot)$ satisfies Assumption 1. Then for $a>\frac{1}{2}$ in (21), for $i=1,\ldots,q_{n}$ and $b\in\mathbb{R}$, we have,
$\sup_{|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b}t_{2i}\to 1,\textrm{ as }n\to\infty,$ (27)
where $t_{2i}$ denotes the type II error corresponding to $i^{th}$ hypothesis.
Observe that for Theorems 1 and 2, the loss function $L(\boldsymbol{\theta},\psi)$ for which the minimax risk has to be obtained, depends on $t_{1i}$ and $t_{2i}$, the type I and type II errors corresponding to the $i^{th}$ hypothesis, respectively. Later, it will be clear from the proofs of these two Theorems that out of type I and type II errors, type II error is the dominating one. Hence, the lower bound on the type II error over the subset $\boldsymbol{\Theta}_{b}$ for $a>\frac{1}{2}$ implies the overall risk also exceeds the minimax risk, $1-\Phi(b)+o(1)$, in this case. On the other hand, when one is interested in studying the minimax risk based on the loss, defined as the sum of FDR and FNR, then also, the proofs of Theorems 4 and 5 imply that those mainly depend on the FNR, which is as same as the type II error. Hence, the conclusion from Proposition 1 goes through here, too. As a consequence, Proposition 1 acts as a guideline for us regarding the choice of $a$ given in (21) in the subsequent subsections in order to obtain the minimax risk.
3.1 Results related to the classification loss
Motivated by the result of Proposition 1, from this Subsection, We are interested in the special case when $a=\frac{1}{2}$ in (21). This subclass of priors is named as horseshoe-type priors. Note that it is a rich subclass which contains the three parameter beta normal mixtures with $\alpha=0.5$ and $\beta>0$ (e.g. horseshoe, Strawderman-Berger), the generalized double Pareto priors with $\alpha=1$ (e.g. standard double Pareto), and the inverse–gamma priors with $\alpha=0.5$, etc.
As stated earlier, we are interested in studying the optimality of our decision rule (23) based on horseshoe-type priors in the sense that whether (23) can attain the minimax risk over the set $\boldsymbol{\Theta}_{b}$ or not. At first, we assume the level of sparsity is known and use the global parameter $\tau$ as a tuning one. We propose a condition on $\tau$, depending on the level of sparsity. The first result of our work is as follows:
Theorem 1.
Suppose $X_{i}\stackrel{{\scriptstyle ind}}{{\sim}}\mathcal{N}(\theta_{i},1),i=1,2,\cdots,n$. We want to test hypotheses $H_{0i}:\theta_{i}=0$ vs $H_{1i}:\theta_{i}\neq 0$ simultaneously based on the decision rule (23) using one-group priors. Assume that $\tau\to 0$ as $n\to\infty$ such that $\frac{n\tau}{q_{n}}\to C\in(0,\infty)$. Also assume that slowly varying function $L(\cdot)$ satisfies Assumption 1. Then for horseshoe-type priors (i.e. for $a=\frac{1}{2}$), we have
$\frac{1}{q_{n}}\sup_{\boldsymbol{\theta}\in\boldsymbol{\Theta}_{b}}\mathbb{E}L(\boldsymbol{\theta},\psi)=1-\Phi(b)+o(1),$ (28)
where $L(\boldsymbol{\theta},\psi)$ is same as defined in (15) and $\boldsymbol{\Theta}_{b}$ is defined in (7). This result holds for any $b\in\mathbb{R}$ or $b=b_{n}\to\infty$.
Remark 1.
The importance of Theorem 1 lies in proposing a decision rule based on one-group priors which attains the minimax risk corresponding to the classification or Hamming loss. To the best of our knowledge, it is the first result in the literature of global-local priors in this context. In order to establish this result, the only condition one needs on $\tau$ is that $\tau$ is of the same order as $\frac{q_{n}}{n}$, i.e., the proportion of non-zero means, asymptotically. The same condition was previously used by Ghosh et al. (2016) while they obtained the optimality of the same testing rule with respect to the Bayes risk for the normal means model assuming the observations are generated from a two-groups spike and slab distribution. Note that, the result is established via providing bounds on the type I and type II errors. For type I error, Theorem 6 of Ghosh et al. (2016) comes in handy. However, the calculation for the type II error involves the use of some technical observations and many non-trivial arguments. In a nutshell, this result illustrates that, like the BH procedure of Benjamini and Hochberg (1995) and the $\ell$-value approach based on two-groups spike-and-slab priors, a decision rule (23) based on a broad class of global-local prior can also attain the minimax risk for the normal means model under the assumption that the sparsity level is known.
The main drawback of Theorem 1 is that it assumes the sparsity of the underlying model to be known, which holds seldom. This motivates us to consider the adaptive decision rule (24) and study the same in terms of the minimax risk. Now consider the following Theorem.
Theorem 2.
Consider the setup of Theorem 1. Suppose we want to test hypotheses $H_{0i}:\theta_{i}=0$ vs $H_{1i}:\theta_{i}\neq 0$ simultaneously based on the decision rule (24) using one-group priors, where $\widehat{\tau}$ is the same as (25). Assume that the number of non-zero means $q_{n}$ satisfies $n^{\beta_{1}}\leq q_{n}\leq n^{\beta_{2}}$ for some $10<\beta_{1}<\beta_{2}<$ 1. Also assume that slowly varying function $L(\cdot)$ satisfies Assumption 1. Then for horseshoe-type priors (i.e. for $a=\frac{1}{2}$), we have
$\frac{1}{q_{n}}\sup_{\boldsymbol{\theta}\in\boldsymbol{\Theta}_{b}}\mathbb{E}L(\boldsymbol{\theta},\psi)=1-\Phi(b)+o(1),$ (29)
where $L(\boldsymbol{\theta},\psi)$ and $\Theta_{b}$ are same as defined in Theorem 1. This result holds for any $b\in\mathbb{R}$ or $b=b_{n}\to\infty$.
Remark 2.
The key importance of Theorem 2 lies in providing a positive answer to the question that the minimax risk based on classification loss can still be attained using the decision rule, even if the level of sparsity is unknown. Though the two results (Theorem 1 and 2) attain the same risk, but the calculations are quite different. Note that, due to the introduction of $\widehat{\tau}$ in place of $\tau$ as stated in (24), the decision rule corresponding to rejecting $i^{th}$ null hypothesis depends on the entire dataset $\mathbf{X}$, not on $X_{i}$ only, and also has a complicated form. Hence, in order to bypass the calculations based on $\mathbb{E}(1-\kappa_{i}|X_{i},\widehat{\tau})$ directly, we try to incorporate some salient features of $\mathbb{E}(1-\kappa_{i}|x_{i},{\tau})$ for any fixed $x_{i}\in\mathbb{R}$. Observe that, for any fixed $x_{i}\in\mathbb{R}$, $\mathbb{E}(1-\kappa_{i}|x_{i},{\tau})$ is non-decreasing in $\tau$. Hence, for calculations corresponding to the type I error, first, we truncate $\widehat{\tau}$ in a range such that the monotonicity of $\tau$ can be used. In that range, after using monotonicity, one can employ the same arguments used in Theorem 1 for controlling type I error in that part. On the complementary part of the range, we need to provide nontrivial arguments by exploiting the structure of $\widehat{\tau}$. For controlling type II error, we apply the procedure on $\mathbb{E}(\kappa_{i}|x_{i},{\tau})$, using the fact that for any for any fixed $x_{i}\in\mathbb{R}$, $\mathbb{E}(\kappa_{i}|x_{i},{\tau})$ is non-increasing in $\tau$. Nevertheless, we should mention that, though the stated approach looks similar to those used in Theorems 10 and 11 of Ghosh et al. (2016), our procedure is much different from those, since observations generated in this case are different from that problem, as there is no assumption of an underlying two-groups structure here, to be specific. In this sense, the techniques used to derive Theorems 10 and 11 of Ghosh et al. (2016) can not be applied in this context. Finally, we should not forget that this result implies the decision rule (24) acts as an alternative to the famous BH procedure of Benjamini and Hochberg (1995) and the $\ell$-value approach based on two-groups spike-and-slab priors, even if the sparsity level is unknown.
Theorem 3.
Consider the setup of Theorem 1. Suppose we want to test hypotheses $H_{0i}:\theta_{i}=0$ vs $H_{1i}:\theta_{i}\neq 0$ simultaneously based on the decision rule (26) using one-group priors, where the prior on $\tau$ satisfies (C4) with $\alpha_{n}\propto\frac{(\log n)^{\delta_{2}}}{n}$ for some $0<\delta_{2}<\frac{1}{2}$. Assume that the number of non-zero means $q_{n}$ satisfies $q_{n}\propto(\log n)^{\delta_{3}}$ for some $\delta_{3}>0$. Also assume that slowly varying function $L(\cdot)$ satisfies Assumption 1. Then for horseshoe-type priors (i.e. for $a=\frac{1}{2}$), we have
$\frac{1}{q_{n}}\sup_{\boldsymbol{\theta}\in\boldsymbol{\Theta}_{b}}\mathbb{E}L(\boldsymbol{\theta},\psi)=1-\Phi(b)+o(1),$ (30)
where $L(\boldsymbol{\theta},\psi)$ and $\Theta_{b}$ are same as defined in Theorem 1. This result holds for any $b\in\mathbb{R}$ or $b=b_{n}\to\infty$.
3.2 Results related to minimax risk for FDR+FNR
As alluded to earlier, in this Subsection, too, we confine our interest in studying whether the decision rules based on horseshoe-type priors (23) and (24) can achieve the minimax risk, asymptotically, irrespective of the level of sparsity to be known or not. Similar to the previous subsection, here, too, at first we assume the level of sparsity to be known and try to model $\tau$ accordingly. Our result provides an affirmative conclusion that, similar to the BH and $l$-value procedures, as stated by Abraham et al. (2024), the decision rule (23) can also attain the minimax risk assuming the proportion of non-zero means to be known. The result is stated formally below.
Theorem 4.
Suppose $X_{i}\stackrel{{\scriptstyle ind}}{{\sim}}\mathcal{N}(\theta_{i},1),i=1,2,\cdots,n$. We want to test hypotheses $H_{0i}:\theta_{i}=0$ vs $H_{1i}:\theta_{i}\neq 0$ simultaneously based on the decision rule (23) using one-group priors. Assume that $\tau\to 0$ as $n\to\infty$ such that $\frac{n\tau}{q_{n}}\to C\in(0,\infty)$. Also assume that slowly varying function $L(\cdot)$ satisfies Assumption 1. Then for horseshoe-type priors (i.e. for $a=\frac{1}{2}$), we have
$\sup_{\boldsymbol{\theta}\in\boldsymbol{\Theta}_{b}}R(\boldsymbol{\theta},\psi)=1-\Phi(b)+o(1),$ (31)
where $\Theta_{b}$ is same as defined in Theorem 1 and $R(\boldsymbol{\theta},\psi)$ is defined in (4). This result continues to hold for any $b\in\mathbb{R}$ or $b=b_{n}\to\infty$.
Remark 3.
Similar to the results of the previous Subsection, this result also justifies the selection of the decision rule (23) based on one-group priors since it attains the minimax risk for the loss function defined as the sum of FDR and FNR. As stated earlier, the only work in the context of obtaining an upper bound to the minimax risk based on this loss induced by one-group priors was considered by Salomond (2017). However, their choice of the minimum value of the non-zero means exceeds the universal threshold. Not only this, they did not provide any answer regarding whether their proposed decision rule can attain the minimax risk. This result provides a positive answer to the question for the horseshoe-type priors corresponding to the minimax risk when the supremum of $\boldsymbol{\theta}$ is considered over a set that contains the universal threshold. An advantage of this result is that the upper bound can be made arbitrarily small by taking $b=b_{n}\to\infty$. The proof is done by providing bounds on FDR and FNR. The truncation used in bounding FDR is inspired by their approach. However, we have to do delicate calculations after that. We need to use Hoeffding’s inequality to bound some quantity related to FDR. The technique used in providing a bound of FNR is quite nontrivial from that of Salomond (2017). In that context, bound on type II error, as derived in previous results, comes in useful. Our proof also demonstrates that some calculations used in Salomond (2017) can be done using a simpler approach, too.
Given that (23) achieves the minimax risk, it is natural to ask whether the adaptive decision rule (24), too, can provide the same result, when the number of non-zero means is unknown. The following result answers the question in an affirmative sense.
Theorem 5.
Consider the setup of Theorem 1. Suppose we want to test hypotheses $H_{0i}:\theta_{i}=0$ vs $H_{1i}:\theta_{i}\neq 0$ simultaneously based on the decision rule (24) using one-group priors, where $\widehat{\tau}$ is the same as defined in van der Pas et al. (2014). Assume that the number of non-zero means $q_{n}$ satisfies $n^{\beta_{1}}\leq q_{n}\leq n^{\beta_{2}}$ for some $10<\beta_{1}<\beta_{2}<$ 1. Also assume that slowly varying function $L(\cdot)$ satisfies Assumption 1. Then for horseshoe-type priors (i.e. for $a=\frac{1}{2}$), we have
$\sup_{\boldsymbol{\theta}\in\boldsymbol{\Theta}_{b}}R(\boldsymbol{\theta},\psi)=1-\Phi(b)+o(1),$ (32)
where $\Theta_{b}$ is same as defined in Theorem 1 and $R(\boldsymbol{\theta},\psi)$ is defined in (4). This result holds for any $b\in\mathbb{R}$ or $b=b_{n}\to\infty$.
Remark 4.
This result elucidates that, like the BH procedure of Benjamini and Hochberg (1995) and the $\ell$-value approach based on two-groups spike-and-slab priors, a decision rule (24) based on a broad class of global-local prior can also attain the minimax risk for the normal means model, even if the sparsity level is unknown. This type of result was earlier considered by Salomond (2017) for a broader class of priors, which contains our chosen class of priors. They proved that the upper bound of the minimax risk for their proposed decision rule converges to 00 as $n\to\infty$. However, their work did not provide any answer to the question of whether their decision rule can attain the minimax risk or not. On the contrary, our result does provide a positive answer to that question. As a byproduct of our result, the minimax risk can be made arbitrarily small by choosing $b=b_{n}\to\infty$, even for an unknown sparse situation. In this way, their result becomes a special case of ours under our chosen class of priors under study.
Theorem 6.
Consider the setup of Theorem 1. Suppose we want to test hypotheses $H_{0i}:\theta_{i}=0$ vs $H_{1i}:\theta_{i}\neq 0$ simultaneously based on the decision rule (26) using one-group priors, where the prior on $\tau$ satisfies (C4) with $\alpha_{n}\propto\frac{(\log n)^{\delta_{2}}}{n}$ for some $0<\delta_{2}<\frac{1}{2}$. Assume that the number of non-zero means $q_{n}$ satisfies $q_{n}\propto(\log n)^{\delta_{3}}$ for some $\delta_{3}>0$. Also assume that slowly varying function $L(\cdot)$ satisfies Assumption 1. Then for horseshoe-type priors (i.e. for $a=\frac{1}{2}$), we have
$\sup_{\boldsymbol{\theta}\in\boldsymbol{\Theta}_{b}}R(\boldsymbol{\theta},\psi)=1-\Phi(b)+o(1),$ (33)
where $\Theta_{b}$ is same as defined in Theorem 1 and $R(\boldsymbol{\theta},\psi)$ is defined in (4). This result holds for any $b\in\mathbb{R}$ or $b=b_{n}\to\infty$.
4 Concluding Remarks
Due to the availability of large datasets related to several fields like Astronomy, Bioinformatics, Image Processing, and many more, multiple hypothesis testing is a serious problem of concern for the researchers. Hence, the main task is to control the overall type I error of the problem, the most famous one being the FDR. Given that a procedure controls FDR, it is obvious to ask about its performance in controlling FNR, i.e., the optimal value of the risk corresponding to the loss defined as the sum of FDR and FNR. Another related problem of concern is the optimal risk of a testing procedure based on classification loss. For the normal means model, Abraham et al. (2024) obtained the minimax risk for both these problems and established that the BH procedure of Benjamini and Hochberg (1995) and $\ell$-value procedure using spike and slab prior attain the risk, even if the sparsity pattern is unknown. This raises the question of whether any continuous one-group prior can attain the minimax risk or not. In this work, we provide a positive answer to this question by considering a broad class of one-group priors under study. To the best of our knowledge, these are the first results of its kind. As a consequence of our results, this gives freedom to practitioners to choose one-group priors as an alternative to its two-groups counterpart with more confidence.
Though our results are promising and ensure a general class of continuous priors can be considered as an alternative to the ’gold standard’ spike-and-slab priors, there are some questions that still need to be addressed from a decision theoretic viewpoint. First, when the level of sparsity is unknown, the most general estimate of $\tau$ as stated by van der Pas et al. (2017) is the MMLE of $\tau$, defined as the maximizer of the marginal likelihood function of $\mathbf{X}$ given $\tau$. So, it is natural to ask whether the decision rule (24) using the MMLE of $\tau$ can attain the minimax risk for the two loss functions considered here. Hence, one might be interested to study the optimality of this procedure in this context.
Another possible extension of our work is studying the class of priors proposed by van der Pas et al. (2016). For that class of priors, first, it is needed to prove that, assuming the level of sparsity to be known, whether any multiple testing rule can attain the minimax risk, asymptotically, under the beta-min condition of Abraham et al. (2024). The question regarding the optimality of a testing procedure for an unknown level of sparsity is open till now. All of these can be considered as some problems of interest to be discussed in detail in the future.
5 Proofs
Proof of Proposition 1.
From the definition of type II error, we have
$t_{2i}$ $=\mathbb{P}_{\theta_{i}\neq 0}(\mathbb{E}(\kappa_{i}|X_{i},\tau)>\frac{1}{2})$
$=\mathbb{P}_{\theta_{i}\neq 0}(\mathbb{E}(1-\kappa_{i}|X_{i},\tau)\leq\frac{1}{2})$
$\geq\mathbb{P}_{\theta_{i}\neq 0}(K_{1}\tau^{2a}\exp(\frac{X^{2}_{i}}{2})(1+o(1))\leq\frac{1}{2})$
$=\mathbb{P}_{\theta_{i}\neq 0}(X^{2}_{i}\leq 4a\log(\frac{1}{\tau})(1+o(1)))$
$=\Phi(\sqrt{4a\log(\frac{1}{\tau})}(1+o(1)-\theta_{i})-\Phi(-\sqrt{4a\log(\frac{1}{\tau})}(1+o(1)-\theta_{i}).$ (34)
The inequality in the third line holds by using Theorem 4 of Ghosh et al. (2016).
Note that in order to provide a lower bound on the type II error on the subset of $\boldsymbol{\Theta}_{b}$, it is enough to consider only one candidate of $\theta_{i}$ from that set and show that the lower bound of (5) for that choice of $\theta_{i}$ tends to $1$ as $n\to\infty$. Here, take $b=0$ and $\theta_{i}=\sqrt{2\log(\frac{n}{q_{n}})}$. Next, note that under the assumption on $\tau$ that $\frac{n\tau}{q_{n}}\to C\in(0,\infty)$, we obtain
$\log(\frac{1}{\tau})=\log(\frac{n}{q_{n}})+\log(\frac{1}{C})+\log(1+o(1))=\log(\frac{n}{q_{n}})(1+o(1)).$
This implies, for all sufficiently large $n$,
$\sup_{|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b}t_{2i}$ $\geq\Phi((\sqrt{4a}-\sqrt{2})\log(\frac{n}{q_{n}}))(1+o(1))-\Phi((-\sqrt{4a}-\sqrt{2})\log(\frac{n}{q_{n}}))(1+o(1))$
$=\Phi((\sqrt{4a}-\sqrt{2})\log(\frac{n}{q_{n}}))(1+o(1))$
$\to 1,\textrm{ as }n\to\infty.$
Here the equality holds since for $a>\frac{1}{2}$, $\Phi((\sqrt{4a}-\sqrt{2})\log(\frac{n}{q_{n}}))(1+o(1))\to 1$ as $n\to\infty$ and for any $a>0$, $\Phi((-\sqrt{4a}-\sqrt{2})\log(\frac{n}{q_{n}}))(1+o(1))\to 0$ as $n\to\infty$. ∎
Proof of Theorem 1.
The lower bound of the left hand side of (28) holds due to Theorem 7 of Abraham et al. (2024). Hence, we need to obtain the upper bound of it. Using arguments similar to Theorem 7 of Abraham et al. (2024), note that
$\mathbb{E}L(\theta,\psi)=\sum_{i:\theta_{i}=0}t_{1i}+\sum_{i:\theta_{i}\neq 0}t_{2i},$ (35)
where $t_{1i}$ and $t_{2i}$ denote the type I and type II errors corresponding to $i^{th}$ hypothesis, respectively. Hence, it is enough to show that
$\frac{1}{q_{n}}\sum_{i:\theta_{i}=0}t_{1i}=o(1),$ (36)
and
$\frac{1}{q_{n}}\sum_{i:\theta_{i}\neq 0}\sup_{|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b}t_{2i}\leq 1-\Phi(b)+o(1).$ (37)
First, we prove (35). In this case using Theorem 6 of Ghosh et al. (2016), we have
$t_{1i}\equiv t_{1}\leq K_{1}\tau^{2a}\frac{L(\frac{1}{\tau^{2}})}{\sqrt{\log(\frac{1}{\tau^{2}})}}(1+o(1)),$ (38)
where $K_{1}$ is a positive constant independent of $\tau$ and $i$ and $o(1)$ is such that $\lim_{n\to\infty}o(1)=0$ and is independent of any $i$. Hence, using (37) and under the assumption that $L(\cdot)$ satisfies Assumption 1, for $a\geq\frac{1}{2}$,
$\frac{1}{q_{n}}\sum_{i:\theta_{i}=0}t_{1i}$ $\leq K_{1}M\frac{(n-q_{n})}{q_{n}}\tau\frac{(\tau^{2})^{a-\frac{1}{2}}}{\sqrt{\log(\frac{1}{\tau^{2}})}}(1+o(1))$
$\leq K_{1}M\frac{n\tau}{q_{n}}\frac{(\tau^{2})^{a-\frac{1}{2}}}{\sqrt{\log(\frac{1}{\tau^{2}})}}(1+o(1))=o(1),\textrm{ as }n\to\infty.$
Here equality in the last line follows due to the assumption on $\tau$. This completes the proof of (36). Now, we need to provide an upper bound on $t_{2i}$ for each $i=1,2,\cdots,n$.
[…]
$t_{2i}$ $=P_{\theta_{i}}(E(\kappa_{i}|X_{i},\tau)>\frac{1}{2})$
$\leq P_{\theta_{i}}(B_{n})\leq P_{\theta_{i}}(B_{n}\cap C_{n})+P_{\theta_{i}}(C^{c}_{n}).$ (39)
Observe that, for any $\theta_{i}$
$P_{\theta_{i}}(g(X_{i},\tau,\eta,\delta)>\frac{1}{2},|X_{i}|>\sqrt{\rho\log(\frac{1}{\tau})})$
$\leq P_{\theta_{i}}(g_{1}(X_{i})>\frac{1}{4},|X_{i}|>\sqrt{\rho\log(\frac{1}{\tau})})+P_{\theta_{i}}(g_{2}(X_{i},\tau,\eta,\delta)>\frac{1}{4},|X_{i}|>\sqrt{\rho\log(\frac{1}{\tau})})$
$\leq P_{\theta_{i}}(g_{1}(X_{i})>\frac{1}{4},|X_{i}|>\sqrt{2\log(\frac{1}{\tau})})+P(\tau^{(\log(\frac{1}{\tau}))^{-\delta_{1}}}>\frac{1}{4})$
$\leq P(\log(\frac{1}{\tau})(1-\tau^{\frac{1}{1+t_{0}}})<C_{1})+P(\tau^{(\log(\frac{1}{\tau}))^{-\delta_{1}}}>\frac{1}{4})$
$\to 0\textrm{ as }\tau\to 0.$ (40)
Here in the chain of inequalities, the second one holds since $\rho>2$. Next we use that both $g_{1}(x)$ and $g_{2}(x)$ are decreasing function of $|x|$. Hence this implies for any $\theta_{i}$ and for all sufficiently large $n$
$P_{\theta_{i}}(B_{n}\cap C_{n})=0.$ (41)
As a result, (5) and (41) combine for all sufficiently large $n$
$t_{2i}$ $\leq P_{\theta_{i}}(|X_{i}|\leq\sqrt{\rho\log(\frac{1}{\tau})})$
$=\Phi(T_{n}-\theta_{i})-\Phi(-T_{n}-\theta_{i}),$ (42)
where $T_{n}=\sqrt{\rho\log(\frac{1}{\tau})}$. Next, we divide our calculations depending on $\theta_{i}$ is positive or not.
Case-1: $\theta_{i}>0$ for some $i$. Then, using the lower bound on $\theta_{i}$ the upper bound of $t_{2i}$ obtained in (5) modifies to
$t_{2i}$ $\leq\Phi(T_{n}-\theta_{i})$
$\leq\Phi(\sqrt{\rho\log(\frac{1}{\tau})}-\sqrt{2\log(\frac{n}{q_{n}})}-b),$
where the choice of $\rho$ is same as discussed before. Observe that the decision rule does not depend on how $\eta\in(0,1),\delta\in(0,1)$ and $\rho>\frac{2}{\eta(1-\delta)}$ is chosen. Hence, taking infimum over all such $\rho$’s and subsequently over all possible choices of $(\eta,\delta)\in(0,1)\times(0,1)$, and finally using continuity of $\Phi(\cdot)$, we have
$t_{2i}$ $\leq\Phi(\sqrt{2(1+g(\tau))\log(\frac{1}{\tau})}-\sqrt{2\log(\frac{n}{q_{n}})}-b).$ (43)
Next, under the assumption $\frac{n\tau}{q_{n}}\to C\in(0,\infty)$, we obtain
$\log(\frac{1}{\tau})=\log(\frac{n}{q_{n}})+\log(\frac{1}{C})+\log(1+o(1))$ (44)
and
$g(\tau))\log(\frac{1}{\tau})=(\log(\frac{n}{q_{n}}))^{1-\delta_{1}}(1+o(1)).$ (45)
Combining all these observations, we have
$t_{2i}$ $\leq\Phi(\sqrt{a_{n}+d_{n}}-\sqrt{a_{n}}-b),$ (46)
where $a_{n}=2\log(\frac{n}{q_{n}})$ and $d_{n}=2[\log(\frac{1}{C})+\log(1+o(1))]+2(\log(\frac{n}{q_{n}}))^{1-\delta_{1}}(1+o(1))$. Now, note that
$\Phi(\sqrt{a_{n}+d_{n}}-\sqrt{a_{n}}-b)-\Phi(-b)$
[…]
$\leq\sqrt{a_{n}+d_{n}}-\sqrt{a_{n}}$
$=\frac{(\sqrt{a_{n}+d_{n}}-\sqrt{a_{n}})\cdot(\sqrt{a_{n}+d_{n}}+\sqrt{a_{n}})}{(\sqrt{a_{n}+d_{n}}+\sqrt{a_{n}})}$
$=\frac{d_{n}}{(\sqrt{a_{n}+d_{n}}+\sqrt{a_{n}})}$
$\leq\frac{d_{n}}{2\sqrt{a_{n}}}.$ (47)
Note that, for $d_{n}=(\log(\frac{n}{q_{n}}))^{1-\delta_{1}}(1+o(1)),\frac{1}{2}<\delta_{1}<1$, we have $d_{n}=o(\sqrt{a_{n}})$, as $n\to\infty$. As a result, we obtain
$t_{2i}$ $\leq\Phi(\sqrt{a_{n}+d_{n}}-\sqrt{a_{n}}-b)$
$=\Phi(-b)+o(1).$ (48)
Note that, this $o(1)$ is independent of $i$.
Case-2: $\theta_{i}<0$ for some $i$. Then, using the lower bound on $-\theta_{i}$ the upper bound of $t_{2i}$ obtained in (5) modifies to
$t_{2i}$ $\leq 1-\Phi(-T_{n}-\theta_{i})$
$\leq 1-\Phi(-\sqrt{\rho\log(\frac{1}{\tau})}+\sqrt{2\log(\frac{n}{q_{n}})}+b).$ (49)
Again using arguments similar to those of (43)-(46), in this case, an upper bound on the probability of type II error is obtained as
$t_{2i}$ $\leq 1-\Phi(-\sqrt{a_{n}+d_{n}}+\sqrt{a_{n}}+b),$ (50)
where $a_{n}$ and $d_{n}$ are same as before. Next, again using the same arguments as used in (5) helps to establish that
$\Phi(-\sqrt{a_{n}+d_{n}}+\sqrt{a_{n}}+b)=\Phi(b)+o(1),$ (51)
where $o(1)$ is independent of $i$. Now, combining (5)-(51),
$t_{2i}$ $\leq 1-\Phi(b)+o(1).$ (52)
Finally (5) and (52) imply, for each $i=1,2,\cdots,q_{n}$ and $|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b$ for each $\theta_{i}\neq 0$,
$t_{2i}$ $\leq 1-\Phi(b)+o(1).$ (53)
This also ensures the upper bound of the left hand side of (37) is of the order of $1-\Phi(b)+o(1)$ and completes the proof of Theorem 1. ∎
Proof of Theorem 2.
Similar to Theorem 1, here too, we need to obtain an upper bound of the left hand side of (29). Hence, it suffices to prove that
$\sum_{i:\theta_{i}=0}t^{EB}_{1i}=o(q_{n})$ (54)
and
$\frac{1}{q_{n}}\sum_{i:\theta_{i}\neq 0}\sup_{|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b}t^{EB}_{2i}\leq 1-\Phi(b)+o(1),$ (55)
where $t^{EB}_{1i}$ and $t^{EB}_{2i}$ are the type I and type II errors for the empirical Bayes approach corresponding to $i^{th}$ hypothesis, respectively. For any $\alpha_{n}>0$, we have
$t^{EB}_{1i}$ $=\mathbb{P}_{\theta_{i}=0}(\mathbb{E}(1-\kappa_{i}|X_{i},\widehat{\tau})>\frac{1}{2})$
$=\mathbb{P}_{\theta_{i}=0}(\mathbb{E}(1-\kappa_{i}|X_{i},\widehat{\tau})>\frac{1}{2},\widehat{\tau}>2\alpha_{n})+\mathbb{P}_{\theta_{i}=0}(\mathbb{E}(1-\kappa_{i}|X_{i},\widehat{\tau})>\frac{1}{2},\widehat{\tau}\leq 2\alpha_{n})$
$=I+II\textrm{ say, }$, (56)
Recall that, by definition, for any given $x\in\mathbb{R}$, $\mathbb{E}(1-\kappa|x,\tau)$ is increasing in $\tau$. Hence, using (38), we have
$II$ $\leq\mathbb{P}_{\theta_{i}=0}(\mathbb{E}(1-\kappa_{i}|X_{i},2\alpha_{n})>\frac{1}{2})$
$\leq K_{1}M\frac{(2\alpha_{n})^{2a}}{\sqrt{\log(\frac{1}{4\alpha^{2}_{n}})}}(1+o(1)).$ (57)
Next, note that
$I$ $\leq\mathbb{P}_{\theta_{i}=0}(\widehat{\tau}>2\alpha_{n})$
$\leq\mathbb{P}_{\theta_{i}=0}(\widehat{\tau}_{1}>2\alpha_{n})+\mathbb{P}_{\theta_{i}=0}(\widehat{\tau}_{2}>2\alpha_{n}),$ (58)
where $\widehat{\tau}_{1}=\frac{1}{n}$ and $\widehat{\tau}_{2}=\frac{1}{c_{2}n}\sum_{i=1}^{n}\mathbf{1}(|X_{i}|>\sqrt{c_{1}\log n})$. Now we choose $\alpha_{n}>0$ such that $\widehat{\tau}_{1}\leq 2\alpha_{n}$ for all sufficiently large $n$. Hence, (5) implies, for all sufficiently large $n$
$I$ $\leq\mathbb{P}_{\theta_{i}=0}(\widehat{\tau}_{2}>2\alpha_{n})$
$\leq\mathbb{P}_{\theta_{i}=0}(\widehat{\tau}_{3}>\alpha_{n})+\mathbb{P}_{\theta_{i}=0}(\widehat{\tau}_{4}>\alpha_{n}),$ (59)
where $\widehat{\tau}_{3}=\frac{1}{c_{2}n}\sum_{i:\theta_{i}\neq 0}\mathbf{1}(|X_{i}|>\sqrt{c_{1}\log n})$ and $\widehat{\tau}_{4}=\frac{1}{c_{2}n}\sum_{i:\theta_{i}=0}\mathbf{1}(|X_{i}|>\sqrt{c_{1}\log n})$. Observe that,
$\widehat{\tau}_{3}$ $\leq\frac{1}{c_{2}}\frac{q_{n}}{n}.$
Hence we choose $\alpha_{n}$ such that $\alpha_{n}\geq\frac{1}{c_{2}}\frac{q_{n}}{n}$ for all sufficiently large $n$. This implies
$I$ $\leq\mathbb{P}_{\theta_{i}=0}(\widehat{\tau}_{4}>\alpha_{n})$
$=\mathbb{P}_{\theta_{i}=0}(\frac{1}{(n-q_{n})}\sum_{i:\theta_{i}=0}\mathbf{1}(|X_{i}|>\sqrt{c_{1}\log n})>c_{2}\frac{n\alpha_{n}}{(n-q_{n})})$
$\leq\mathbb{P}_{\theta_{i}=0}(\frac{1}{(n-q_{n})}\sum_{i:\theta_{i}=0}\mathbf{1}(|X_{i}|>\sqrt{c_{1}\log n})\geq{\alpha_{n}})$
$=\mathbb{P}_{\theta_{i}=0}(U_{n}-\mathbb{E}(U_{n})\geq\alpha_{n}(n-q_{n})-\mathbb{E}(U_{n})),$ (60)
where $U_{n}=\sum_{i:\theta_{i}=0}\mathbf{1}(|X_{i}|>\sqrt{c_{1}\log n})$. Next we consider the generalized version of Chernoff-Hoeffding bound for independent but non i.i.d. sequence of random variables, which is stated in the next lemma.
Lemma 1.
Let $Z_{1},Z_{2},\cdots,Z_{m}$ be $m$ independent $0-1$ random variables with $\mathbb{E}(Z_{i})=p_{i}$, $i=1,2,\cdots,m$. Let $Z=\sum_{i=1}^{m}Z_{i}$, $\mu=\mathbb{E}(Z)=\sum_{i=1}^{m}p_{i}$ and $p=\frac{\mu}{m}$. Then
[…]
and
$\mathbb{P}(Z\leq\mu-\lambda)\leq\exp{-\{mH_{1-p}(1-p+\frac{\lambda}{m})\}},\hskip 8.5359pt\text{for}\hskip 8.5359pt0<\lambda<\mu,$
where $H_{p}(x)=x\log(\frac{x}{p})+(1-x)\log(\frac{1-x}{1-p})$ is the relative entropy of $x$ w.r.t. $p$.
Here, note that $Z_{i}=\mathbf{1}(|X_{i}|>\sqrt{c_{1}\log n})$, $E(Z_{i})=p_{i}\lesssim n^{-\frac{c_{1}}{2}}(\log n)^{-\frac{1}{2}}$, and
$\mu$ $=E(Z)\lesssim n^{1-\frac{c_{1}}{2}}(\log n)^{-\frac{1}{2}},$
and
$p=\frac{\mu}{n-q_{n}}\lesssim n^{-\frac{c_{1}}{2}}(\log n)^{-\frac{1}{2}}(1+o(1)).$
For $\lambda=\alpha_{n}(n-q_{n})-\mu$, we have $0<\lambda<m-\mu$. Also, we have $p+\frac{\lambda}{n-q_{n}}=\alpha_{n}$. Hence,
$H_{p}(p+\frac{\lambda}{n-q_{n}})$ $=\alpha_{n}\log(\frac{\alpha_{n}}{p})+(1-\alpha_{n})\log(\frac{1-\alpha_{n}}{1-p}).$ (61)
Also, note that, since $c_{1}\geq 2$, $\frac{p}{\alpha_{n}}\to 0$ as $n\to\infty$.
Recall that, $\frac{\log(\frac{1}{1-y})}{y}\to 1$ as $y\to 0$. Hence, with $y=\frac{p-\alpha_{n}}{1-\alpha_{n}}$, the second term in the right hand side of (61) is of the form
$\log(\frac{1-\alpha_{n}}{1-p})$ $=\frac{p-\alpha_{n}}{1-\alpha_{n}}(1+o(1)),$ (62)
where $o(1)$ depends only on $n$ such that $\lim_{n\to\infty}o(1)=0$. Hence, using (61) and (62), an lower bound of $H_{p}(\alpha_{n})$ is given by
$H_{p}(\alpha_{n})$ $=\alpha_{n}\log(\frac{\alpha_{n}}{p})+(1-\alpha_{n})\cdot\frac{(p-\alpha_{n})}{(1-\alpha_{n})}(1+o(1))$
$=\alpha_{n}\log(\frac{\alpha_{n}}{p})(1+o(1))$
$\gtrsim\alpha_{n}(1+o(1)).$ (63)
With the use of Lemma 1 and (5) with the assumption $q_{n}=o(n)$, the upper bound of (5) is obtained as
$P_{\theta_{i}=0}({\widehat{\tau}_{4}}>\alpha_{n})$ $\leq e^{-{n\alpha_{n}}(1+o(1))}$
$\leq e^{-\frac{q_{n}}{c_{2}}(1+o(1))},$ (64)
where inequality in the last line holds due to. By choosing $\alpha_{n}=c_{2}\frac{q_{n}}{n}$, we immediately see that $\alpha_{n}\to 0$ and satisfies and. Thus, with the above choice of $\alpha_{n}$, and combining (5) and (5), we obtain
$I$ $\leq e^{-\frac{q_{n}}{c_{2}}(1+o(1))}.$ (65)
Hence, combining (5),(5) and (65), we have
$\frac{1}{q_{n}}\sum_{\theta_{i}=0}t^{EB}_{1i}$ $\lesssim\frac{n\alpha_{n}}{q_{n}}\frac{(\alpha_{n})^{2a-1}}{\sqrt{\log(\frac{1}{\alpha_{n}})}}(1+o(1))+\frac{n}{q_{n}}e^{-\frac{q_{n}}{c_{2}}(1+o(1))}.$ (66)
Note that the choice of $\alpha_{n}$ satisfies $\frac{n\alpha_{n}}{q_{n}}\to C_{3}$ for some $0<C_{3}<\infty$. Hence, the first term in the right hand side of (66) goes to zero as $n\to\infty$. On the other hand, since $q_{n}$ satisfies $n^{\beta_{1}}\leq q_{n}\leq n^{\beta_{2}}$ for some $10<\beta_{1}<\beta_{2}<$ 1, the second term too tends to zero as $n\to\infty$. This implies (54) holds.
Now we move towards establishing (55). In this context, note that for any $\gamma\in(0,\frac{1}{c_{2}})$ and for any $\zeta_{n}>0$,
$t^{EB}_{2i}$ $=\mathbb{P}_{\theta_{i}\neq 0}(\mathbb{E}(\kappa_{i}|X_{i},\widehat{\tau})>\frac{1}{2})$
$=\mathbb{P}_{\theta_{i}\neq 0}(\mathbb{E}(\kappa_{i}|X_{i},\widehat{\tau})>\frac{1}{2},\widehat{\tau}>\gamma\zeta_{n})+\mathbb{P}_{\theta_{i}\neq 0}(\mathbb{E}(\kappa_{i}|X_{i},\widehat{\tau})>\frac{1}{2},\widehat{\tau}\leq\gamma\zeta_{n})$
$=III+IV\textrm{ say, }$, (67)
Recall that, by definition, for any given $x\in\mathbb{R}$, $\mathbb{E}(\kappa|x,\tau)$ is non-increasing in $\tau$. Hence,
$III$ $\leq\mathbb{P}_{\theta_{i}\neq 0}(\mathbb{E}(\kappa_{i}|X_{i},\gamma\zeta_{n})>\frac{1}{2}).$ (68)
Now, assume that $\zeta_{n}\to 0$ such that $\frac{n\zeta_{n}}{q_{n}}\to C$ for some $C\in(0,\infty)$. Hence, replacing $\tau$ by $\gamma\zeta_{n}$ in (5)-(5) and (68), we have for each $\theta_{i}\neq 0,i=1,2,\cdots,n$,
$\sup_{|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b}III$ $\leq 1-\Phi(b)+o(1),$ (69)
where the $o(1)$ term is independent of $i$. Hence, our next target is to show that $IV\to 0$ as $n\to\infty$. Note that
$IV$ $\leq\mathbb{P}_{\theta_{i}\neq 0}(\widehat{\tau}\leq\gamma\zeta_{n})$
$\leq\mathbb{P}_{\theta_{i}\neq 0}(\frac{1}{c_{2}n}\sum_{j=1(\neq i)}^{n}\mathbf{1}(|X_{j}|>\sqrt{c_{1}\log n})\leq\gamma\zeta_{n})$
$\leq\mathbb{P}(|\frac{1}{n-1}\sum_{j=1(\neq i)}^{n}\mathbf{1}(|X_{j}|>\sqrt{c_{1}\log n})-\nu_{n}|\geq\nu_{n}-c_{2}\gamma\frac{n}{n-1}\zeta_{n})$
$\leq\frac{\nu_{n}(1-\nu_{n})}{(n-1)(\nu_{n}-c_{2}\gamma\frac{n}{n-1}\zeta_{n})^{2}},$ (70)
where $\nu_{n}=\mathbb{P}(|X_{i}|>\sqrt{c_{1}\log n})$. Since, $\nu_{n}(1-\nu_{n})\leq\frac{1}{4}$ for any $10<\nu_{n}<$ 1, it is enough to show that the denominator in the right hand side of (5) goes to $\infty$ as $n\to\infty$. Here in the chain of inequalities, the first one holds since by definition $\widehat{\tau}\geq\frac{1}{c_{2}n}\sum_{j=1(\neq i)}^{n}\mathbf{1}(|X_{j}|>\sqrt{c_{1}\log n})$. Next we use the fact that the distribution of the remaining $X_{j}$’s does not depend on the distribution of $X_{i}$. The final one follows from the use of Markov’s inequality. Observe that
$\nu_{n}$ $=1-\Phi(\sqrt{c_{1}\log n}-\theta_{i})+1-\Phi(\sqrt{c_{1}\log n}+\theta_{i}).$ (71)
Next, we divide our calculations depending on $\theta_{i}$ is positive or not.
Case-1: $\theta_{i}>0$ for some $i$. Hence, using the assumption on the growth of $q_{n}$,
$\nu_{n}$ $\geq 1-\Phi(\sqrt{c_{1}\log n}-\theta_{i})$
$\geq 1-\Phi(K_{2}\sqrt{\log n}-b),$ (72)
where $K_{2}=\sqrt{c_{1}}-\sqrt{2(1-\beta_{2})}>0$. Now we again divide our calculations depending on $b$ being positive or negative.
Subcase-1: $b\geq 0$. Then the lower bound of $\nu_{n}$ modifies to
$\nu_{n}$ $\geq 1-\Phi(K_{2}\sqrt{\log n})$
$=\frac{1}{K_{2}\sqrt{\log n}}\exp(-\frac{\log n}{2}\cdot K^{2}_{2})(1+o(1))$
$=\frac{1}{K_{2}\sqrt{\log n}}n^{-\frac{K^{2}_{2}}{2}}(1+o(1)).$ (73)
Hence,
$(n-1)(\nu_{n}-c_{2}\gamma\frac{n}{n-1}\zeta_{n})^{2}$
$\geq(n-1)(\frac{1}{K_{2}\sqrt{\log n}}n^{-\frac{K^{2}_{2}}{2}}(1+o(1))-Cc_{2}\gamma n^{-(1-\beta_{2})})^{2}.$ (74)
This implies (5) goes to $\infty$ if
$\frac{K^{2}_{2}}{2}<(1-\beta_{2})\textrm{ and }K^{2}_{2}<1.$ (75)
Simplification yields,
$c_{1}<8(1-\beta_{2})\textrm{ and }\sqrt{\frac{c_{1}}{2}}-\frac{1}{\sqrt{2}}<1-\beta_{2}<\sqrt{\frac{c_{1}}{2}}+\frac{1}{\sqrt{2}}.$
Combining these two, we obtain that the denominator of (5) goes to $\infty$ as $n\to\infty$ for any $10<\beta_{2}<$ 1. This implies, when $\theta_{i}>0$ for some $i$ and $\theta_{i}\geq\sqrt{2\log(\frac{n}{q_{n}})}+b$ with $b\geq 0$,
$IV=o(1),\textrm{ as }n\to\infty.$ (76)
Subcase-1: $b<0$. In this case, we have
$-b<K_{2}\sqrt{\log n}.$
Hence lower bound of $\nu_{n}$ is of the form
$\nu_{n}$ $\geq 1-\Phi(2K_{2}\sqrt{\log n})$
$=\frac{\exp(-2K^{2}_{2}\log n)}{2K_{2}\sqrt{\log n}}(1+o(1)).$
As a result $(n-1)(\nu_{n}-c_{2}\gamma\frac{n}{n-1}\zeta_{n})^{2}\to\infty$ as $n\to\infty$ if $2K^{2}_{2}<(1-\beta_{2})$ and $14K^{2}_{2}<$ 1. This also holds for any $10<\beta_{2}<$ 1. This implies, when $\theta_{i}>0$ for some $i$ and $\theta_{i}\geq\sqrt{2\log(\frac{n}{q_{n}})}+b$ with $b<0$,
$IV=o(1),\textrm{ as }n\to\infty.$ (77)
Combining (76) and (77) imply that when $\theta_{i}>0$ for some $i$ and $\theta_{i}\geq\sqrt{2\log(\frac{n}{q_{n}})}+b$,
$IV=o(1),\textrm{ as }n\to\infty.$ (78)
Hence, when $\theta_{i}>0$ for some $i$ and $\theta_{i}\geq\sqrt{2\log(\frac{n}{q_{n}})}+b$, combining (5), (68) and (78) ensures that
$t_{2i}$ $\leq 1-\Phi(b)+o(1),$ (79)
where $o(1)$ term is independent of $i$.
Case-2: $\theta_{i}<0$ for some $i$. Hence using the assumption on the growth of $q_{n}$,
$\nu_{n}$ $\geq 1-\Phi(\sqrt{c_{1}\log n}+\theta_{i})$
$\geq 1-\Phi(K_{2}\sqrt{\log n}-b),$ (80)
where $K_{2}$ is same as before. Note that the lower bound of $\nu_{n}$ obtained in (5) is exactly same as that of (5). Hence, applying the same calculations done in (5)-(77), we again obtain that, when $\theta_{i}<0$ for some $i$ and $-\theta_{i}\geq\sqrt{2\log(\frac{n}{q_{n}})}+b$,
$t_{2i}$ $\leq 1-\Phi(b)+o(1),$ (81)
where $o(1)$ term is independent of $i$.
Finally, (79) and (81) imply that
$\sup_{|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b}t^{EB}_{2i}\leq 1-\Phi(b)+o(1),$ (82)
with $o(1)$ being independent of any $i$. This implies (55) and completes the proof of Theorem 2. ∎
Proof of Theorem 3.
Similar to Theorem 2, here too, we need to obtain an upper bound of the left-hand side of (30). Hence, it suffices to prove that
$\sum_{i:\theta_{i}=0}t^{FB}_{1i}=o(q_{n})$ (83)
and
$\frac{1}{q_{n}}\sum_{i:\theta_{i}\neq 0}\sup_{|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b}t^{FB}_{2i}\leq 1-\Phi(b)+o(1),$ (84)
where $t^{FB}_{1i}$ and $t^{FB}_{2i}$ are the type I and type II errors for the full Bayes approach corresponding to $i^{th}$ hypothesis, respectively. Note that,
[…]
$=\int_{\frac{1}{n}}^{\alpha_{n}}E(1-\kappa_{i}|{X}_{i},\tau)\pi(\tau|\mathbf{X})d\tau\leq E(1-\kappa_{i}|{X}_{i},\alpha_{n})\cdot$ (85)
For proving the inequality above, we first use the fact that given any $x_{i}\in\mathbb{R}$, $E(1-\kappa_{i}|x_{i},\tau)$ is non-decreasing in $\tau$, and next we use (C4). Now, using Theorem 4 of Ghosh et al. (2016), we have for $a=\frac{1}{2}$, as $n\to\infty$
$E(1-\kappa_{i}|{X}_{i},\alpha_{n})\leq{K_{1}}e^{\frac{X^{2}_{i}}{2}}{\alpha_{n}}(1+o(1)).$ (86)
Here $o(1)$ depends only on $n$ such that $\lim_{n\to\infty}o(1)=0$ and is independent of $i$ and $K_{1}$ is a constant depending on $M$. Hence, we have the following :
$t^{\text{FB}}_{1i}$ $=P_{H_{0i}}(E(1-\kappa_{i}|\mathbf{X})>\frac{1}{2})$
$\leq P_{H_{0i}}\bigg{(}\frac{X^{2}_{i}}{2}>\log(\frac{1}{\alpha_{n}})-\log K_{1}-\log(1+o(1))\bigg{)}$
$=P_{H_{0i}}\bigg{(}|X_{i}|>\sqrt{2\log(\frac{1}{\alpha_{n}})}\bigg{)}(1+o(1)).$
Note that, as $\alpha_{n}<1$ for all $n\geq 1$, $02\log(\frac{1}{\alpha_{n}})>$ 0 for all $n\geq 1$. Next, using the fact that under $H_{0i},X_{i}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,1)$ with $1-\Phi(t)<\frac{\phi(t)}{t}$ for $t>0$, we get
$t^{\text{FB}}_{1i}$ $\leq 2\frac{\phi(\sqrt{2\log(\frac{1}{\alpha_{n}})})}{\sqrt{2\log(\frac{1}{\alpha_{n}})}}(1+o(1))=\sqrt{\frac{1}{\pi}}\frac{{\alpha}_{n}}{\sqrt{\log(\frac{1}{\alpha^{2}_{n}})}}(1+o(1)),$ (87)
where $o(1)$ depends only on $n$ such that $\lim_{n\to\infty}o(1)=0$. This implies, for sufficiently large $n$
$\frac{1}{q_{n}}\sum_{i:\theta_{i}=0}t^{\text{FB}}_{1i}$ $\lesssim\frac{n\alpha_{n}}{q_{n}\sqrt{\log(\frac{1}{\alpha_{n}})}}(1+o(1))$
$=\frac{(\log n)^{\delta_{2}}}{(\log n)^{\delta_{3}}\sqrt{\log n-\delta_{2}\log\log n}}(1+o(1))$
$=(\log n)^{\delta_{2}-\frac{1}{2}-\delta_{3}}(1+o(1))\to 0,\textrm{ as }n\to\infty,$ (88)
since $\delta_{2}<\frac{1}{2}$ and $\delta_{3}>0$. This implies (83) holds. Next, in order to establish (84) holds, it is enough to show that, for each $i=1,\cdots,q_{n}$, $\sup_{|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b}t^{\text{FB}}_{2i}\leq 1-\Phi(b)+o(1)$, where $o(1)$ is independent of $i$.
In order to provide an upper bound on the probability of type-II error induced by the decision rule (26), $t^{\text{FB}}_{2i}=P_{H_{1i}}(E(\kappa_{i}|\mathbf{X})>\frac{1}{2})$, we first note that,
[…]
where the inequality in the last line follows due to the fact that given any $x_{i}\in\mathbb{R}$, $E(\kappa_{i}|x_{i},\tau)$ is non-increasing in $\tau$. Hence, using arguments similar to (5)-(43) with $\tau=\frac{1}{n}$, when $\theta_{i}>0$ for $i=1,\cdots,q_{n}$, we have,
$t^{\text{FB}}_{2i}$ $\leq\Phi(\sqrt{2(1+(\log n)^{-\delta_{1}})\log n}-\sqrt{2\log(\frac{n}{q_{n}})}-b).$ (90)
Hence, in order to obtain the desired upper bound on $t^{\text{FB}}_{2i}$, it is sufficient to show that $\sqrt{(1+(\log n)^{-\delta_{1}})\log n}-\sqrt{\log(\frac{n}{q_{n}})}\to 0$ as $n\to\infty$. Note that, for $q_{n}\propto(\log n)^{\delta_{3}}$,
$\sqrt{(1+(\log n)^{-\delta_{1}})\log n}-\sqrt{\log(\frac{n}{q_{n}})}$ $=\sqrt{(1+(\log n)^{-\delta_{1}})\log n}-\sqrt{\log n-\delta_{2}\log\log n}>0,$
for all $n$. Hence, the upper bound converging to zero is enough. Observe that
$\sqrt{(1+(\log n)^{-\delta_{1}})\log n}-\sqrt{\log(\frac{n}{q_{n}})}$ $=\sqrt{(1+(\log n)^{-\delta_{1}})\log n}-\sqrt{\log n-\delta_{2}\log\log n}$
$=\frac{(\log n)^{(1-\delta_{1})}+\delta_{2}\log\log n}{\sqrt{(1+(\log n)^{-\delta_{1}})\log n}+\sqrt{\log n-\delta_{2}\log\log n}}$
$\leq(\log n)^{(\frac{1}{2}-\delta_{1})}+\delta_{2}\frac{\log\log n}{\sqrt{\log n}}(1+o(1))$
$\to 0,\textrm{ as }n\to\infty,$ (91)
since $\frac{1}{2}<\delta_{1}<1$. Combining (90) and (5) implies the desired upper bound on $t^{\text{FB}}_{2i}$ is established for $\theta_{i}>0$. The proof when $\theta_{i}<0$ holds using similar arguments. This completes the proof of Theorem 3. ∎
Proof of Theorem 4.
The lower bound to the left hand side of (31) follows from Theorem 1 of Abraham et al. (2024). Hence, it is enough to establish that the upper bound to the left hand side of (31) is of the order $1-\Phi(b)+o(1)$.
Using the definition of $FDR(\theta,\psi)$, we have for any $\lambda\in(0,1)$,
$FDR(\theta,\psi)$ $=\mathbb{E}_{\theta}\bigg{[}\frac{\sum_{i:\theta_{i}=0}\psi_{i}}{max(\sum_{i=1}^{n}\psi_{i},1)}\bigg{]}$
$=\mathbb{E}_{\theta}\bigg{[}\frac{\sum_{i:\theta_{i}=0}\psi_{i}}{max(\sum_{i=1}^{n}\psi_{i},1)}\mathbf{1}_{\sum_{i=1}^{n}\psi_{i}>\lambda q_{n}}\bigg{]}+\mathbb{E}_{\theta}\bigg{[}\frac{\sum_{i:\theta_{i}=0}\psi_{i}}{max(\sum_{i=1}^{n}\psi_{i},1)}\mathbf{1}_{\sum_{i=1}^{n}\psi_{i}\leq\lambda q_{n}}\bigg{]}$
$=V+VI\textrm{, say}.$ (92)
Observe that, for the set $\mathbf{1}_{\sum_{i=1}^{n}\psi_{i}>\lambda q_{n}}$, $max(\sum_{i=1}^{n}\psi_{i},1)>\lambda q_{n}$. Hence $V$ can be bounded as
$V$ $\leq\frac{\sum_{i:\theta_{i}=0}t_{1i}}{\lambda q_{n}}=\frac{(n-q_{n})t_{1}}{\lambda q_{n}}$
$\leq\frac{M}{\lambda}\frac{n\tau}{q_{n}}\frac{(\tau^{2})^{a-\frac{1}{2}}}{\sqrt{\log(\frac{1}{\tau^{2}})}}(1+o(1))=o(1),\textrm{ as }n\to\infty.$ (93)
The equality in the second line holds since by definition the type I error for $i^{th}$ hypothesis $t_{1i}$ is the same for all $i$, and denoted as $t_{1}$. The second inequality follows due to the use of (38) and the assumption on $L(\cdot)$. Now, note that
$\{\sum_{i=1}^{n}\psi_{i}\leq\lambda q_{n}\}\subseteq\{\sum_{i:\theta_{i}\neq 0}\psi_{i}\leq\lambda q_{n}\}.$ (94)
This implies
$VI$ $\leq\mathbb{P}_{\theta}(\sum_{i:\theta_{i}\neq 0}\psi_{i}\leq\lambda q_{n}).$ (95)
Since, $\psi_{i}$’s are independent Bernoulli random variables, hence, we use Hoeffding’s inequality. Choose $0<\lambda<\Phi(b)$. As a result, for sufficiently large $n$, $t$ defined in Hoeffding’s inequality is obtained as
$t$ $=q_{n}-\sum_{i:\theta_{i}\neq 0}t_{2i}-\lambda q_{n}$
$=(1-\lambda)q_{n}-\sum_{i:\theta_{i}\neq 0}t_{2i}$
$\geq[\Phi(b)-\lambda+o(1)]q_{n}.$ (96)
Note that, the choice of $\lambda$ implies $t>0$ for sufficiently large $n$. In order to obtain (5), we use a uniform upper bound on $t_{2i},i:\theta_{i}\neq 0$ as derived in (53). Using (5), we obtain, for all $|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b$, $i=1,2,\cdots,q_{n}$
$\mathbb{P}_{\theta}(\sum_{i:\theta_{i}\neq 0}\psi_{i}\leq\lambda q_{n})$ $\leq\exp(-\frac{2t^{2}}{q_{n}})$
$\leq\exp(-2[\Phi(b)-\lambda+o(1)]^{2}q_{n}).$ (97)
Next, combining (95)-(5), for sufficiently large $n$
$\sup_{\theta\in\Theta_{b}}VI=o(1),\textrm{ as }n\to\infty.$ (98)
Now, (5), (5) and (98) combine to prove that
$\sup_{\theta\in\Theta_{b}}FDR(\theta,\psi)=o(1),\textrm{ as }n\to\infty.$ (99)
On the other hand, for $FNR(\theta,\psi)$, from definition, we obtain
$FNR(\theta,\psi)$ $=\frac{\mathbb{E}_{\theta}(\sum_{i:\theta_{i}\neq 0}(1-\psi_{i}))}{q_{n}}$
$=\frac{1}{q_{n}}\sum_{i:\theta_{i}\neq 0}t_{2i}.$ (100)
Recall that, in (53), we have already derived that, for each $|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b$, $i=1,2,\cdots,q_{n}$,
$t_{2i}$ $\leq 1-\Phi(b)+o(1),$
where the $o(1)$ term is independent of any $i$. This combining with (5) implies for all sufficiently large $n$
$\sup_{\theta\in\Theta_{b}}FNR(\theta,\psi)$ $\leq 1-\Phi(b)+o(1).$ (101)
Finally, (99) and (101) imply the upper bound to the left hand side of (31) is of the order $1-\Phi(b)+o(1)$ and completes the proof. ∎
Proof of Theorem 5.
Here we follow the basic architecture of the proof of Theorem 4. Hence, we only need to show that the upper bound to the left-hand side of (32) is of the order $1-\Phi(b)+o(1)$.
Next, using the argument of (5), we have
$V$ $\leq\frac{\sum_{i:\theta_{i}=0}t^{EB}_{1i}}{\lambda q_{n}}$
$\lesssim\frac{n\alpha_{n}}{q_{n}}\frac{(\alpha_{n})^{2a-1}}{\sqrt{\log(\frac{1}{\alpha_{n}})}}(1+o(1))+\frac{n}{q_{n}}e^{-\frac{q_{n}}{c_{2}}(1+o(1))}.$ (102)
Here inequality in the second line holds due to (66). Note that the choice of $\alpha_{n}$ satisfies $\frac{n\alpha_{n}}{q_{n}}\to C_{3}$ for some $0<C_{3}<\infty$. Hence, the first term to the right hand side of (5) goes to zero as $n\to\infty$. On the other hand, since $q_{n}$ satisfies $n^{\beta_{1}}\leq q_{n}\leq n^{\beta_{2}}$ for some $10<\beta_{1}<\beta_{2}<$ 1, the second term too tends to zero as $n\to\infty$. This implies
$V=o(1),\textrm{ as }n\to\infty.$ (103)
Again employing (94) and (95), we obtain for sufficiently large $n$
$t$ $=q_{n}-\sum_{i:\theta_{i}\neq 0}t^{EB}_{2i}-\lambda q_{n}$
$\geq[\Phi(b)-\lambda+o(1)]q_{n},$ (104)
where the upper bound on $t^{EB}_{2i}$ as obtained in (82) is used. This lower bound with Hoeffding’s inequality confirms, for sufficiently large $n$
$\sup_{\theta\in\Theta_{b}}VI=o(1),\textrm{ as }n\to\infty.$ (105)
Combining (103) and (105) ensures
$\sup_{\theta\in\Theta_{b}}FDR(\theta,\psi)=o(1),\textrm{ as }n\to\infty.$ (106)
For $FNR(\theta,\psi)$, from definition, we obtain
$FNR(\theta,\psi)$ $=\frac{\mathbb{E}_{\theta}(\sum_{i:\theta_{i}\neq 0}(1-\psi_{i}))}{q_{n}}$
$=\frac{1}{q_{n}}\sum_{i:\theta_{i}\neq 0}t^{EB}_{2i}.$ (107)
Recall that, in (82), we have already derived that, for each $|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b$, $i=1,2,\cdots,q_{n}$,
$t^{EB}_{2i}$ $\leq 1-\Phi(b)+o(1),$
where the $o(1)$ term is independent of any $i$. This combining with (5) implies for all sufficiently large $n$
$\sup_{\theta\in\Theta_{b}}FNR(\theta,\psi)$ $\leq 1-\Phi(b)+o(1).$ (108)
Finally, (106) and (108) imply the upper bound to the left hand side of (32) is of the order $1-\Phi(b)+o(1)$ and completes the proof. ∎
Proof of Theorem 6.
Here also we follow the basic architecture of the proof of Theorem 4. Hence, in this case, too, it is sufficient to show that the upper bound to the left-hand side of (33) is of the order $1-\Phi(b)+o(1)$.
Note that, using the argument of (5) and (5), we have
$V$ $\leq\frac{\sum_{i:\theta_{i}=0}t^{FB}_{1i}}{\lambda q_{n}}$
$\lesssim\frac{n\alpha_{n}}{q_{n}\sqrt{\log(\frac{1}{\alpha_{n}})}}(1+o(1))$
$=o(1),\textrm{ as }n\to\infty.$ (109)
Again employing arguments of (94) and (95), we obtain for sufficiently large $n$
$t$ $=q_{n}-\sum_{i:\theta_{i}\neq 0}t^{FB}_{2i}-\lambda q_{n}$
$\geq[\Phi(b)-\lambda+o(1)]q_{n},$ (110)
where the upper bound on $t^{FB}_{2i}$ as obtained in Theorem 3 is used. This lower bound with Hoeffding’s inequality confirms, for sufficiently large $n$
$\sup_{\theta\in\Theta_{b}}VI=o(1),\textrm{ as }n\to\infty.$ (111)
Combining (5) and (111) ensures
$\sup_{\theta\in\Theta_{b}}FDR(\theta,\psi)=o(1),\textrm{ as }n\to\infty.$ (112)
For $FNR(\theta,\psi)$, from definition, we obtain
$FNR(\theta,\psi)$ $=\frac{\mathbb{E}_{\theta}(\sum_{i:\theta_{i}\neq 0}(1-\psi_{i}))}{q_{n}}$
$=\frac{1}{q_{n}}\sum_{i:\theta_{i}\neq 0}t^{FB}_{2i}.$ (113)
Recall that, in Theorem 3, we have already derived that, for each $|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b$, $i=1,2,\cdots,q_{n}$,
$t^{FB}_{2i}$ $\leq 1-\Phi(b)+o(1),$
where the $o(1)$ term is independent of any $i$. This combining with (5) implies for all sufficiently large $n$
$\sup_{\theta\in\Theta_{b}}FNR(\theta,\psi)$ $\leq 1-\Phi(b)+o(1).$ (114)
Finally, (112) and (114) imply the upper bound to the left hand side of (33) is of the order $1-\Phi(b)+o(1)$ and completes the proof. ∎