### v3

# Sharp Asymptotic Minimaxity for One-Group Shrinkage Priors in Sparse Gaussian Sequence Model

Abstract
This paper investigates the asymptotic properties of Bayesian multiple testing rules in the sparse Gaussian sequence model when the mean vector is modeled using a broad class of global-local priors, expressed as scale mixtures of normals. We focus on the minimax risk for two important loss functions: the standard misclassification (Hamming) loss and the composite risk given by the sum of the false discovery rate (FDR) and false non-discovery rate (FNR). Under the beta-min separation condition of Abraham, Castillo, and Roquain (2024), and assuming the sparsity level is known, we establish conditions on the global shrinkage parameter ensuring that decision rules based on this class of priors achieve the minimax risk for both losses. When sparsity is unknown, we propose both empirical Bayes and fully Bayesian strategies for the global parameter, and demonstrate that, under suitable assumptions, the resulting rules also achieve minimax risk. Our results reveal that minimaxity is attainable only within the subclass of horseshoe-type priors (with tail parameter $a=0.5$), encompassing the horseshoe, Strawderman–Berger, standard double Pareto, and inverse-gamma priors with $\alpha=0.5$, among others. Beyond this subclass, minimaxity fails for either loss. In summary, alongside classical methods such as the Benjamini–Hochberg procedure (Benjamini and Hochberg, 1995) and $\ell$-value rules based on spike-and-slab priors, multiple testing procedures built on one-group shrinkage priors also achieve the sharp minimax boundary. To the best of our knowledge, these are the first results showing that global–local priors can exactly attain optimal minimax risk in high-dimensional multiple testing.
Keywords— Multiple testing, False Discovery Rate, Global-local priors, Classification Loss.
1 Introduction
Multiple hypothesis testing has emerged as a critical focal point in statistics, particularly in the context of high-dimensional data analysis. Its relevance stretches across a multitude of scientific disciplines, including genomics, bioinformatics, medicine, economics, and finance. For example, in microarray experiments, researchers conduct thousands of tests simultaneously to pinpoint differentially expressed genes—those whose expression levels correlate with specific biological traits of interest. This highlights just one canonical scenario where analyzing sparse, high-dimensional data is essential, as the primary objective is to uncover significant signals amidst a vast pool of noise. Multiple hypothesis testing stands out as a powerful and effective strategy for achieving this goal while balancing the overall type I and type II errors at scale. The groundbreaking work of Benjamini and Hochberg (1995) significantly propelled research in this area, igniting interest and investigation from both frequentist and Bayesian perspectives. The ongoing exploration of multiple hypothesis testing not only advances statistical methods but also enhances our ability to extract meaningful insights from complex datasets, underscoring its importance in today’s scientific landscape.
Over the past three decades, numerous multiple testing procedures have been proposed in the literature, primarily aimed at controlling an overall measure of type I error at a predetermined level $\alpha\in(0,1)$. Historically, frequentist multiple testing procedures sought to control the family-wise error rate (FWER), defined as the probability of making at least one false rejection. The Bonferroni correction and its refinements ensure strong FWER control, but are often overly conservative in high-dimensional settings, resulting in very low statistical power. This motivated the search for more flexible error criteria. The seminal work of Benjamini and Hochberg (1995) introduced the false discovery rate (FDR), the expected proportion of false positives among all rejections, as a pragmatic alternative to the family-wise error rate, sparking a vast literature on multiple testing from both frequentist and Bayesian perspectives. Following this landmark work, numerous other FDR-controlling procedures have been proposed in the literature. Relevant references include works by Benjamini and Liu (1999), Benjamini and Yekutieli (2001), Sarkar (2002), Storey (2002), Storey, Taylor and Siegmund (2004), Benjamini, Krieger and Yekutieli (2006), Storey (2007), Sarkar (2007), Romano and Wolf (2007), Romano, Shaikh and Wolf (2008), Sarkar (2008), Sarkar and Guo (2010), Gavrilov, Benjamini and Sarkar (2009), Blanchard and Roquain (2009), Sun and Cai (2009), and Guo, He and Sarkar (2014), among others. Collectively, these works have established FDR as the de facto standard for error control in large-scale testing, offering a better balance between overall type I and type II errors than traditional FWER methods.
Beyond FDR, recent research has examined broader decision-theoretic risk criteria, such as the expected Hamming loss (misclassification rate) and the combined risk defined as the sum of the FDR and the false non-discovery rate (FNR). These measures allow for a sharper characterization of the trade-off between false discoveries and missed detections. A natural yardstick in this context would be to consider the minimum probable value of this combined error measure. Classical works in hypothesis testing theory such as Ingster and Suslina (2012), Baraud (2002), and Nickl and Van De Geer (2013) explored minimax testing rates in nonparametric settings, while more recent contributions (Arias-Castro and Chen, 2017; Belitser and Nurushev, 2020; Rabinovich et al., 2020; Fromont et al., 2016) extended these ideas to high-dimensional multiple testing.
In minimax analyses of multiple testing, one starts with a parameter space that is as broad as possible, bearing in mind that the trivial rule that never rejects any null hypothesis has a worst-case risk of one. To do better, one must both employ procedures that lower this risk and assume signals are not vanishingly small. The standard way to formalize the latter is via a beta-min separation, which requires non-null effects to exceed a data-driven strength threshold. Working in the sparse Gaussian sequence model with equal signal magnitudes, Arias-Castro and Chen (2017) showed a sharp dichotomy: below a universal (logarithmic) threshold, no thresholding rule—including BH-type methods—achieves a nontrivial minimax guarantee; above it, the Benjamini–Hochberg procedure, with an appropriately tuned level, drives the minimax risk to zero asymptotically. This establishes the beta-min condition as essential for meaningful minimax performance.
Building on this insight, Abraham, Castillo, and Roquain (2024) fully characterized the phase transition for sparse multiple testing. Under a beta-min assumption and across a wide range of sparsity regimes, they derived explicit asymptotic formulas for the minimax risk under both a composite error criterion (FDR+FNR) and the classification (Hamming) loss, revealing a clean boundary between the impossible and possible regimes of reliable testing. Crucially, they also proved that practical procedures—BH with a vanishing rejection level and empirical-Bayes $\ell$-value rules under spike–and–slab priors—attain these sharp limits adaptively, without prior knowledge of the sparsity. Their results extend beyond equal-effect settings, covering heterogeneous signal strengths and more general error models, underscoring the robustness of the boundary phenomenon.
In contrast to the traditional frequentist approaches, Bayesian methods provided a natural and powerful framework for multiple testing by combining prior information through hierarchical Bayes formulation coupled with decision-theoretic principles. The most natural Bayesian formulation of a multiple-hypothesis testing problem is to use a two-groups model, which models the parameters of interest with a two-groups prior or a spike-and-slab prior (Mitchell and Beauchamp (1988)), in which each parameter is modeled as a mixture of a point mass at zero (the spike part) and a continuous (typically) heavy-tailed distribution (the slab part). Under this setup, a natural Bayesian decision rule is to reject null hypotheses whose posterior inclusion probabilities fall below a pre-specified threshold, leading to the optimal Bayes rule, which minimizes the posterior expected loss. See, for instance, Efron (2004), Müller et al. (2004), Scott and Berger (2006), Müller, Parmigiani, and Rice (2006), Efron (2007), and Bogdan, Ghosh and Tokdar (2008), among others.
A major theoretical advancement along this path was made by Bogdan et al. (2011), who demonstrated that under varying levels of sparsity, both frequentist and Bayesian multiple testing procedures can achieve asymptotic Bayes optimality under sparsity (ABOS), meaning that the risk of the procedure converges to that of the optimal Bayes rule. In particular, they showed that the procedures of Benjamini and Hochberg (1995) and Bonferroni attain the ABOS property under mild conditions. Theoretical properties of two-groups formulations have also been studied by Johnstone and Silverman (2004), Castillo, and van der Vaart (2012), Castillo, Schmidt-Heiber, and van der Vaart (2015), Martin and Walker (2014), Cai and Jin (2010), and Neuvial and Roquain (2012), only to name a few.
Despite these advantages, two-group models often face daunting computational challenges in high-dimensional problems and may be less suitable when most parameters are close to, but not exactly, zero, and posterior sampling or marginal likelihood optimization must be carried out at scale. This has motivated the development of continuous one-group shrinkage priors, which are computationally more tractable and offer greater flexibility. In this formulation, each parameter is modeled as a scale mixture of normals, with a global shrinkage parameter controlling overall sparsity and local parameters allowing for signal-specific adaptivity. Prominent examples include the inverse Gamma priors (Tipping, 2001), the horseshoe prior (Carvalho, Polson and Scott, 2009, 2010), the three-parameter beta normal mixtures (Armagan, Clyde and Dunson, 2011), the generalized double Pareto prior (Armagan, Dunson and Lee, 2013), the Dirichlet–Laplace prior (Bhattacharya et al., 2015), and the horseshoe++ prior (Bhadra et al., 2017). These priors share the property of tail robustness: they strongly shrink small signals toward zero while leaving large effects essentially unshrunk, enabling estimation risks that are within logarithmic factors of the minimax benchmark without prior knowledge of sparsity. See, for instance, van der Pas, Kleijn and van der Vaart (2014), van der Pas, Salomond, and Schmidt-Heiber (2016), van der Pas, Szabo and van der Vaart (2017), Ghosh and Chakrabarti (2017), Bhadra et al. (2017), and Paul and Chakrabarti (2025), among others.
In the context of multiple hypothesis testing, Carvalho, Polson and Scott (2009) proposed an intuitive decision rule based on posterior shrinkage factors, which behave analogously to posterior inclusion probabilities in spike-and-slab models. Datta and Ghosh (2013) and Ghosh et al. (2016) studied Bayes risk under 011 loss, showing that decision rules based on a broad class of one-group shrinkage priors, including the horseshoe, can asymptotically achieve near optimal Bayes risk under sparsity. Subsequent work of Ghosh and Chakrabarti (2017), and Paul and Chakrabarti (2025) refined these results and established theoretically that the aforesaid multiple testing rule based on the broad class of horseshoe-type priors exactly attains asymptotic Bayes-optimal performance in terms of type I and type II error trade-offs, paralleling the sharp detection boundaries known for spike-and-slab priors. Their heavy-tailed local scales also yield robustness to large signals, avoiding the over-shrinkage typical of classical $\ell_{1}$ penalization. Collectively, these results establish one-group shrinkage priors as theoretically optimal, adaptive, and computationally attractive alternatives to discrete mixture models in high-dimensional inference.
Salomond (2017) investigated the asymptotic behavior of Bayesian multiple-testing procedures that utilize thresholding rules using a broad class of Gaussian scale-mixture priors. Drawing inspiration from the findings of van der Pas, Salomond, and Schmidt-Heiber (2016), they modeled each mean parameter using a broad category of one-group shrinkage priors, specifically a scale mixture of normals. Under conditions akin to the beta-min conditions for non-zero means, they derived an upper bound for the minimax risk, corresponding to the sum of the false discovery rate (FDR) and the false non-discovery rate (FNR). Their results indicated that if the non-zero signals are sufficiently large—referred to as the “separation condition”—the minimax risk approaches zero asymptotically, which suggests consistent detection capability. However, the detection boundary for non-zero signals identified by Salomond (2017) is larger than the oracle threshold proposed by Abraham, Castillo, and Roquain (2024). This raises the question of whether any one-group prior can asymptotically achieve the minimax risk under the beta-min condition put forth by Abraham, Castillo, and Roquain (2024) for non-zero signals, a question that remains unresolved. Additionally, Salomond (2017) did not examine the scenario involving misclassification loss. When the level of sparsity is unknown, one potential solution is to use the full Bayesian approach, which their paper did not address.
Motivated by the work of Abraham, Castillo, and Roquain (2024) and driven by the unresolved questions from Salomond (2017), we explore the multiple hypothesis testing problem in the sparse Gaussian sequence model where we have an observation vector, say, $\mathbf{X}=(X_{1},\dots,X_{n})\in\mathbb{R}^{n}$, such that each $X_{i}$ can be expressed as
$X_{i}=\theta_{i}+\epsilon_{i},\textrm{ for }i=1,\dots,n.$ (1)
In (1) above, the parameters $\theta_{1},\dots,\theta_{n}$ are unknown mean parameters, and the unobserved residuals $\epsilon_{i}$’s are assumed to be independent and identically distributed $N(0,\sigma^{2})$ random variables. For theoretical developments, the residual variance $\sigma^{2}$ is typically assumed to be known. Hence, without any loss of generality, one may assume $\sigma^{2}$ to be $1$. For the above normal means model, we are interested in testing the hypothesis
$\textrm{H}_{0i}:\theta_{i}=0\textrm{ vs. }\textrm{H}_{1i}:\theta_{i}\neq 0,\textrm{ simultaneously for }i=1,\ldots,n.$
We adopt the asymptotic framework outlined by Abraham, Castillo, and Roquain (2024) where the parameter vector $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{n})$ is assumed to be sparse, that is, it belongs to the set
[…]
consisting of vectors that have at most $q_{n}$ many nonzero coordinates, where $0\leq q_{n}\leq n$. Throughout the paper, we consider the sparse asymptotic framework where
$q_{n}\rightarrow\infty,\ \textrm{and}\ \dfrac{q_{n}}{n}\rightarrow 0\textrm{ as }n\rightarrow\infty.$
We focus on the same broad class of priors previously considered in Ghosh et al. (2016); Ghosh and Chakrabarti (2017) when each unknown $\theta_{i}$ is modeled as a scale mixture of normals, as,
$\theta_{i}|\lambda_{i},\tau$ $\stackrel{{\scriptstyle ind}}{{\sim}}N(0,\lambda^{2}_{i}\sigma^{2}\tau^{2})$
$\lambda^{2}_{i}$ $\stackrel{{\scriptstyle ind}}{{\sim}}\pi_{1}(\lambda^{2}_{i})=K(\lambda^{2}_{i})^{-a-1}L(\lambda^{2}_{i})$ (2)
$\tau$ $\sim\pi_{2}(\tau),$
for some choices of $\pi_{2}$. In (1) above, $K>0$ is the constant of proportionality, $a$ is a positive real number, and $L\colon(0,\infty)\mapsto(0,\infty)$ is a measurable, continuous, and slowly varying function. Considering the composite loss $\textrm{FDR}+\textrm{FNR}$ and the hamming loss, our aim in this paper is to investigate whether the multiple testing rules (formally stated in (18), (19), and (21)), based on the above class of one-group shrinkage priors, asymptotically achieve the minimax risk under appropriate conditions. Our theoretical results indicate that one-group shrinkage priors of the form described in (1) cannot achieve the desired minimax risk—regardless of the loss functions mentioned—if $a>0.5$. Therefore, we confine our attention to the subclass of horseshoe-type priors, where $a=0.5$. When the number of non-zero signals is known, we theoretically demonstrate that the decision rule based on horseshoe-type priors can asymptotically attain the minimax risk with a suitable choice of the global shrinkage parameter, dependent on the proportion of non-zero means, for both the $\textrm{FDP}+\textrm{FNP}$ and classification loss functions examined by Abraham, Castillo, and Roquain (2024). In cases where the level of sparsity is unknown, we consider adaptive empirical Bayes and full Bayes decision rules based on the horseshoe-type priors and establish that these adaptive decision rules achieve the corresponding minimax risks asymptotically, as the dimension grows to infinity. Our findings provide important guidelines for selecting priors, showing that beyond a certain subclass of our chosen class—specifically, the horseshoe-type priors—the minimax risk cannot be achieved concerning either of the two loss functions discussed in this study. Nevertheless, this subclass of horseshoe-type priors is quite extensive, encompassing various prior distributions such as the horseshoe, Strawderman-Berger, standard double Pareto, and inverse-gamma priors with $\alpha=0.5$, among others. In summary, alongside well-known methods such as the Benjamini-Hochberg (BH) procedure and the $\ell$-value-based approach utilizing spike and slab priors, a multiple testing rule based on one-group shrinkage priors also achieves the optimal boundary. To the best of our knowledge, these findings represent the first results in the literature demonstrating that global-local priors can guarantee the exact achievement of optimal minimax risk in the context of simultaneous hypothesis testing problem.
The remainder of the paper is organized as follows. Section 2 reviews existing minimax results in Gaussian sequence models. Section 3 presents our main theoretical contributions, with Subsection 3 focusing on classification loss and Subsection 3.2 on composite loss (FDR+FNR). Section 4 provides concluding remarks and future directions. Proofs are deferred to Section 5.
1.1 Notations
For any two sequences $\{a_{n}\}$ and $\{b_{n}\}$ with $\{b_{n}\}\neq 0$, we write $a_{n}\asymp b_{n}$ to denote that there exist two constants $c_{1}$ and $c_{2}$ such that $0<c_{1}\leq a_{n}/b_{n}\leq c_{2}<\infty$ for sufficiently large $n$. Similarly, $a_{n}\lesssim b_{n}$ denotes that for sufficiently large $n$, there exists a constant $c>0$ such that $a_{n}\leq cb_{n}$. Similarly, $a_{n}\propto b_{n}$ implies there exists some constant $C>0$ such that $a_{n}=Cb_{n}$ for all $n$. Finally, $a_{n}=o(b_{n})$ denotes $\lim_{n\to\infty}\frac{a_{n}}{b_{n}}=0$. $\phi(\cdot)$ denotes the density of a standard normal distribution.
2 Existing Results on Relevant Minimax Risk Theory
The most recent and influential contributions on minimax risk for the Gaussian sequence model are due to Abraham, Castillo, and Roquain (2024). Working under Gaussian assumptions, they introduced the beta-min condition, designed to separate the null from the alternative, defined as follows: given $\tilde{a}\in\mathbb{R}$, set
$\Theta=\Theta(\tilde{a},q_{n})=\{\theta\in\ell_{0}[q_{n}]:|\theta_{i}|\geq\tilde{a}\ \text{for all }i\in S_{\theta},\ |S_{\theta}|=q_{n}\},$ (3)
where $S_{\theta}=\{i:\theta_{i}\neq 0\}$. Motivated by earlier works of Arias-Castro and Chen (2017) and Rabinovich et al. (2020), they chose $\tilde{a}$ to be close to the universal threshold $\sqrt{2\log(n/q_{n})}$ and considered
$\Theta_{b}=\Theta(\tilde{a}_{b},q_{n}),\qquad\tilde{a}_{b}=\sqrt{2\log\left(\frac{n}{q_{n}}\right)}+b,$ (4)
with $b\in\mathbb{R}$. They established that for any $b\in\mathbb{R}$,
$\inf_{\psi}\sup_{\theta\in\Theta_{b}}R(\theta,\psi)=1-\Phi(b)+o(1),$ (5)
where $\Phi(\cdot)$ is the standard normal c.d.f. and $R(\theta,\psi)$ is defined as
$R(\boldsymbol{\theta,\psi})$ $=FDR(\boldsymbol{\theta,\psi})+FNR(\boldsymbol{\theta,\psi}).$ (6)
This result holds even when $b=b_{n}\to\infty$ or $b=b_{n}\to$-$\infty$. As a byproduct, they showed that the oracle thresholding rule
$\psi_{i}=\mathbf{1}\left\{|X_{i}|\geq\sqrt{2\log(n/q_{n})}\right\},\ \textrm{ for }i=1,\dots,n,$
is asymptotically minimax. Moreover, they identified the sharp boundary for nontrivial testing: when $b\to$-$\infty$, no multiple testing procedure performs better than trivial rules (rejecting all or none). In contrast, finite values of $b$ yield nontrivial minimax risk.
Since the oracle rule depends on $q_{n}$, which is typically unknown, an important question is whether there exist data-driven multiple testing rules that achieve the minimax risk in (5). Assuming $q_{n}\lesssim n^{c}$ for some $0<c<1$, Abraham, Castillo, and Roquain (2024) proved that both the Benjamini–Hochberg (BH) procedure of Benjamini and Hochberg (1995) with appropriately chosen $\alpha_{n}\to 0$, and the $\ell$-value procedure under spike-and-slab priors with a suitably chosen slab, achieve the minimax bound for finite $b$ or diverging $b=b_{n}\to\infty$.
Specifically, under the two-groups spike-and-slab prior
$\theta_{i}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}(1-p)\delta_{0}+pF,\qquad i=1,\dots,n,$ (7)
where $\delta_{0}$ is a point mass at zero and $F$ is a non-degenerate slab distribution, the marginal distribution of $X_{i}$ is
$X_{i}\sim(1-p)\varphi(x)+p\int_{\mathbb{R}}\varphi(x-\theta)F(\theta)\,d\theta,$ (8)
with $\varphi(\cdot)$ denoting the standard normal density. Choosing a quasi-Cauchy slab, they defined the posterior $\ell$-value
$\ell_{i,p}(X)=\Pi_{p}(\theta_{i}=0\mid X)=\frac{(1-p)\varphi(X_{i})}{(1-p)\varphi(X_{i})+p\tilde{f}(X_{i})},$
and rejected $H_{0i}$ when $\ell_{i,p}(X)<t$ for some threshold $t\in(0,1)$. Since $p$ is unknown, they used the marginal maximum likelihood estimate (MMLE) $\hat{p}=\arg\max_{p\in[1/n,1]}L(p)$, where $L(p)$ is the likelihood of $X$. The decision rule $\psi_{\hat{\ell}}(X)$ was shown to achieve the minimax risk.
Abraham, Castillo, and Roquain (2024) also relaxed Gaussian assumptions. For example, under weaker noise conditions — assuming observations follow a continuous distribution with a Lipschitz cumulative distribution function (c.d.f.)— and for heterogeneous signals, they extended the minimax framework. For signals of varying strengths,
$\Theta(\tilde{a},q_{n})=\Big\{\theta\in\ell_{0}[q_{n}]:\textrm{there exists }i_{1},\dots,i_{q_{n}}\ \text{distinct},\ |\theta_{i_{j}}|\geq\tilde{a}_{j},\ j=1,\dots,q_{n}\Big\},$ (9)
with $\tilde{a}=(\tilde{a}_{1},\dots,\tilde{a}_{q_{n}})$, they introduced
$\Lambda_{n}(\tilde{a})=\frac{1}{q_{n}}\sum_{j=1}^{q_{n}}F_{\tilde{a}_{j}}(a_{n}^{*}),\qquad a_{n}^{*}=\left(\zeta\log\left(\frac{n}{q_{n}}\right)\right)^{1/\zeta},\quad\zeta>1,$
and proved that
$\inf_{\psi}\sup_{\theta\in\Theta(\tilde{a},q_{n})}R(\theta,\psi)$ $=\Lambda_{n}(\tilde{a})+o(1),$ (10)
$\inf_{\psi}\sup_{\theta\in\Theta(\tilde{a},q_{n})}R(\theta,\psi)$ $=2\Lambda_{n}(\tilde{a})-1+o(1),\quad\text{(scale models).}$ (11)
For the normal means model, $\Lambda_{n}(\tilde{a})=\frac{1}{q_{n}}\sum_{j=1}^{q_{n}}[1-\Phi(\tilde{a}_{j}-a_{n}^{*})]$ with $a_{n}^{*}=\sqrt{2\log(n/q_{n})}$. They showed that both BH (with $\alpha_{n}\to 0$) and $\ell$-value rules achieve this bound, even when $q_{n}$ is unknown.
Furthermore, Abraham, Castillo, and Roquain (2024) studied minimax risk for classification (Hamming) loss:
$L(\theta,\psi)=\sum_{i=1}^{n}\left[\mathbf{1}\{\theta_{i}=0,\psi_{i}\neq 0\}+\mathbf{1}\{\theta_{i}\neq 0,\psi_{i}=0\}\right].$ (12)
Earlier, Butucea et al. (2018, 2023) introduced the concept of almost full recovery under Hamming loss, defined as
$\frac{1}{q_{n}}\sup_{\theta\in\Theta_{q_{n}}}\mathbb{E}L(\theta,\psi)=o(1),\qquad n\to\infty.$ (13)
For the Gaussian location model, Abraham, Castillo, and Roquain (2024) established
$\frac{1}{q_{n}}\inf_{\psi}\sup_{\theta\in\Theta_{b}}\mathbb{E}L(\theta,\psi)=1-\Phi(b)+o(1),$ (14)
valid for both BH and $\ell$-value rules under polynomial sparsity.
The only prior work on one-group priors and minimax risk in normal means is due to Salomond (2017). Modeling each $\theta_{i}$ under
$\theta_{i}\mid\sigma_{i}^{2}\sim N(0,\sigma_{i}^{2}),\quad\sigma_{i}^{2}\sim\pi(\sigma_{i}^{2}),$ (15)
they proposed thresholding rules of the form $\psi_{i}=\mathbf{1}\{\mathbb{E}(\phi_{i}\mid X)>\alpha\}$ with $\phi_{i}=\sigma_{i}^{2}/(1+\sigma_{i}^{2})$. Inspired by Arias-Castro and Chen (2017) and Rabinovich et al. (2020), they bounded minimax risk for FDR+FNR under additional assumptions on sparsity. Their results showed consistency when signals exceeded a “separation condition,” but the boundary was stricter than that of Abraham, Castillo, and Roquain (2024). Moreover, they did not consider classification loss, heterogeneous signals, or unknown sparsity. As such, whether one-group priors can achieve minimaxity under the sharp beta-min condition remains an open problem. Revisiting this question is the focus of our work in the next section.
3 Major Theoretical Results
Inspired by the works of Abraham, Castillo, and Roquain (2024) and Salomond (2017), we were interested in studying the multiple testing problem $H_{1i}:\theta_{i}\neq 0$ simultaneously, when each $\theta_{i}$ is modeled as a scale mixture of normals as described in (1). Note that, in the hierarchical formulation of (1), the posterior mean of $\theta_{i}$ is obtained as
$\mathbb{E}(\theta_{i}|\mathbf{X},\tau,\sigma)=(1-\mathbb{E}(\kappa_{i}|X_{i},\tau,\sigma))X_{i},$ (16)
where $\kappa_{i}=1/(1+\lambda^{2}_{i}\tau^{2})$. As a result, in the literature of one-group priors, $\mathbb{E}(\kappa_{i}|X_{i},\tau,\sigma)$ is interpreted as a shrinkage coefficient and hence such priors are also known as one-group shrinkage priors. Out of the two sets of hyperparameters ($\lambda_{i}$ and $\tau$) present in the hierarchy (1), $\lambda_{i}$’s are named as local shrinkage parameters and $\tau$ as global shrinkage parameter. Hence, alternatively, one-group priors are also termed as global-local shrinkage priors. Different choices of $\pi_{1}(\cdot)$ proposed by several authors have enriched the literature over the years. Some of the examples of such priors are due to t-prior (Tipping (2001)), the Laplace prior (Park and Casella (2008)), the normal-exponential-gamma prior (Griffin and Brown (2005)), horseshoe prior (Carvalho, Polson and Scott (2009), Carvalho, Polson and Scott (2010), Polson and Scott (2010), Polson and Scott (2012)), three parameter beta normal priors (Armagan, Clyde and Dunson (2011)), generalized double pareto priors (Armagan, Dunson and Lee (2013)), the Dirichlet-laplace prior (Bhattacharya et al. (2015)), horseshoe+ prior (Bhadra et al. (2017)) etc.
In order to mimic the optimal properties (either in terms of estimation or testing) of its two-groups counterparts, this class of priors should have the dual nature of squelching the noises towards zero along with keeping intact the true signals. In this hierarchical formulation, the global parameter $\tau$ plays the level of sparsity and hence a smaller value of it ensures a large mass around the origin; on the other hand, $\lambda_{i}$’s are used to detect the signals present in the model. This results in noise observations being shrunk towards zero while the real signals are left unshrunk. This is known as tail robustness and hence the priors are named as tail robust priors.
Using a simulation study, Carvalho, Polson and Scott (2009) at first provides an indication that $\mathbb{E}(1-\kappa_{i}|X_{i},\tau)$ with $\kappa_{i}=1/(1+\lambda^{2}_{i}\tau^{2})$ mimics the role of the posterior inclusion probability ($\Pi(\theta_{i}\neq 0|\mathbf{X})$) corresponding to the spike and slab prior. They proposed the following decision rule intuitively,
$\text{Reject}\hskip 1.42271ptH_{0i}\text{ if }\mathbb{E}(1-\kappa_{i}|X_{i},\tau)>\frac{1}{2},\text{ for }i=1,\dots,n.$ (17)
Later Datta and Ghosh (2013) and Ghosh and Chakrabarti (2017) established optimality of the decision rule (17) in the context of minimizing the Bayes risk under 010-1 loss function. Motivated by these works, in order to study the minimax risk of different loss functions based on one-group priors for the normal means problem, we also consider the decision rule $\psi=(\psi_{1},\psi_{2},\cdots,\psi_{n})$, where for $i=1,\dots,n$,
$\psi_{i}\equiv\psi_{i}(X_{i})=\mathbf{1}\{\mathbb{E}(1-\kappa_{i}|X_{i},\tau)>\frac{1}{2}\}.$ (18)
Based on the decision rule (18), we want to investigate whether any global-local prior of the form (1) can attain the minimax risk either based on Hamming loss or based on the sum of FDR and FNR, irrespective of the level of sparsity being known or not.
For the theoretical study of the decision rule (18), we assume the slowly varying function $L(\cdot)$ defined in (1) satisfies Assumption 1.
Assumption 1.
For $a\geq\frac{1}{2}$:
(A1) There exists some $c_{0}(>0)$ such that $L(t)\geq c_{0}\;\forall t\geq t_{0}$, for some $t_{0}>0$, which depends on both $L$ and $c_{0}$.
(A2) There exists some $M\in(0,\infty)$ such that $\sup_{t\in(0,\infty)}L(t)\leq M$.
It is evident from earlier results of Datta and Ghosh (2013), van der Pas, Kleijn and van der Vaart (2014), Ghosh et al. (2016), and Ghosh and Chakrabarti (2017) that the global shrinkage parameter $\tau$ plays a pivotal role in capturing the underlying sparsity of the model. However, the sparsity level is usually unknown. Since the decision rule (18) also contains $\tau$, it is of equal importance to propose a modified version of it when the level of sparsity is unknown. Using the empirical Bayes estimate of $\tau$, proposed by van der Pas, Kleijn and van der Vaart (2014), the decision rule modifies to, for $i=1,\dots,n$,
$\psi_{i}\equiv\psi^{\text{EB}}_{i}(\mathbf{X})=\mathbf{1}\{\mathbb{E}(1-\kappa_{i}|X_{i},\widehat{\tau})>\frac{1}{2}\},$ (19)
where
$\widehat{\tau}=\max\{\frac{1}{n},\frac{1}{c_{2}n}\sum_{i=1}^{n}\mathbf{1}(|X_{i}|>\sqrt{c_{1}\log n})\},$ (20)
with $c_{1}\geq 2$ and $c_{2}\geq 1$. An alternative approach is to model $\tau$ by an absolutely continuous prior $\pi(\tau)$ on it, known as the full Bayes approach. Motivated by the previous work of Paul and Chakrabarti (2025), here, we consider a general class of priors $\pi(\cdot)$ satisfying the following condition:
(C4) $\int_{\frac{1}{n}}^{\alpha_{n}}\pi(\tau)d\tau=1$ for some $\alpha_{n}$ such that $\alpha_{n}\to 0$ and $n\alpha_{n}\to\infty$ as $n\to\infty$.
In this procedure, the decision rule is modified as, for $i=1,\dots,n$,
$\psi_{i}\equiv\psi^{\text{FB}}_{i}(\mathbf{X})=\mathbf{1}\{\mathbb{E}(1-\kappa_{i}|\mathbf{X})>\frac{1}{2}\},$ (21)
where the full Bayes posterior mean of $1-\kappa_{i}$ is defined as,
$\mathbb{E}(1-\kappa_{i}|\mathbf{X})$ $=\int_{\frac{1}{n}}^{\alpha_{n}}\mathbb{E}(1-\kappa_{i}|\mathbf{X},\tau)\pi(\tau|\mathbf{X})d\tau=\int_{\frac{1}{n}}^{\alpha_{n}}\mathbb{E}(1-\kappa_{i}|X_{i},\tau)\pi(\tau|\mathbf{X})d\tau.$
We will study the optimality of the decision rules (18), (19), and (21) depending on the level of sparsity to be known or not, for both classification loss and the loss measured as the sum of FDR and FNR. Results of the minimax risk related to the Hamming or classification loss using one-group priors are discussed in Subsection 3.1. On the other hand, Subsection 3.2 contains results based on minimax risk corresponding to the loss defined as the sum of FDR and FNR using one-group shrinkage priors.
We end this Section with a Proposition based on a lower bound on the probability of type II error corresponding to the $i^{th}$ hypothesis. This shows that why $a>\frac{1}{2}$ in (1) can not be used as a good choice for the broad class of one-group priors considered here.
Proposition 1.
Suppose $X_{i}\stackrel{{\scriptstyle ind}}{{\sim}}N(\theta_{i},1),i=1,\dots,n$. We want to test hypotheses $H_{0i}:\theta_{i}=0$ vs $H_{1i}:\theta_{i}\neq 0$ simultaneously based on the decision rule (18) using one-group priors. Assume that $\tau\to 0$ as $n\to\infty$ such that $\frac{n\tau}{q_{n}}\to C\in(0,\infty)$. Also assume that slowly varying function $L(\cdot)$ satisfies Assumption 1. Then for $a>\frac{1}{2}$ in (1), for $i=1,\ldots,q_{n}$ and $b\in\mathbb{R}$, we have,
$\sup_{|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b}t_{2i}\to 1,\textrm{ as }n\to\infty,$ (22)
where $t_{2i}$ denotes the type II error measure corresponding to $i^{th}$ hypothesis.
Observe that for Theorems 1 and 2, the loss function $L(\boldsymbol{\theta},\psi)$ for which the minimax risk has to be obtained, depends on $t_{1i}$ and $t_{2i}$, the type I and type II error measures corresponding to the $i^{th}$ hypothesis, respectively. Later, it will be clear from the proofs of these two theorems that out of type I and type II errors, the type II error is the dominating one. Hence, the lower bound on the type II error measure over the subset $\boldsymbol{\Theta}_{b}$ for $a>\frac{1}{2}$ implies the overall risk also exceeds the minimax risk, $1-\Phi(b)+o(1)$, in this case. On the other hand, when one is interested in studying the minimax risk based on the loss, defined as the sum of FDR and FNR, then also, the proofs of Theorems 4 and 5 imply that those mainly depend on the FNR, which is as same as the type II error. Hence, the conclusion from Proposition 1 goes through here, too. As a consequence, Proposition 1 acts as a guideline for us regarding the choice of $a$ given in (1) in the subsequent subsections in order to obtain the minimax risk.
3.1 Results Based On the Classification Loss
Motivated by the result of Proposition 1, from this Subsection, We are interested in the special case when $a=\frac{1}{2}$ in (1). This subclass of priors is named as horseshoe-type priors. Note that it is a rich subclass which contains the three parameter beta normal mixtures with $\alpha=0.5$ and $\beta>0$ (e.g. horseshoe, Strawderman-Berger), the generalized double Pareto priors with $\alpha=1$ (e.g. standard double Pareto), and the inverse–gamma priors with $\alpha=0.5$, etc.
As stated earlier, we are interested in studying the optimality of our decision rule (18) based on horseshoe-type priors in the sense that whether (18) can attain the minimax risk over the set $\boldsymbol{\Theta}_{b}$ or not. At first, we assume the level of sparsity is known and use the global parameter $\tau$ as a tuning one. We propose a condition on $\tau$, depending on the level of sparsity. The first result of our work is as follows:
Theorem 1.
Suppose $X_{i}\stackrel{{\scriptstyle ind}}{{\sim}}N(\theta_{i},1),i=1,\dots,n$. We want to test hypotheses $H_{0i}:\theta_{i}=0$ vs $H_{1i}:\theta_{i}\neq 0$ simultaneously based on the decision rule (18) using one-group priors. Assume that $\tau\to 0$ as $n\to\infty$ such that $\frac{n\tau}{q_{n}}\to C\in(0,\infty)$. Also assume that slowly varying function $L(\cdot)$ satisfies Assumption 1. Then for the class of horseshoe-type priors (i.e. for $a=\frac{1}{2}$), we have
$\frac{1}{q_{n}}\sup_{\boldsymbol{\theta}\in\boldsymbol{\Theta}_{b}}\mathbb{E}L(\boldsymbol{\theta},\psi)=1-\Phi(b)+o(1),$ (23)
where $L(\boldsymbol{\theta},\psi)$ is same as defined in (12) and $\boldsymbol{\Theta}_{b}$ is defined in (4). This result holds for any $b\in\mathbb{R}$ or $b=b_{n}\to\infty$.
Remark 1.
The importance of Theorem 1 lies in proposing a decision rule based on one-group priors which attains the minimax risk corresponding to the classification or Hamming loss. To the best of our knowledge, it is the first result in the literature of global-local priors in this context. In order to establish this result, the only condition one needs on $\tau$ is that $\tau$ is of the same order as $\frac{q_{n}}{n}$, i.e., the proportion of non-null signals, asymptotically. The same condition was previously used by Ghosh et al. (2016) while they obtained the optimality of the same testing rule with respect to the Bayes risk for the normal means model assuming the observations are generated from a two-groups spike and slab distribution. Note that, the result is established via providing bounds on the type I and type II error rates. For the type I error rate, Theorem 6 of Ghosh et al. (2016) comes in handy. However, the calculation for the type II error rate involves the use of some technical observations and many non-trivial arguments. In a nutshell, this result illustrates that, like the BH procedure of Benjamini and Hochberg (1995) and the $\ell$-value approach based on two-groups spike-and-slab priors, a decision rule (18) based on a broad class of global-local prior can also attain the minimax risk for the normal means model under the assumption that the sparsity level is known.
The main drawback of Theorem 1 is that it assumes the sparsity of the underlying model to be known, which holds seldom. This motivates us to consider the adaptive decision rule (19) and study the same in terms of the minimax risk. Now consider the following Theorem.
Theorem 2.
Consider the setup of Theorem 1. Suppose we want to test hypotheses $H_{0i}:\theta_{i}=0$ vs $H_{1i}:\theta_{i}\neq 0$ simultaneously based on the decision rule (19) using one-group priors, where $\widehat{\tau}$ is the same as (20). Assume that the number of non-zero means $q_{n}$ satisfies $q_{n}\propto(\log n)^{\delta_{2}}$ for some $\delta_{2}>1$. Also assume that slowly varying function $L(\cdot)$ satisfies Assumption 1. Then for the class of horseshoe-type priors (i.e. for $a=\frac{1}{2}$), we have
$\frac{1}{q_{n}}\sup_{\boldsymbol{\theta}\in\boldsymbol{\Theta}_{b}}\mathbb{E}L(\boldsymbol{\theta},\psi)=1-\Phi(b)+o(1),$ (24)
where $L(\boldsymbol{\theta},\psi)$ and $\Theta_{b}$ are same as defined in Theorem 1. This result holds for any $b\in\mathbb{R}$ or $b=b_{n}\to\infty$.
Remark 2.
[…]
Theorem 3.
Consider the setup of Theorem 1. Suppose we want to test hypotheses $H_{0i}:\theta_{i}=0$ vs $H_{1i}:\theta_{i}\neq 0$ simultaneously based on the decision rule (21) using one-group priors, where the prior on $\tau$ satisfies (C4) with $\alpha_{n}\propto\frac{(\log n)^{\delta_{3}}}{n}$ for some $0<\delta_{3}<\frac{1}{2}$. Assume that the number of non-zero means $q_{n}$ satisfies $q_{n}\propto(\log n)^{\delta_{2}}$ for some $\delta_{2}>0$. Also assume that slowly varying function $L(\cdot)$ satisfies Assumption 1. Then for horseshoe-type priors (i.e. for $a=\frac{1}{2}$), we have
$\frac{1}{q_{n}}\sup_{\boldsymbol{\theta}\in\boldsymbol{\Theta}_{b}}\mathbb{E}L(\boldsymbol{\theta},\psi)=1-\Phi(b)+o(1),$ (25)
where $L(\boldsymbol{\theta},\psi)$ and $\Theta_{b}$ are same as defined in Theorem 1. This result holds for any $b\in\mathbb{R}$ or $b=b_{n}\to\infty$.
3.2 Results Based On the Composite Loss FDR+FNR
As alluded to earlier, in this Subsection, too, we confine our interest to studying whether the decision rules based on horseshoe-type priors (18) and (19) can achieve the minimax risk asymptotically, irrespective of whether the level of sparsity is known or not. Similar to the previous subsection, here, too, at first we assume the level of sparsity to be known and try to model $\tau$ accordingly. Our result provides an affirmative conclusion that, similar to the BH and $l$-value procedures, as stated by Abraham, Castillo, and Roquain (2024), the decision rule (18) can also attain the minimax risk assuming the proportion of non-zero means to be known. The result is stated formally below.
Theorem 4.
Suppose $X_{i}\stackrel{{\scriptstyle ind}}{{\sim}}N(\theta_{i},1),i=1,\dots,n$. We want to test hypotheses $H_{0i}:\theta_{i}=0$ vs $H_{1i}:\theta_{i}\neq 0$ simultaneously based on the decision rule (18) using one-group priors. Assume that $\tau\to 0$ as $n\to\infty$ such that $\frac{n\tau}{q_{n}}\to C\in(0,\infty)$. Also assume that slowly varying function $L(\cdot)$ satisfies Assumption 1. Then for horseshoe-type priors (i.e. for $a=\frac{1}{2}$), we have
$\sup_{\boldsymbol{\theta}\in\boldsymbol{\Theta}_{b}}R(\boldsymbol{\theta},\psi)=1-\Phi(b)+o(1),$ (26)
where $\Theta_{b}$ is same as defined in Theorem 1 and $R(\boldsymbol{\theta},\psi)$ is defined in (6). This result continues to hold for any $b\in\mathbb{R}$ or $b=b_{n}\to\infty$.
Remark 3.
Similar to the results of the previous Subsection, this result also justifies the selection of the decision rule (18) based on one-group priors since it attains the minimax risk for the loss function defined as the sum of FDR and FNR. As stated earlier, the only work in the context of obtaining an upper bound to the minimax risk based on this loss induced by one-group priors was considered by Salomond (2017). However, their choice of the minimum value of the non-zero means exceeds the universal threshold. Not only this, but they also did not provide any answer regarding whether their proposed decision rule can achieve the minimax risk. This result provides a positive answer to the question for the horseshoe-type priors corresponding to the minimax risk when the supremum of $\boldsymbol{\theta}$ is considered over a set that contains the universal threshold. An advantage of this result is that the upper bound can be made arbitrarily small by taking $b=b_{n}\to\infty$. The proof is done by providing bounds on FDR and FNR. The truncation used in bounding FDR is inspired by their approach. However, we have to do delicate calculations after that. We need to use Hoeffding’s inequality to bound some quantity related to FDR. The technique used in providing a bound of FNR is quite nontrivial from that of Salomond (2017). In that context, bound on type II error, as derived in previous results, comes in useful. Our proof also demonstrates that some calculations used in Salomond (2017) can be performed using a simpler approach as well.
Given that (18) achieves the minimax risk, it is natural to ask whether the adaptive decision rule (19), too, can provide the same result, when the number of non-zero means is unknown. The following result provides an affirmative answer to the question.
Theorem 5.
Consider the setup of Theorem 1. Suppose we want to test hypotheses $H_{0i}:\theta_{i}=0$ vs $H_{1i}:\theta_{i}\neq 0$ simultaneously based on the decision rule (19) using one-group priors, where $\widehat{\tau}$ is the same as defined in van der Pas, Kleijn and van der Vaart (2014). Assume that the number of non-zero means $q_{n}$ satisfies $q_{n}\propto(\log n)^{\delta_{2}}$ for some $\delta_{2}>1$. Also assume that slowly varying function $L(\cdot)$ satisfies Assumption 1. Then for horseshoe-type priors (i.e. for $a=\frac{1}{2}$), we have
$\sup_{\boldsymbol{\theta}\in\boldsymbol{\Theta}_{b}}R(\boldsymbol{\theta},\psi)=1-\Phi(b)+o(1),$ (27)
where $\Theta_{b}$ is same as defined in Theorem 1 and $R(\boldsymbol{\theta},\psi)$ is defined in (6). This result holds for any $b\in\mathbb{R}$ or $b=b_{n}\to\infty$.
Remark 4.
This result elucidates that, like the BH procedure of Benjamini and Hochberg (1995) and the $\ell$-value approach based on two-groups spike-and-slab priors, a decision rule (19) based on a broad class of global-local prior can also attain the minimax risk for the normal means model, even if the sparsity level is unknown. This type of result was earlier considered by Salomond (2017) for a broader class of priors, which contains our chosen class of priors. They proved that the upper bound of the minimax risk for their proposed decision rule converges to 0 as $n\to\infty$. However, their work did not provide any answer to the question of whether their decision rule can attain the minimax risk or not. On the contrary, our result does provide a positive answer to that question. As a byproduct of our result, the minimax risk can be made arbitrarily small by choosing $b=b_{n}\to\infty$, even in the case of an unknown sparse situation. In this way, their result becomes a special case of ours under our chosen class of priors under study.
Theorem 6.
Consider the setup of Theorem 1. Suppose we want to test hypotheses $H_{0i}:\theta_{i}=0$ vs $H_{1i}:\theta_{i}\neq 0$ simultaneously based on the decision rule (21) using one-group priors, where the prior on $\tau$ satisfies (C4) with $\alpha_{n}\propto\frac{(\log n)^{\delta_{3}}}{n}$ for some $0<\delta_{3}<\frac{1}{2}$. Assume that the number of non-zero means $q_{n}$ satisfies $q_{n}\propto(\log n)^{\delta_{2}}$ for some $\delta_{2}>0$. Also assume that slowly varying function $L(\cdot)$ satisfies Assumption 1. Then for horseshoe-type priors (i.e. for $a=\frac{1}{2}$), we have
$\sup_{\boldsymbol{\theta}\in\boldsymbol{\Theta}_{b}}R(\boldsymbol{\theta},\psi)=1-\Phi(b)+o(1),$ (28)
where $\Theta_{b}$ is same as defined in Theorem 1 and $R(\boldsymbol{\theta},\psi)$ is defined in (6). This result holds for any $b\in\mathbb{R}$ or $b=b_{n}\to\infty$.
4 Concluding Remarks
Due to the availability of large datasets related to several fields like Astronomy, Bioinformatics, Image Processing, and many more, multiple hypothesis testing is a serious problem of concern for researchers. Hence, the main task is to control the overall type I error of the problem, the most famous one being the FDR. Given that a procedure controls FDR, it is obvious to ask about its performance in controlling FNR, i.e., the optimal value of the risk corresponding to the loss defined as the sum of FDR and FNR. Another related problem of concern is the optimal risk of a testing procedure based on classification loss. For the normal means model, Abraham, Castillo, and Roquain (2024) obtained the minimax risk for both problems and established that the BH procedure of Benjamini and Hochberg (1995) and the $\ell$-value procedure using a spike and slab prior attain the risk, even if the sparsity pattern is unknown. This raises the question of whether any continuous one-group prior can attain the minimax risk or not. In this work, we provide a positive answer to this question by considering a broad class of one-group priors under study. To the best of our knowledge, these are the first results of their kind. As a consequence of our results, practitioners are now free to choose one-group priors as an alternative to their two-group counterparts with greater confidence.
Though our results are promising and ensure a general class of continuous priors can be considered as an alternative to the ’gold standard’ spike-and-slab priors, there are some questions that still need to be addressed from a decision theoretic viewpoint. First, when the level of sparsity is unknown, the most general estimate of $\tau$ as stated by van der Pas, Szabo and van der Vaart (2017) is the MMLE of $\tau$, defined as the maximizer of the marginal likelihood function of $\mathbf{X}$ given $\tau$. So, it is natural to ask whether the decision rule (19) using the MMLE of $\tau$ can attain the minimax risk for the two loss functions considered here. Hence, one might be interested in studying the optimality of this procedure in this context.
Another possible extension of our work is studying the class of priors proposed by van der Pas, Salomond, and Schmidt-Heiber (2016). For that class of priors, first, it is needed to prove that, assuming the level of sparsity to be known, whether any multiple testing rule can attain the minimax risk, asymptotically, under the beta-min condition of Abraham, Castillo, and Roquain (2024). The question regarding the optimality of a testing procedure for an unknown level of sparsity remains open to this day. All of these can be considered as some problems of interest to be discussed in detail in the future.
5 Proofs
Proof of Proposition 1.
Let us fix any $i\geq 1$. We consider the cases $a\in(\frac{1}{2},1)$ and $a\geq 1$, separately.
Case 1: Let us first consider the case when $a\in(\frac{1}{2},1)$. Using Theorem 4 of Ghosh et al. (2016), and the fact that the slowly varying function $L(\cdot)$ is bounded above by some constant M(>0)M\ (>0), we have
$E(1-\kappa_{i}\big|X_{i},\tau)\leq K_{1}e^{\frac{X_{i}^{2}}{2}}\tau^{2a}(1+o(1)),$ (29)
where the $o(1)$ term above is independent of both the index $i$ and the data point $X_{i}$, but depends on $\tau$ in such a way that $\lim_{\tau\rightarrow 0}o(1)=0$. Now, applying the definition of type II error coupled with (29), the type II error rate for the $i^{th}$ hypothesis testing problem is given by
$t_{2i}$ $=\mathbb{P}_{\theta_{i}\neq 0}\left(\mathbb{E}(1-\kappa_{i}|X_{i},\tau)\leq\frac{1}{2}\right)$
$\geq\mathbb{P}_{\theta_{i}\neq 0}\left(K_{1}\tau^{2a}\exp(\frac{X^{2}_{i}}{2})(1+o(1))\leq\frac{1}{2}\right)$
$=\mathbb{P}_{\theta_{i}\neq 0}\left(X^{2}_{i}\leq 4a\log(\frac{1}{\tau})(1+o(1))\right)$
$=\Phi\left(\sqrt{4a\log(\frac{1}{\tau})}(1+o(1)-\theta_{i}\right)-\Phi\left(-\sqrt{4a\log(\frac{1}{\tau})}(1+o(1)-\theta_{i}\right).$ (30)
Note that, to provide a lower bound to the type II error rate on the set $\boldsymbol{\Theta}_{b}$, it would be enough to consider only one candidate of $\theta_{i}$ from that set and show that the lower bound of (5) for that choice of $\theta_{i}$ tends to $1$ as $n\to\infty$. Let us fix any $b\in\mathbb{R}$ and choose $\theta_{i}=\sqrt{2\log(\frac{n}{q_{n}})}$. Then under the assumption that $\frac{n\tau}{q_{n}}\to C\in(0,\infty)$ as $n\rightarrow\infty$, we obtain
$\log\left(\frac{1}{\tau}\right)=\log\left(\frac{n}{q_{n}}\right)+\log\left(\frac{1}{C}\right)+\log(1+o(1))=\log\left(\frac{n}{q_{n}}\right)(1+o(1)),$ (31)
where the $o(1)$ term above is independent of both the index $i$ and the data point $X_{i}$, but depends on $\tau$ in such a way that $\lim_{\tau\rightarrow 0}o(1)=0$. This implies that for all sufficiently large $n$, we have
$\sup_{|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b}t_{2i}$ $\geq\Phi\left((\sqrt{4a}-\sqrt{2})\log\left(\frac{n}{q_{n}}\right)\right)(1+o(1))-\Phi\left((-\sqrt{4a}-\sqrt{2})\log\left(\frac{n}{q_{n}}\right)\right)(1+o(1))$
$=\Phi\left((\sqrt{4a}-\sqrt{2})\log\left(\frac{n}{q_{n}}\right)\right)(1+o(1))$
$\to 1,\textrm{ as }n\to\infty.$ (32)
In the above chain of inequalities, observe that for any fixed $a\in(\frac{1}{2},1)$, $\Phi\left((\sqrt{4a}-\sqrt{2})\log(\frac{n}{q_{n}})\right)(1+o(1))\to 1$ as $n\to\infty$ and $\Phi\left((-\sqrt{4a}-\sqrt{2})\log(\frac{n}{q_{n}})\right)(1+o(1))\to 0$ as $n\to\infty$.
Case 2: Next, we consider the case where $a\geq 1$. First, we derive an upper bound to $E(1-\kappa_{i}|X_{i},\tau)$ for any fixed $\tau\in(0,1)$ and $X_{i}\in\mathbb{R}$.
Observe that, the posterior distribution of $\kappa_{i}$ given $X_{i}$ and $\tau$ is,
[…]
Using the transformation $t=\frac{1}{\tau^{2}}(\frac{1}{\kappa}-1)$, $E(1-\kappa_{i}|X_{i},\tau)$ becomes
$E(1-\kappa_{i}|X_{i},\tau)=\frac{\tau^{2}\int_{0}^{\infty}(1+t\tau^{2})^{-\frac{3}{2}}t^{-a}L(t)\exp\bigg({\frac{X^{2}_{i}}{2}\cdot\frac{t\tau^{2}}{1+t\tau^{2}}}\bigg.)dt}{\int_{0}^{\infty}(1+t\tau^{2})^{-\frac{1}{2}}t^{-a-1}L(t)\exp\bigg({\frac{X^{2}_{i}}{2}\cdot\frac{t\tau^{2}}{1+t\tau^{2}}}\bigg.)dt}.$
Note that,
$\int_{0}^{\infty}(1+t\tau^{2})^{-\frac{1}{2}}t^{-a-1}L(t)\exp\bigg({\frac{X^{2}_{i}}{2}\cdot\frac{t\tau^{2}}{1+t\tau^{2}}}\bigg.)dt$ $\geq\int_{0}^{\infty}(1+t\tau^{2})^{-\frac{1}{2}}t^{-a-1}L(t)dt=K^{-1}(1+o(1)),$
where $o(1)$ depends only on $\tau$ and tends to zero as $\tau\to 0$. The equality in the last line follows due to the Dominated Convergence Theorem. Hence
$E(1-\kappa_{i}|X_{i},\tau)\leq K(A_{1}+A_{2})(1+o(1))\hskip 2.84544pt,$ (33)
where
$A_{1}$ $=\int_{0}^{1}\frac{t\tau^{2}}{1+t\tau^{2}}\cdot\frac{1}{\sqrt{1+t\tau^{2}}}t^{-a-1}L(t)\exp\bigg({\frac{X^{2}_{i}}{2}\cdot\frac{t\tau^{2}}{1+t\tau^{2}}}\bigg.)dt.$
and $A_{2}=\int_{1}^{\infty}\frac{t\tau^{2}}{1+t\tau^{2}}\cdot\frac{1}{\sqrt{1+t\tau^{2}}}t^{-a-1}L(t)\exp\bigg({\frac{X^{2}_{i}}{2}\cdot\frac{t\tau^{2}}{1+t\tau^{2}}}\bigg.)dt$. Since, for any $t\leq 1$ and $\tau\leq 1$, $\frac{t\tau^{2}}{1+t\tau^{2}}\leq\frac{1}{2}$, using the fact $\int_{0}^{\infty}t^{-a-1}L(t)dt=K^{-1}$,
[…]
For providing an upper bound on $A_{2}$, we divide our calculations depending on $a=1$ and $a>1$.
Next, note that, $1+t\tau^{2}\geq t\tau^{2}$ for any $t>0$ and $t\tau^{2}\geq\sqrt{t}$ if and only if $t\geq\frac{1}{\tau^{4}}$. As a result, for $a=1$, we have
$\int_{1}^{\frac{1}{\tau^{4}}}\frac{t\tau^{2}}{1+t\tau^{2}}\cdot t^{-2}L(t)dt\leq M\tau^{2}\int_{1}^{\frac{1}{\tau^{4}}}t^{-1}dt=M\tau^{2}\log(\frac{1}{\tau^{4}}),$ (35)
and
$\int_{\frac{1}{\tau^{4}}}^{\infty}\frac{t\tau^{2}}{1+t\tau^{2}}\cdot t^{-2}L(t)dt\leq M\tau^{2}\int_{\frac{1}{\tau^{4}}}^{\infty}t^{-\frac{3}{2}}dt=2M\tau^{4}.$ (36)
Hence, combining (35) and (36), for $a=1$, we obtain
$A_{2}$ $\leq 2M\exp\bigg(\frac{X^{2}_{i}}{2}\bigg.)\tau^{2}[\log(\frac{1}{\tau^{4}})+\tau^{2}]=8M\exp\bigg(\frac{X^{2}_{i}}{2}\bigg.)\tau^{2}\log(\frac{1}{\tau})(1+o(1)).$ (37)
On the other hand, for $a>1$,
$\int_{1}^{\infty}\frac{t\tau^{2}}{1+t\tau^{2}}\cdot t^{-a-1}L(t)dt\leq M\tau^{2}\int_{1}^{\infty}t^{-a}dt=\bigg(\frac{M}{(a-1)}\bigg)\tau^{2}.$ (38)
As a result, for $a>1$,
[…]
Therefore, combining (37) and (39), for $a\geq 1$, we have
$A_{2}$ $\leq K_{1}\exp\bigg(\frac{X^{2}_{i}}{2}\bigg.)\tau^{2}\log(\frac{1}{\tau})(1+o(1)),$ (40)
where $K_{1}$ is a constant depending on $M$ and $a$. Now, combining (33), (34) and (40), we obtain, for $a\geq 1$,
$E(1-\kappa_{i}|X_{i},\tau)$ $\leq K_{2}\exp\bigg(\frac{X^{2}_{i}}{2}\bigg.)\tau^{2}\log(\frac{1}{\tau})(1+o(1)),$ (41)
where $K_{2}$ is a constant depending on $K^{-1}$ and $K_{1}$. Next, again using the same technique used when $a\in(\frac{1}{2},1)$ with (41), a lower bound of $t_{2i}$ is obtained as,
$t_{2i}$ $=\mathbb{P}_{\theta_{i}\neq 0}(\mathbb{E}(\kappa_{i}|X_{i},\tau)>\frac{1}{2})$
$\geq\mathbb{P}_{\theta_{i}\neq 0}\left(K_{2}\exp\bigg(\frac{X^{2}_{i}}{2}\bigg.)\tau^{2}\log(\frac{1}{\tau})(1+o(1))\leq\frac{1}{2}\right)$
$=\mathbb{P}_{\theta_{i}\neq 0}\left(X^{2}_{i}\leq 4\log(\frac{1}{\tau})-4\log\log(\frac{1}{\tau})-o(1)-2\log 2\right)$
$=\mathbb{P}_{\theta_{i}\neq 0}\left(X^{2}_{i}\leq 4\log(\frac{1}{\tau})(1+o(1))\right)$
$=\Phi\left(\sqrt{4\log(\frac{1}{\tau})}(1+o(1)-\theta_{i}\right)-\Phi\left(-\sqrt{4\log(\frac{1}{\tau})}(1+o(1)-\theta_{i}\right).$ (42)
Next, again using the arguments used in the previous case, we have, for all sufficiently large $n$,
$\sup_{|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b}t_{2i}$ $\geq\Phi((\sqrt{4}-\sqrt{2})\log(\frac{n}{q_{n}}))(1+o(1))-\Phi((-\sqrt{4}-\sqrt{2})\log(\frac{n}{q_{n}}))(1+o(1))$
$=\Phi((\sqrt{4}-\sqrt{2})\log(\frac{n}{q_{n}}))(1+o(1))$
$\to 1,\textrm{ as }n\to\infty.$ (43)
Proof of Theorem 1.
The desired lower bound to the left-hand side of (23) follows from Theorem 7 of Abraham, Castillo, and Roquain (2024). Hence, we only need to obtain an upper bound for it. Using arguments similar to Theorem 7 of Abraham, Castillo, and Roquain (2024), note that
$\mathbb{E}L(\theta,\psi)=\sum_{i:\theta_{i}=0}t_{1i}+\sum_{i:\theta_{i}\neq 0}t_{2i},$ (44)
where $t_{1i}$ and $t_{2i}$ denote the type I and type II error rates corresponding to the $i^{th}$ hypothesis testing problem, respectively. Hence, it is enough to show that
$\frac{1}{q_{n}}\sum_{i:\theta_{i}=0}t_{1i}=o(1),$ (45)
and
$\frac{1}{q_{n}}\sum_{i:\theta_{i}\neq 0}\sup_{|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b}t_{2i}\leq 1-\Phi(b)+o(1).$ (46)
First, we prove (45). Using the architecture of the proof of Theorem 6 of Ghosh et al. (2016) with $a=\frac{1}{2}$, we have
$t_{1i}\leq\tilde{K}_{1}\frac{\tau L\left(\frac{1}{\tau^{2}}\right)}{\sqrt{\log\left(\frac{1}{\tau^{2}}\right)}}(1+o(1)),$ (47)
where $\tilde{K}_{1}$ is a positive constant that is independent of $\tau$ and $i$, and the $o(1)$ term is such that $\lim_{n\to\infty}o(1)=0$ and is independent of any $i$. Hence, using (46) and under the assumption Assumption 1 on the slowly varying function $L(\cdot)$, we get
$\frac{1}{q_{n}}\sum_{i:\theta_{i}=0}t_{1i}$ $\leq K_{1}M\frac{(n-q_{n})}{q_{n}}\frac{\tau}{\sqrt{\log(\frac{1}{\tau^{2}})}}(1+o(1))$
$\leq K_{1}M\frac{n\tau}{q_{n}}\frac{1}{\sqrt{\log(\frac{1}{\tau^{2}})}}(1+o(1))$
$\rightarrow 0,\textrm{ as }n\to\infty.$ (48)
In the last line of the above chain of inequalities, we used the facts $\frac{n\tau}{q_{n}}\rightarrow C\in(0,\infty)$ as $n\rightarrow\infty$, $\tau\rightarrow 0$ as $n\rightarrow\infty$, and $\log(\frac{1}{\tau^{2}})\rightarrow\infty$ as $\tau\rightarrow 0$. This completes the proof of (45).
Our next goal is to provide an upper bound to $t_{2i}$ that holds uniformly for each $i=1,\dots,n$. Towards that, we first use Lemma 3 of Ghosh and Chakrabarti (2017) that asserts that, for each fixed $x\in\mathbb{R}$ and any fixed $0<\tau<1$, the posterior shrinkage coefficient $\mathbb{E}(\kappa|x,\tau)$ can be bounded above by a non-negative, continuously decreasing, and measurable function $g(x,\tau,\eta,\delta)$ given by
$g(x,\tau,\eta,\delta)=g_{1}(x,\tau,\eta,\delta)+g_{2}(x,\tau,\eta,\delta),$
where
$g_{1}(x,\tau,\eta,\delta)=C_{1}\left[x^{2}\int_{0}^{\frac{x^{2}}{1+t_{0}}}\exp(-\frac{u}{2})du\right]^{-1}=C_{1}\left[x^{2}\left\{1-\exp(-\frac{x^{2}}{2(1+t_{0})})\right\}^{-1}\right],$
and
$g_{2}(x,\tau,\eta,\delta)=C_{2}\tau^{-1}\exp(-\frac{\eta(1-\delta)}{2}x^{2}),$
for any given $\eta,\delta\in(0,1)$. Here, $C_{1}$ and $C_{2}$ are two finite positive constants that do not depend on both $x$ and $\tau$ (but they do depend on $\eta$ and $\delta$), and $t_{0}$ was previously defined in Assumption 1. In addition, the above nonnegative function $g(x,\tau,\eta,\delta)$ satisfies the following:
For any $\rho>\frac{2}{\eta(1-\delta)}$,
$\lim_{\tau\rightarrow 0}\sup_{|x|>\sqrt{\rho\log(\frac{1}{\tau})}}g(x,\tau,\eta,\delta)=0.$
Let us fix any $\eta\in(0,1)$ and $\delta\in(0,1)$. Choose $\rho=\frac{2}{\eta(1-\delta)}[1+\tilde{g}(\tau)]$ where $\tilde{g}(\tau)>0$ and tends to zero as $\tau\to 0$ such that ${\tau}^{\tilde{g}(\tau)}\to 0$ as $\tau\to 0$. One possible choice of $\tilde{g}(\tau)=(\log(\frac{1}{\tau}))^{-\delta_{1}}$ for some $0<\delta_{1}<1$. Later, we may choose a more specific choice of $\tilde{g}(\tau)$, as required.
Next, we define two events $B_{n}$ and $C_{n}$ as $B_{n}=\{g(X_{i},\tau,\eta,\delta)>\frac{1}{2}\}$ and $C_{n}=\{|X_{i}|>\sqrt{\rho\log(\frac{1}{\tau})}\}$ with $\rho$ as defined before. Then, using the preceding observations, the type II error rate for the $i^{th}$ hypothesis testing problem is given by
$t_{2i}$ $=P_{\theta_{i}\neq 0}(E(\kappa_{i}|X_{i},\tau)>\frac{1}{2})$
$\leq P_{\theta_{i}\neq 0}(B_{n})$
$\leq P_{\theta_{i}\neq 0}(B_{n}\cap C_{n})+P_{\theta_{i}}(C^{c}_{n}).$ (49)
Observe that, for any $\theta_{i}\neq 0$ ,
$P_{\theta_{i}\neq 0}(B_{n}\cap C_{n})$ $=P_{\theta_{i}\neq 0}\left(g(X_{i},\tau,\eta,\delta)>\frac{1}{2},|X_{i}|>\sqrt{\rho\log(\frac{1}{\tau})}\right)$
$\leq P_{\theta_{i}\neq 0}\left(g_{1}(X_{i},\tau,\eta,\delta)>\frac{1}{4},|X_{i}|>\sqrt{\rho\log(\frac{1}{\tau})}\right)$
$+P_{\theta_{i}\neq 0}\left(g_{2}(X_{i},\tau,\eta,\delta)>\frac{1}{4},|X_{i}|>\sqrt{\rho\log(\frac{1}{\tau})}\right).$ (50)
Since, for given $\tau,\eta,\delta$, the function $g_{1}(x,\tau,\eta,\delta)$ steadily decreases in $x^{2}$, for any $|x|>\sqrt{\rho\log(\frac{1}{\tau})}$, we have
$g_{1}(x,\tau,\eta,\delta)\geq\frac{{C}_{1}}{\rho}\left[\log\left(\frac{1}{\tau}\right)\left\{1-\tau^{\frac{\rho}{2(1+t_{0})}}\right\}\right]^{-1}\rightarrow 0\textrm{ as }\tau\rightarrow 0.$
Since $\tau\rightarrow 0$ as $n\rightarrow\infty$, we have, $g_{1}(X_{i},\tau,\eta,\delta)\leq\frac{1}{4}$ with probability $1$ over the set $C_{n}$ for all sufficiently large $n$, and this would be true for any fixed $\eta\in(0,1)$ and $\delta\in(0,1)$, and any $\theta_{i}\neq 0$. Therefore,
$P_{\theta_{i}\neq 0}\left(g_{1}(X_{i},\tau,\eta,\delta)>\frac{1}{4}\mid|X_{i}|>\sqrt{\rho\log(\frac{1}{\tau})}\right)=0,\textrm{ for all sufficiently large }n.$ (51)
Again, we observe that, for given $\tau,\eta,\delta$, the function $g_{2}(x,\tau,\eta,\delta)$ steadily decreases in $x^{2}$. Therefore, using a similar argument, it can be shown easily that,
$P_{\theta_{i}\neq 0}\left(g_{2}(X_{i},\tau,\eta,\delta)>\frac{1}{4}\mid|X_{i}|>\sqrt{\rho\log(\frac{1}{\tau})}\right)=0,\textrm{ for all sufficiently large }n.$ (52)
Combining (5), (51) and (52), it therefore follows that
$P_{\theta_{i}\neq 0}(B_{n}\cap C_{n})=0,\textrm{ for all sufficiently large }n.$ (53)
(5) and (53) together imply that, for all sufficiently large $n$, we have
$t_{2i}$ $\leq P_{\theta_{i}\neq 0}\left(|X_{i}|\leq\sqrt{\rho\log(\frac{1}{\tau})}\right)$
$=\Phi(T_{n}-\theta_{i})-\Phi(-T_{n}-\theta_{i}),$ (54)
where $T_{n}=\sqrt{\rho\log(\frac{1}{\tau})}$. Next, we divide our calculations depending on whether $\theta_{i}$ is positive or not.
Case: Let us assume $\theta_{i}>0$, that is, $\theta_{i}>\sqrt{2\log\left(\frac{n}{q_{n}}\right)}+b$. Then, using the lower bound on $\theta_{i}$ coupled with (5), the type II error rate $t_{2i}$ can be bounded above as
$t_{2i}$ $\leq\Phi(T_{n}-\theta_{i})$
$\leq\Phi\left(\sqrt{\rho\log(\frac{1}{\tau})}-\sqrt{2\log(\frac{n}{q_{n}})}-b\right),$ (55)
where the choice of $\rho$ is same as discussed before. Observe that the decision rule does not depend on how $\eta\in(0,1),\delta\in(0,1)$ and $\rho>\frac{2}{\eta(1-\delta)}$ is chosen. Hence, taking infimum over all such $\rho$’s and subsequently over all possible choices of $(\eta,\delta)\in(0,1)\times(0,1)$, and finally using the continuity of $\Phi(\cdot)$, we have
$t_{2i}$ $\leq\Phi\left(\sqrt{2(1+g(\tau))\log(\frac{1}{\tau})}-\sqrt{2\log(\frac{n}{q_{n}})}-b\right).$ (56)
Next, under the assumption $\frac{n\tau}{q_{n}}\to C\in(0,\infty)$, we obtain
$\log(\frac{1}{\tau})=\log(\frac{n}{q_{n}})+\log(\frac{1}{C})+\log(1+o(1))$ (57)
and
$g(\tau)\log(\frac{1}{\tau})=\left(\log(\frac{n}{q_{n}})\right)^{1-\delta_{1}}(1+o(1)).$ (58)
Combining all these observations, we have
$t_{2i}$ $\leq\Phi(\sqrt{a_{n}+d_{n}}-\sqrt{a_{n}}-b),$ (59)
where $a_{n}=2\log(\frac{n}{q_{n}})$ and $d_{n}=2[\log(\frac{1}{C})+\log(1+o(1))]+2\left(\log(\frac{n}{q_{n}})\right)^{1-\delta_{1}}(1+o(1))$. Now, note that
$\Phi(\sqrt{a_{n}+d_{n}}-\sqrt{a_{n}}-b)-\Phi(-b)$
$=\frac{1}{\sqrt{2\pi}}\int_{-b}^{\sqrt{a_{n}+d_{n}}-\sqrt{a_{n}}-b}\exp(-\frac{x^{2}}{2})dx$
$\leq\sqrt{a_{n}+d_{n}}-\sqrt{a_{n}}$
$=\frac{(\sqrt{a_{n}+d_{n}}-\sqrt{a_{n}})(\sqrt{a_{n}+d_{n}}+\sqrt{a_{n}})}{\sqrt{a_{n}+d_{n}}+\sqrt{a_{n}}}$
$=\frac{d_{n}}{\sqrt{a_{n}+d_{n}}+\sqrt{a_{n}}}$
$\leq\frac{d_{n}}{2\sqrt{a_{n}}}.$ (60)
Note that for $d_{n}=(\log(\frac{n}{q_{n}}))^{1-\delta_{1}}(1+o(1))$, with $\frac{1}{2}<\delta_{1}<1$, we have $d_{n}=o(\sqrt{a_{n}})$, as $n\to\infty$. Using this observation, coupled with (5), we obtain
$t_{2i}$ $\leq\Phi(\sqrt{a_{n}+d_{n}}-\sqrt{a_{n}}-b)$
$=\Phi(-b)+\left[\Phi(\sqrt{a_{n}+d_{n}}-\sqrt{a_{n}}-b)-\Phi(-b)\right]$
$=\Phi(-b)+o(1).$ (61)
Note that, this $o(1)$ is independent of $i$.
Case 2: Let us assume that $\theta_{i}<0$, that is, $\theta_{i}<-\sqrt{2\log\left(\frac{n}{q_{n}}\right)}-b$. Then, using the lower bound on $\theta_{i}$ the upper bound of $t_{2i}$ obtained in (5) modifies to
$t_{2i}$ $\leq 1-\Phi(-T_{n}-\theta_{i})$
$\leq 1-\Phi\left(-\sqrt{\rho\log(\frac{1}{\tau})}+\sqrt{2\log(\frac{n}{q_{n}})}+b\right).$ (62)
Again using arguments similar to those used in establishing (56)-(59), an upper bound to the probability of type II error in this case is obtained as
$t_{2i}$ $\leq 1-\Phi(-\sqrt{a_{n}+d_{n}}+\sqrt{a_{n}}+b),$ (63)
where $a_{n}$ and $d_{n}$ are same as before. Next, using the same arguments used to derive (5), we obtain
$\Phi(-\sqrt{a_{n}+d_{n}}+\sqrt{a_{n}}+b)=\Phi(b)+o(1),$ (64)
where $o(1)$ is independent of $i$. Now, combining (5)-(64),
$t_{2i}$ $\leq 1-\Phi(b)+o(1).$ (65)
Finally (5) and (65) imply, for each $i$ with $|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b$,
$t_{2i}$ $\leq 1-\Phi(b)+o(1).$ (66)
This also ensures the upper bound to the left-hand side of (46) is of the order of $1-\Phi(b)+o(1)$ and completes the proof of Theorem 1. ∎
Proof of Theorem 2.
Let $t^{EB}_{1i}$ and $t^{EB}_{2i}$ denote the probabilities of the type I and type II errors for the empirical Bayes decision rule corresponding to the $i^{th}$ hypothesis, respectively. Then the overall risk associated with the empirical Bayes decision rule is given by
$\mathbb{E}L(\theta,\psi)=\sum_{i:\theta_{i}=0}t^{EB}_{1i}+\sum_{i:\theta_{i}\neq 0}t^{EB}_{2i},$ (67)
Similar to Theorem 1, here too, we need to obtain an upper bound to the left hand side of (24). Hence, it suffices to prove that
$\sum_{i:\theta_{i}=0}t^{EB}_{1i}=o(q_{n})$ (68)
and
$\frac{1}{q_{n}}\sum_{i:\theta_{i}\neq 0}\sup_{|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b}t^{EB}_{2i}\leq 1-\Phi(b)+o(1).$ (69)
For any $\alpha_{n}>0$, we have
$t^{EB}_{1i}$ $=\mathbb{P}_{\theta_{i}=0}(\mathbb{E}(1-\kappa_{i}|X_{i},\widehat{\tau})>\frac{1}{2})$
$=\mathbb{P}_{\theta_{i}=0}(\mathbb{E}(1-\kappa_{i}|X_{i},\widehat{\tau})>\frac{1}{2},\widehat{\tau}>2\alpha_{n})+\mathbb{P}_{\theta_{i}=0}(\mathbb{E}(1-\kappa_{i}|X_{i},\widehat{\tau})>\frac{1}{2},\widehat{\tau}\leq 2\alpha_{n})$
$=u_{1n}+u_{2n},\textrm{ say, }$ (70)
where $u_{1n}=\mathbb{P}_{\theta_{i}=0}(\mathbb{E}(1-\kappa_{i}|X_{i},\widehat{\tau})>\frac{1}{2},\widehat{\tau}>2\alpha_{n})$ and $u_{2n}=\mathbb{P}_{\theta_{i}=0}(\mathbb{E}(1-\kappa_{i}|X_{i},\widehat{\tau})>\frac{1}{2},\widehat{\tau}\leq 2\alpha_{n})$. Next, we observe that, for any given $x\in\mathbb{R}$, $\mathbb{E}(1-\kappa|x,\tau)$ is non-decreasing in $\tau$. Hence, using (47), we have
$u_{2n}$ $=\mathbb{P}_{\theta_{i}=0}(\mathbb{E}(1-\kappa_{i}|X_{i},\widehat{\tau})>\frac{1}{2},\widehat{\tau}\leq 2\alpha_{n})$
$=\mathbb{P}_{\theta_{i}=0}(\mathbb{E}(1-\kappa_{i}|X_{i},\widehat{\tau})>\frac{1}{2}\mid\widehat{\tau}\leq 2\alpha_{n})\mathbb{P}_{\theta_{i}=0}(\widehat{\tau}\leq 2\alpha_{n})$
$\leq\mathbb{P}_{\theta_{i}=0}(\mathbb{E}(1-\kappa_{i}|X_{i},2\alpha_{n})>\frac{1}{2})$
$\leq K_{1}M\frac{2\alpha_{n}}{\sqrt{\log(\frac{1}{4\alpha^{2}_{n}})}}(1+o(1)).$ (71)
Let us define, $\widehat{\tau}_{1}=\frac{1}{n}$ and $\widehat{\tau}_{2}=\frac{1}{c_{2}n}\sum_{i=1}^{n}\mathbf{1}(|X_{i}|>\sqrt{c_{1}\log n})$, with $c_{1}\geq 2$ and $c_{2}\geq 1$. Then, $\widehat{\tau}=\max\{\widehat{\tau}_{1},\widehat{\tau}_{2}\}$. Based on this observation, and using the definition of $u_{1n}$, we obtain
$u_{1n}$ $\leq\mathbb{P}_{\theta_{i}=0}(\widehat{\tau}>2\alpha_{n})$
$\leq\mathbb{P}_{\theta_{i}=0}(\widehat{\tau}_{1}>2\alpha_{n})+\mathbb{P}_{\theta_{i}=0}(\widehat{\tau}_{2}>2\alpha_{n}).$ (72)
Let us choose $\alpha_{n}>0$ such that $\widehat{\tau}_{1}=\dfrac{1}{n}\leq 2\alpha_{n}$ for all sufficiently large $n$, so that $\mathbb{P}_{\theta_{i}=0}(\widehat{\tau}_{1}>2\alpha_{n})=0$, provided $n$ is sufficiently large. Hence, (5) implies, for all sufficiently large $n$
$u_{1n}$ $\leq\mathbb{P}_{\theta_{i}=0}(\widehat{\tau}_{2}>2\alpha_{n})$
$\leq\mathbb{P}_{\theta_{i}=0}(\widehat{\tau}_{3}>\alpha_{n})+\mathbb{P}_{\theta_{i}=0}(\widehat{\tau}_{4}>\alpha_{n}),$ (73)
where $\widehat{\tau}_{3}=\dfrac{1}{c_{2}n}\sum\limits_{i:\theta_{i}\neq 0}\mathbf{1}(|X_{i}|>\sqrt{c_{1}\log n})$ and $\widehat{\tau}_{4}=\dfrac{1}{c_{2}n}\sum\limits_{i:\theta_{i}=0}\mathbf{1}(|X_{i}|>\sqrt{c_{1}\log n})$. Observe that,
$\widehat{\tau}_{3}$ $\leq\frac{1}{c_{2}}\frac{q_{n}}{n}.$ (74)
Let us choose $\alpha_{n}>0$ such that $\alpha_{n}\geq\frac{1}{c_{2}}\frac{q_{n}}{n}$ for all sufficiently large $n$, whence $\mathbb{P}_{\theta_{i}=0}(\widehat{\tau}_{3}>\alpha_{n})=0$.
Therefore, for all sufficiently large $n$, we obtain
$u_{1n}$ $\leq\mathbb{P}_{\theta_{i}=0}(\widehat{\tau}_{4}>\alpha_{n})$
$=\mathbb{P}_{\theta_{i}=0}\left(\frac{1}{n-q_{n}}\sum_{i:\theta_{i}=0}\mathbf{1}(|X_{i}|>\sqrt{c_{1}\log n})>c_{2}\frac{n\alpha_{n}}{n-q_{n}}\right)$
$\leq\mathbb{P}_{\theta_{i}=0}\left(\frac{1}{n-q_{n}}\sum_{i:\theta_{i}=0}\mathbf{1}(|X_{i}|>\sqrt{c_{1}\log n})>{\alpha_{n}}\right)$
$=\mathbb{P}_{\theta_{i}=0}\left(\frac{1}{n-q_{n}}\sum_{i:\theta_{i}=0}\mathbf{1}(|X_{i}|>\sqrt{c_{1}\log n})>p_{n}+\epsilon_{n}\right),$ (75)
where $p_{n}=2\mathbb{P}_{\theta_{i}=0}\left(|X_{i}|>\sqrt{c_{1}\log n}\right)$ and $\epsilon_{n}=\alpha_{n}-p_{n}$.
Using Mill’s ratio, we obtain,
$p_{n}=2\mathbb{P}_{\theta_{i}=0}\left(X_{i}>\sqrt{c_{1}\log n}\right)\propto\frac{n^{-c_{1}/2}}{\sqrt{\log n}}.$
Again, by our choice of $\alpha_{n}$, we have
$\epsilon_{n}$ $\geq$ $\dfrac{1}{c_{2}}\dfrac{q_{n}}{n}\left(1-\dfrac{\widetilde{C}}{\sqrt{\log n}}\cdot\dfrac{n^{1-c_{1}/2}}{q_{n}}\right)>0,$ (76)
for all sufficiently large $n$.
Using the preceding observations and applying Hoeffding’s Inequality, we obtain
$u_{1n}$ $\leq\mathbb{P}_{\theta_{i}=0}\left(\frac{1}{n-q_{n}}\sum_{i:\theta_{i}=0}\mathbf{1}(|X_{i}|>\sqrt{c_{1}\log n})>p_{n}+\epsilon_{n}\right),$
$\leq e^{-(n-q_{n})D\left(p_{n}+\epsilon_{n}||p_{n}\right)}$
$=e^{-(n-q_{n})D\left(\alpha_{n}||p_{n}\right)},$ (77)
where $D\left(p||q\right)=p\log\left(\dfrac{p}{q}\right)+(1-p)\log\left(\dfrac{1-p}{1-q}\right)$ denotes the Kullback-Leibler divergence between two Bernoulli distributed random variables with parameters $p$ and $q$, respectively.
Now, using the facts $\alpha_{n}\geq\frac{1}{c_{2}}\frac{q_{n}}{n}$, and $p_{n}\propto\frac{n^{-c_{1}/2}}{\sqrt{\log n}}$, we obtain
$D\left(\alpha_{n}||p_{n}\right)$ $=\alpha_{n}\log\left(\dfrac{\alpha_{n}}{p_{n}}\right)+(1-\alpha_{n})\log\left(\dfrac{1-\alpha_{n}}{1-p_{n}}\right)$
$=\alpha_{n}\log\left(\dfrac{\alpha_{n}}{p_{n}}\right)-(1-\alpha_{n})\left(\dfrac{\alpha_{n}-p_{n}}{1-p_{n}}\right)(1+o(1))$
$\gtrsim\frac{q_{n}}{n}\log(n)(1+o(1)),$ (78)
whence
$u_{1n}\leq e^{-(n-q_{n})D\left(\alpha_{n}||p_{n}\right)}\lesssim e^{-(n-q_{n})\frac{q_{n}}{n}\log(n)(1+o(1))}\lesssim e^{-q_{n}\log(n)}(1+o(1)).$ (79)
Let us choose $\alpha_{n}=c_{2}\frac{q_{n}}{n}$, so that $\alpha_{n}\to 0$ as $n\rightarrow\infty$. The aforesaid choice of $\alpha_{n}$ also satisfies the condition $\alpha_{n}\geq\frac{1}{c_{2}}\frac{q_{n}}{n}$ for all sufficiently large $n$. Hence, combining (5),(5) and (79), we have
$\frac{1}{q_{n}}\sum_{i\colon\theta_{i}=0}t^{EB}_{1i}$ $\lesssim\frac{(n-q_{n})\alpha_{n}}{q_{n}}\frac{1}{\sqrt{\log(\frac{1}{\alpha_{n}})}}(1+o(1))+\frac{n}{q_{n}}e^{-q_{n}\log(n)}(1+o(1)).$ (80)
Note that the choice of $\alpha_{n}$ satisfies $\frac{n\alpha_{n}}{q_{n}}\to C_{3}$ for some $0<C_{3}<\infty$. Hence, the first term in the right-hand side of (80) goes to zero as $n\to\infty$. On the other hand, since $q_{n}$ satisfies $q_{n}\propto{\log n}^{\delta_{2}}$ for some $\delta_{2}>1$, the second term too tends to zero as $n\to\infty$. This implies (68) holds.
Now we move towards establishing (69). Recall that, by definition, for any given $x\in\mathbb{R}$, $\mathbb{E}(\kappa|x,\tau)$ is non-increasing in $\tau$. Since $\widehat{\tau}\geq\frac{1}{n}$, hence,
$t^{EB}_{2i}$ $=\mathbb{P}_{\theta_{i}\neq 0}(\mathbb{E}(\kappa_{i}|X_{i},\widehat{\tau})>\frac{1}{2})$
$\leq\mathbb{P}_{\theta_{i}\neq 0}(\mathbb{E}(\kappa_{i}|X_{i},\frac{1}{n})>\frac{1}{2})$ (81)
Hence, using arguments similar to (5)-(56) with $\tau=\frac{1}{n}$, when $\theta_{i}>0$ for $i\in S_{\boldsymbol{\theta}}$, we have,
$t^{\text{EB}}_{2i}$ $\leq\Phi(\sqrt{2(1+(\log n)^{-\delta_{1}})\log n}-\sqrt{2\log(\frac{n}{q_{n}})}-b).$ (82)
Hence, in order to obtain the desired upper bound on $t^{\text{EB}}_{2i}$, it is sufficient to show that $\sqrt{(1+(\log n)^{-\delta_{1}})\log n}-\sqrt{\log(\frac{n}{q_{n}})}\to 0$ as $n\to\infty$. Note that, for $q_{n}\propto(\log n)^{\delta_{2}}$,
$\sqrt{(1+(\log n)^{-\delta_{1}})\log n}-\sqrt{\log(\frac{n}{q_{n}})}$ $=\sqrt{(1+(\log n)^{-\delta_{1}})\log n}-\sqrt{\log n-\delta_{2}\log\log n}>0,$
for all $n$. Hence, the upper bound converging to zero is enough. Observe that
$\sqrt{(1+(\log n)^{-\delta_{1}})\log n}-\sqrt{\log(\frac{n}{q_{n}})}$ $=\sqrt{(1+(\log n)^{-\delta_{1}})\log n}-\sqrt{\log n-\delta_{2}\log\log n}$
$=\frac{(\log n)^{(1-\delta_{1})}+\delta_{2}\log\log n}{\sqrt{(1+(\log n)^{-\delta_{1}})\log n}+\sqrt{\log n-\delta_{2}\log\log n}}$
$\leq(\log n)^{(\frac{1}{2}-\delta_{1})}+\delta_{2}\frac{\log\log n}{\sqrt{\log n}}(1+o(1))$
$\to 0,\textrm{ as }n\to\infty,$ (83)
since $\frac{1}{2}<\delta_{1}<1$. Combining (82) and (5) implies the desired upper bound on $t^{\text{EB}}_{2i}$ is established for $\theta_{i}>0$ when $i\in S_{\boldsymbol{\theta}}$. The proof when $\theta_{i}<0$ holds using similar arguments. This completes the proof of Theorem 2. ∎
Proof of Theorem 3.
Similar to Theorem 2, here too, we need to obtain an upper bound of the left-hand side of (25). Hence, it suffices to prove that
$\sum_{i:\theta_{i}=0}t^{FB}_{1i}=o(q_{n})$ (84)
and
$\frac{1}{q_{n}}\sum_{i:\theta_{i}\neq 0}\sup_{|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b}t^{FB}_{2i}\leq 1-\Phi(b)+o(1),$ (85)
where $t^{FB}_{1i}$ and $t^{FB}_{2i}$ are the type I and type II errors for the full Bayes approach corresponding to $i^{th}$ hypothesis, respectively. Note that,
$E(1-\kappa_{i}|\mathbf{X})$ $=\int_{\frac{1}{n}}^{\alpha_{n}}E(1-\kappa_{i}|\mathbf{X},\tau)\pi(\tau|\mathbf{X})d\tau$
$=\int_{\frac{1}{n}}^{\alpha_{n}}E(1-\kappa_{i}|{X}_{i},\tau)\pi(\tau|\mathbf{X})d\tau\leq E(1-\kappa_{i}|{X}_{i},\alpha_{n})\cdot$ (86)
For proving the inequality above, we first use the fact that given any $X_{i}\in\mathbb{R}$, $E(1-\kappa_{i}|X_{i},\tau)$ is non-decreasing in $\tau$, and next we use (C4). Now, using Theorem 4 of Ghosh et al. (2016), we have for $a=\frac{1}{2}$, as $n\to\infty$
$E(1-\kappa_{i}|{X}_{i},\alpha_{n})\leq{K_{1}}e^{\frac{X^{2}_{i}}{2}}{\alpha_{n}}(1+o(1)).$ (87)
Here $o(1)$ depends only on $n$ such that $\lim_{n\to\infty}o(1)=0$ and is independent of $i$ and $K_{1}$ is a constant depending on $M$. Hence, we have the following :
$t^{\text{FB}}_{1i}$ $=P_{H_{0i}}(E(1-\kappa_{i}|\mathbf{X})>\frac{1}{2})$
$\leq P_{H_{0i}}\bigg(\frac{X^{2}_{i}}{2}>\log(\frac{1}{\alpha_{n}})-\log K_{1}-\log(1+o(1))\bigg)$
$=P_{H_{0i}}\bigg(|X_{i}|>\sqrt{2\log(\frac{1}{\alpha_{n}})}\bigg)(1+o(1)).$ (88)
Note that, as $\alpha_{n}<1$ for all $n\geq 1$, $2\log(\frac{1}{\alpha_{n}})>0$ for all $n\geq 1$. Next, using the fact that under $H_{0i},X_{i}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ with $1-\Phi(t)<\frac{\phi(t)}{t}$ for $t>0$, we get
$t^{\text{FB}}_{1i}$ $\leq 2\frac{\phi(\sqrt{2\log(\frac{1}{\alpha_{n}})})}{\sqrt{2\log(\frac{1}{\alpha_{n}})}}(1+o(1))=\sqrt{\frac{1}{\pi}}\frac{{\alpha}_{n}}{\sqrt{\log(\frac{1}{\alpha^{2}_{n}})}}(1+o(1)),$ (89)
where $o(1)$ depends only on $n$ such that $\lim_{n\to\infty}o(1)=0$. This implies, for sufficiently large $n$
$\frac{1}{q_{n}}\sum_{i:\theta_{i}=0}t^{\text{FB}}_{1i}$ $\lesssim\frac{n\alpha_{n}}{q_{n}\sqrt{\log(\frac{1}{\alpha_{n}})}}(1+o(1))$
$=\frac{(\log n)^{\delta_{3}}}{(\log n)^{\delta_{2}}\sqrt{\log n-\delta_{2}\log\log n}}(1+o(1))$
$=(\log n)^{\delta_{3}-\frac{1}{2}-\delta_{2}}(1+o(1))\to 0,\textrm{ as }n\to\infty,$ (90)
since $\delta_{3}<\frac{1}{2}$ and $\delta_{2}>0$. This implies (84) holds. Next, in order to establish (85) holds, it is enough to show that, for each $i=1,\cdots,q_{n}$, $\sup_{|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b}t^{\text{FB}}_{2i}\leq 1-\Phi(b)+o(1)$, where $o(1)$ is independent of $i$.
In order to provide an upper bound on the probability of type-II error induced by the decision rule (21), $t^{\text{FB}}_{2i}=P_{H_{1i}}(E(\kappa_{i}|\mathbf{X})>\frac{1}{2})$, we first note that,
$E(\kappa_{i}|\mathbf{X})$ $=\int_{\frac{1}{n}}^{\alpha_{n}}E(\kappa_{i}|\mathbf{X},\tau)\pi(\tau|\mathbf{X})d\tau$
$=\int_{\frac{1}{n}}^{\alpha_{n}}E(\kappa_{i}|{X}_{i},\tau)\pi(\tau|\mathbf{X})d\tau\leq E(\kappa_{i}|{X}_{i},\frac{1}{n}),$ (91)
where the inequality in the last line follows due to the fact that given any $X_{i}\in\mathbb{R}$, $E(\kappa_{i}|X_{i},\tau)$ is non-increasing in $\tau$. Hence, following the arguments of (82)-(5) with $\tau=\frac{1}{n}$, when $\theta_{i}>0$ for $i=1,\cdots,q_{n}$, we have,
$\sup_{|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b}t^{\text{FB}}_{2i}$ $\leq 1-\Phi(b)+o(1),$ (92)
where the $o(1)$ term is independent of any $i$. This completes the proof of Theorem 3. ∎
Proof of Theorem 4.
The lower bound to the left hand side of (26) follows from Theorem 1 of Abraham, Castillo, and Roquain (2024). Hence, it is enough to establish that the upper bound to the left hand side of (26) is of the order $1-\Phi(b)+o(1)$.
Using the definition of $FDR(\theta,\psi)$, we have for any $\lambda\in(0,1)$,
$FDR(\theta,\psi)$ $=\mathbb{E}_{\theta}\bigg[\frac{\sum_{i:\theta_{i}=0}\psi_{i}}{\max(\sum_{i=1}^{n}\psi_{i},1)}\bigg]$
$=\mathbb{E}_{\theta}\bigg[\frac{\sum_{i:\theta_{i}=0}\psi_{i}}{\max(\sum_{i=1}^{n}\psi_{i},1)}\mathbf{1}_{\sum_{i=1}^{n}\psi_{i}>\lambda q_{n}}\bigg]+\mathbb{E}_{\theta}\bigg[\frac{\sum_{i:\theta_{i}=0}\psi_{i}}{\max(\sum_{i=1}^{n}\psi_{i},1)}\mathbf{1}_{\sum_{i=1}^{n}\psi_{i}\leq\lambda q_{n}}\bigg]$
$=u_{3n}+u_{4n},\textrm{ say}.$ (93)
Observe that, for the set $\mathbf{1}_{\sum_{i=1}^{n}\psi_{i}>\lambda q_{n}}$, $\max(\sum_{i=1}^{n}\psi_{i},1)>\lambda q_{n}$. Hence, $u_{3n}$ can be bounded as
$u_{3n}$ $\leq\frac{\sum_{i:\theta_{i}=0}t_{1i}}{\lambda q_{n}}=\frac{(n-q_{n})t_{1}}{\lambda q_{n}}$
$\leq\frac{M}{\lambda}\frac{n\tau}{q_{n}}\frac{(\tau^{2})^{a-\frac{1}{2}}}{\sqrt{\log(\frac{1}{\tau^{2}})}}(1+o(1))=o(1),\textrm{ as }n\to\infty.$ (94)
The equality in the second line holds since by definition the type I error for $i^{th}$ hypothesis $t_{1i}$ is the same for all $i$, and is denoted as $t_{1}$. The second inequality follows due to the use of (47) and the assumption on $L(\cdot)$. Now, note that
$\{\sum_{i=1}^{n}\psi_{i}\leq\lambda q_{n}\}\subseteq\{\sum_{i:\theta_{i}\neq 0}\psi_{i}\leq\lambda q_{n}\}.$ (95)
This implies
$u_{4n}$ $\leq\mathbb{P}_{\theta}(\sum_{i:\theta_{i}\neq 0}\psi_{i}\leq\lambda q_{n}).$ (96)
Since $\psi_{i}$’s are independent Bernoulli random variables, we use Hoeffding’s inequality. Choose $0<\lambda<\Phi(b)$. As a result, for all sufficiently large $n$, the constant $t$ defined in Hoeffding’s inequality is obtained as
$t$ $=q_{n}-\sum_{i:\theta_{i}\neq 0}t_{2i}-\lambda q_{n}$
$=(1-\lambda)q_{n}-\sum_{i:\theta_{i}\neq 0}t_{2i}$
$\geq[\Phi(b)-\lambda+o(1)]q_{n}.$ (97)
Note that the choice of $\lambda$ implies $t>0$ for sufficiently large $n$. In order to obtain (5), we use a uniform upper bound to all those $t_{2i}$ for which $\theta_{i}\neq 0$ as derived in (66). Using (5), we obtain, for all $|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b$, $i\in S_{\boldsymbol{\theta}}$
$\mathbb{P}_{\theta}(\sum_{i:\theta_{i}\neq 0}\psi_{i}\leq\lambda q_{n})$ $\leq\exp(-\frac{2t^{2}}{q_{n}})$
$\leq\exp(-2[\Phi(b)-\lambda+o(1)]^{2}q_{n}).$ (98)
Next, combining (96)-(5), for sufficiently large $n$
$\sup_{\theta\in\Theta_{b}}u_{4n}=o(1),\textrm{ as }n\to\infty.$ (99)
Now, (5), (5) and (99) combine to prove that
$\sup_{\theta\in\Theta_{b}}FDR(\theta,\psi)=o(1),\textrm{ as }n\to\infty.$ (100)
On the other hand, for $FNR(\theta,\psi)$, from definition, we obtain
$FNR(\theta,\psi)$ $=\frac{\mathbb{E}_{\theta}(\sum_{i:\theta_{i}\neq 0}(1-\psi_{i}))}{q_{n}}$
$=\frac{1}{q_{n}}\sum_{i:\theta_{i}\neq 0}t_{2i}.$ (101)
Recall that, in (66), we have already derived that, for each $|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b$, $i\in S_{\boldsymbol{\theta}}$,
$t_{2i}$ $\leq 1-\Phi(b)+o(1),$ (102)
where the $o(1)$ term is independent of any $i$. This, in combination with (5), implies that, for all sufficiently large $n$,
$\sup_{\theta\in\Theta_{b}}FNR(\theta,\psi)$ $\leq 1-\Phi(b)+o(1).$ (103)
Finally, (100) and (103) together imply the upper bound to the left-hand side of (26) is of the order $1-\Phi(b)+o(1)$ and completes the proof. ∎
Proof of Theorem 5.
Here we follow the basic architecture of the proof of Theorem 4. Hence, we only need to show that the upper bound to the left-hand side of (27) is of the order $1-\Phi(b)+o(1)$.
Next, using the argument of (5), we have
$u_{3n}$ $\leq\frac{\sum_{i:\theta_{i}=0}t^{EB}_{1i}}{\lambda q_{n}}$
$\lesssim\frac{n\alpha_{n}}{q_{n}}\frac{(\alpha_{n})^{2a-1}}{\sqrt{\log(\frac{1}{\alpha_{n}})}}(1+o(1))+\frac{n}{q_{n}}e^{-\frac{q_{n}}{c_{2}}(1+o(1))}.$ (104)
Here inequality in the second line holds due to (80). Note that the choice of $\alpha_{n}$ satisfies $\frac{n\alpha_{n}}{q_{n}}\to C_{3}$ for some $0<C_{3}<\infty$. Hence, the first term to the right hand side of (5) goes to zero as $n\to\infty$. On the other hand, since $q_{n}$ satisfies $q_{n}\propto\left({\log n}\right)^{\delta_{2}}$ for some $\delta_{2}>1$, the second term too tends to zero as $n\to\infty$. This implies
$u_{3n}=o(1),\textrm{ as }n\to\infty.$ (105)
Again employing (95) and (96), we obtain for sufficiently large $n$
$t$ $=q_{n}-\sum_{i:\theta_{i}\neq 0}t^{EB}_{2i}-\lambda q_{n}$
$\geq[\Phi(b)-\lambda+o(1)]q_{n},$ (106)
where the upper bound on $t^{EB}_{2i}$ as obtained in (82) is used. This lower bound with Hoeffding’s inequality confirms, for sufficiently large $n$
$\sup_{\theta\in\Theta_{b}}u_{4n}=o(1),\textrm{ as }n\to\infty.$ (107)
Combining (105) and (107) ensures
$\sup_{\theta\in\Theta_{b}}FDR(\theta,\psi)=o(1),\textrm{ as }n\to\infty.$ (108)
For $FNR(\theta,\psi)$, from definition, we obtain
$FNR(\theta,\psi)$ $=\frac{\mathbb{E}_{\theta}(\sum_{i:\theta_{i}\neq 0}(1-\psi_{i}))}{q_{n}}$
$=\frac{1}{q_{n}}\sum_{i:\theta_{i}\neq 0}t^{EB}_{2i}.$ (109)
Recall that, in (82), we have already derived that, for each $|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b$, $i\in S_{\boldsymbol{\theta}}$,
$t^{EB}_{2i}$ $\leq 1-\Phi(b)+o(1),$
where the $o(1)$ term is independent of any $i$. This combining with (5) implies for all sufficiently large $n$
$\sup_{\theta\in\Theta_{b}}FNR(\theta,\psi)$ $\leq 1-\Phi(b)+o(1).$ (110)
Finally, (108) and (110) imply the upper bound to the left hand side of (27) is of the order $1-\Phi(b)+o(1)$ and completes the proof. ∎
Proof of Theorem 6.
Here, we also follow the basic architecture of the proof of Theorem 4. Hence, in this case, too, it is sufficient to show that the upper bound to the left-hand side of (28) is of the order $1-\Phi(b)+o(1)$.
Note that, using the argument of (5) and (5), we have
$u_{3n}$ $\leq\frac{\sum_{i:\theta_{i}=0}t^{FB}_{1i}}{\lambda q_{n}}$
$\lesssim\frac{n\alpha_{n}}{q_{n}\sqrt{\log(\frac{1}{\alpha_{n}})}}(1+o(1))$
$=o(1),\textrm{ as }n\to\infty.$ (111)
Again employing arguments of (95) and (96), we obtain for sufficiently large $n$
$t$ $=q_{n}-\sum_{i:\theta_{i}\neq 0}t^{FB}_{2i}-\lambda q_{n}$
$\geq[\Phi(b)-\lambda+o(1)]q_{n},$ (112)
where the upper bound on $t^{FB}_{2i}$ as obtained in Theorem 3 is used. This lower bound with Hoeffding’s inequality confirms, for sufficiently large $n$
$\sup_{\theta\in\Theta_{b}}u_{4n}=o(1),\textrm{ as }n\to\infty.$ (113)
Combining (5) and (113) ensures
$\sup_{\theta\in\Theta_{b}}FDR(\theta,\psi)=o(1),\textrm{ as }n\to\infty.$ (114)
For $FNR(\theta,\psi)$, from definition, we obtain
$FNR(\theta,\psi)$ $=\frac{\mathbb{E}_{\theta}(\sum_{i:\theta_{i}\neq 0}(1-\psi_{i}))}{q_{n}}$
$=\frac{1}{q_{n}}\sum_{i:\theta_{i}\neq 0}t^{FB}_{2i}.$ (115)
Recall that, in Theorem 3, we have already derived that, for each $|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b$, $i\in S_{\boldsymbol{\theta}}$,
$t^{FB}_{2i}$ $\leq 1-\Phi(b)+o(1),$ (116)
where the $o(1)$ term is independent of any $i$. This combining with (5) implies for all sufficiently large $n$
$\sup_{\theta\in\Theta_{b}}FNR(\theta,\psi)$ $\leq 1-\Phi(b)+o(1).$ (117)
Finally, (114) and (117) imply the upper bound to the left hand side of (28) is of the order $1-\Phi(b)+o(1)$ and completes the proof. ∎