### v4

# Sharp Asymptotic Minimaxity for Multiple Testing Using One-Group Shrinkage Priors

Abstract
This paper investigates asymptotic minimaxity properties of Bayesian multiple testing rules in the sparse Gaussian sequence model using a broad class of global-local scale mixtures of normals as priors for the means. Minimaxity is studied under standard misclassification loss and the composite loss given by the sum of the false discovery proportion (FDP) and false non-discovery proportion (FNP). When the sparsity level is known, we show that by suitably choosing the global shrinkage parameter based on the sparsity level, our proposed testing rule achieves the exact minimax risk asymptotically for both losses under the “beta-min” separation condition. When the sparsity level is unknown, both empirical Bayes and fully Bayesian adaptations of the same rule are shown to achieve exact minimax risk asymptotically under suitable assumptions on sparsity. Our results reveal that minimaxity is attained for “horseshoe-type” priors that are broad enough to include the horseshoe, Strawderman–Berger, standard double Pareto, and certain inverse-gamma priors, among others. For non-“horseshoe-type” priors, minimaxity fails to hold for either loss function. To the best of our knowledge, these are the first results of their kind for multiple hypothesis testing based on global–local shrinkage priors.
Keywords— Multiple testing, False Discovery Rate, Global-local priors, Benjamini-Hochberg Method, Minimaxity, Classification Loss, Beta-min Conditions.
1 Introduction
Multiple hypothesis testing has emerged as a critical focal point in statistics, particularly in high-dimensional data analysis. Its relevance stretches across diverse scientific disciplines, including genomics, bioinformatics, medicine, economics, and finance. For example, in microarray experiments, researchers conduct thousands of tests simultaneously to pinpoint differentially expressed genes—those whose expression levels correlate with specific biological traits of interest. This highlights just one canonical example in which analyzing sparse, high-dimensional data is essential, with the primary objective of uncovering significant signals amidst a vast pool of noise. The ongoing exploration of multiple hypothesis testing in the literature not only advances statistical methods but also enhances our ability to extract meaningful insights from complex datasets, underscoring its importance in today’s scientific landscape.
Over the past three decades, numerous multiple testing procedures have been proposed in the literature, primarily aimed at controlling an overall measure of type I error at a predetermined level $\alpha\in(0,1)$. Historically, frequentist multiple testing procedures sought to control the family-wise error rate (FWER), defined as the probability of making at least one false rejection. The Bonferroni correction and its refinements ensure FWER control, but are often overly conservative in high-dimensional settings, resulting in very low statistical power (i.e., high type II error). This motivated the search for more flexible error criteria. The seminal work of Benjamini and Hochberg (1995) introduced the False Discovery Rate (FDR), defined as the expected proportion of false positives among all rejections, as a pragmatic (and less conservative) alternative to the family-wise error rate. The procedure for controlling the FDR, invented by Benjamini and Hochberg (1995), is henceforth referred to as the BH procedure. This work was a major impetus for modern research on multiple testing, and a vast body of work soon emerged in both the frequentist and Bayesian frameworks.
While FDR control provides guarantees on type I errors, it does not directly quantify the overall decision-theoretic performance of a procedure, especially in regimes where signal detection is intrinsically difficult. This has motivated the study of risk criteria that combine some overall type I and type II measures. Prominent examples include the expected misclassification loss (misclassification rate) and composite risk measures such as the expected weighted average of the False Discovery Rate (FDR) and False Non-discovery Rate (FNR). These criteria aim to quantify a more balanced trade-off between false positives and false negatives than approaches that merely control FDR; see, for example, Genovese and Wasserman (2009), Sun and Cai (2007), Sun and Cai (2009), and Bogdan et al. (2011). The overarching goal in this line of work is to identify procedures that achieve favorable global error trade-offs rather than focusing exclusively on one-sided error control.
Parallel to the above lines of development, there has been considerable interest among researchers in the last decade in studying and evaluating hypothesis testing procedures through the prism of minimax analysis for different loss functions. While minimax testing has been extensively studied in classical and nonparametric settings, recent work has extended this framework to high-dimensional multiple testing problems (Arias-Castro and Chen, 2017; Belitser and Nurushev, 2020; Rabinovich et al., 2020; Fromont et al., 2016), raising fundamental questions about optimal detectability under sparsity. Ours is a modest attempt to address some relevant and unresolved questions (to be described in detail later) in this context.
In minimax analysis, it is essential to consider a sufficiently large parameter space and a broad class of decision rules so that the resulting benchmarks have wide applicability. In hypothesis testing, however, trivial procedures that always reject or never reject yield worst-case risk equal to one under both Hamming loss and FDP++FNP loss. Consequently, to obtain meaningful (nontrivial) minimax risks, one must impose conditions that ensure separation between the null and alternative hypotheses so that signals can be detected with positive probability.
In the context of sparse Gaussian sequence models (1), Arias-Castro and Chen (2017) introduced the beta-min condition, requiring non-null effects to exceed a minimum magnitude. Under the additional assumption of equal signal strengths, they established a sharp dichotomy: below a universal logarithmic threshold, no thresholding rule—including BH-type methods—can achieve nontrivial minimax performance, whereas above this threshold, the BH procedure with a suitably tuned level attains vanishing minimax risk. This highlights the fundamental role of separation conditions in enabling reliable detection. Subsequently, Rabinovich et al. (2020) derived non-asymptotic bounds for the combined risk $\mathrm{FDR}+\mathrm{FNR}$ in generalized Gaussian sequence models under additional signal strength assumptions, and showed that BH-type procedures can attain the corresponding lower bounds. However, their signal strength conditions are stronger than the beta-min regime of Arias-Castro and Chen (2017), leading to asymptotically vanishing risk rather than sharp boundary behavior.
Building on these insights, Abraham, Castillo, and Roquain (2024) fully characterized the phase transition for sparse multiple testing. Under a beta-min assumption and across a wide range of sparsity regimes, they derived explicit asymptotic formulas for the minimax risk under both a composite error criterion (FDP+FNP) and the classification (Hamming) loss, delineating a clean boundary between the impossible and possible regimes of reliable testing. Crucially, they also proved that two popular procedures—BH with a vanishing rejection level and empirical-Bayes $\ell$-value rules (defined in (12)) under spike–and–slab priors—attain the exact minimax risk adaptively, without prior knowledge of the sparsity. Their results extend beyond equal-effect settings, covering heterogeneous signal strengths and more general error models, underscoring the robustness of the boundary phenomenon.
We now turn to Bayesian approaches to multiple testing, which provide a natural decision-theoretic framework through hierarchical modeling and posterior risk minimization. Within this framework, a particularly natural formulation of simultaneous hypothesis testing is based on two-group mixture models, such as spike-and-slab priors (Mitchell and Beauchamp, 1988). In the Gaussian sequence model, where the goal is to identify which mean parameters are exactly zero, and which are non-zero, a spike-and-slab prior models each mean as a mixture of a point mass at zero (the spike) and a continuous, typically heavy-tailed distribution (the slab), with the mixing proportion representing the prior probability of a non-null effect. Under this formulation, the Bayes-optimal decision rule for additive 011 loss rejects null hypotheses whose posterior inclusion probabilities are below $1/2$.
This framework has been widely studied and applied in large-scale testing problems; see, for example, Efron (2004), Müller et al. (2004), Scott and Berger (2006), Müller, Parmigiani, and Rice (2006), Efron (2007), and Bogdan et al. (2011), among others. These works establish both the practical effectiveness and the strong theoretical properties of Bayesian two-group models for large-scale multiple testing, including optimality results under sparsity and several risk criteria. More broadly, theoretical properties of two-group mixture formulations for high-dimensional inference have been studied extensively in related contexts; see, for example, Johnstone and Silverman (2004), Cai and Jin (2010), Castillo and van der Vaart (2012), Neuvial and Roquain (2012), Martin and Walker (2014), and Castillo, Schmidt-Heiber, and van der Vaart (2015), only to name a few.
Although intuitive and theoretically appealing on many counts, two-group models often face daunting computational challenges in high-dimensional problems and may be less suitable when most parameters are close to, but not exactly, zero, and posterior sampling or marginal likelihood optimization must be carried out at scale. This has motivated the development of continuous one-group shrinkage priors, which are computationally more tractable and offer greater flexibility. In this formulation, each parameter is modeled as a scale mixture of normals, with a global shrinkage parameter controlling overall sparsity and local shrinkage parameters allowing signal-specific adaptivity. Prominent examples include the inverse Gamma priors (Tipping, 2001), the horseshoe prior (Carvalho, Polson and Scott, 2009, 2010), the three-parameter beta normal mixtures (Armagan, Clyde and Dunson, 2011), the generalized double Pareto prior (Armagan, Dunson and Lee, 2013), the Dirichlet–Laplace prior (Bhattacharya et al., 2015), and the horseshoe++ prior (Bhadra et al., 2017). These priors share the property of tail robustness: they strongly shrink small signals toward zero while leaving large effects essentially unshrunk, enabling estimation risks that are within logarithmic factors of the minimax benchmark without prior knowledge of sparsity. See, for instance, van der Pas, Kleijn and van der Vaart (2014), van der Pas, Salomond, and Schmidt-Heiber (2016), van der Pas, Szabo and van der Vaart (2017), Ghosh and Chakrabarti (2017), Bhadra et al. (2017), and Paul and Chakrabarti (2025), among others.
In the context of multiple hypothesis testing for the Gaussian sequence model (1), Carvalho, Polson and Scott (2009) proposed an intuitive decision rule based on posterior shrinkage factors arising from the horseshoe prior, which behave analogously to posterior inclusion probabilities in spike-and-slab models. Building on this idea, Datta and Ghosh (2013) and Ghosh et al. (2016) studied such rules from a formal decision-theoretic perspective under misclassification (Hamming) loss. They showed that decision rules based on a broad class of heavy-tailed one-group shrinkage priors, including the horseshoe, can achieve asymptotically near-optimal Bayes risk (in the sense of Bogdan et al., 2011) under sparsity when the data are generated from a two-group model. Subsequent work by Ghosh and Chakrabarti (2017) and Paul and Chakrabarti (2025) refined these results and established that testing procedures based on a broad class of so-called horseshoe-type priors (defined later) in fact attain exact asymptotic Bayes optimality in terms of type I and type II error trade-offs, paralleling the sharp detection boundaries known for spike-and-slab priors. Collectively, these results demonstrate that carefully chosen one-group shrinkage priors can serve as theoretically optimal, adaptive, and computationally attractive alternatives to discrete mixture models for large-scale multiple testing. Since the two terms misclassification loss and Hamming loss are synonymous, we will use only the term misclassification loss throughout the remaining part of the paper.
To the best of our knowledge, the first investigation of minimax risk properties for multiple testing procedures based on one-group priors was undertaken by Salomond (2017). That work studied thresholding rules similar to those described above, using a broad class of Gaussian scale-mixture priors inspired by the analysis of van der Pas, Salomond, and Schmidt-Heiber (2016). Under separation conditions analogous to beta-min assumptions on the non-zero means, Salomond (2017) derived upper bounds on the maximum risk measured by the sum of the false discovery rate (FDR) and the false non-discovery rate (FNR). These results imply that when the non-null signals are sufficiently strong, the corresponding risk tends to zero asymptotically, indicating consistent detection.
However, that work did not establish exact asymptotic minimaxity results for the proposed procedures, and the sharp minimax risk expressions for this problem were obtained only later by Abraham, Castillo, and Roquain (2024). Consequently, it remained unclear whether any multiple testing procedure based on continuous one-group shrinkage priors could attain the exact minimax risk under the beta-min conditions. Moreover, Salomond (2017) did not study minimaxity under misclassification loss, and while an empirical Bayes approach was considered to adapt to unknown sparsity, fully Bayesian procedures were not analyzed. Bridging this gap between sharp minimax theory and Bayesian one-group testing procedures is the central objective of the present work.
Motivated by the sharp minimax results of Abraham, Castillo, and Roquain (2024), the unresolved questions in Salomond (2017), and the strong Bayes optimality results for one-group priors in Datta and Ghosh (2013), Ghosh et al. (2016), Ghosh and Chakrabarti (2017), and Paul and Chakrabarti (2025), we study multiple hypothesis testing using one-group shrinkage priors from the perspective of minimax optimality.
We assume that we have an observation vector, say, $\mathbf{X}=(X_{1},\dots,X_{n})\in\mathbb{R}^{n}$, from the Gaussian sequence model, such that each $X_{i}$ can be expressed as
$X_{i}=\theta_{i}+\epsilon_{i},\textrm{ for }i=1,\dots,n.$ (1)
In (1) above, the parameters $\theta_{1},\dots,\theta_{n}$ are unknown mean parameters, and the unobserved residuals $\epsilon_{i}$’s are assumed to be independent $N(0,\sigma^{2})$ random variables. For theoretical developments, the residual variance $\sigma^{2}$ is typically assumed to be known. Thus, without any loss of generality, one may assume $\sigma^{2}$ to be $1$. For the above normal means model, we are interested in testing the following hypotheses:
$\textrm{H}_{0i}:\theta_{i}=0\textrm{ vs. }\textrm{H}_{1i}:\theta_{i}\neq 0,\textrm{ simultaneously for }i=1,\ldots,n.$ (2)
Following Abraham, Castillo, and Roquain (2024), we work under a sparse asymptotic regime in which the parameter vector $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{n})$ belongs to the set
[…]
consisting of vectors that have at most $q_{n}$ nonzero coordinates. Throughout the paper, we consider the sparse asymptotic framework where
$q_{n}\rightarrow\infty,\ \textrm{and}\ \dfrac{q_{n}}{n}\rightarrow 0\textrm{ as }n\rightarrow\infty.$
We focus on the same broad class of priors previously considered in Ghosh et al. (2016) and Ghosh and Chakrabarti (2017) when each unknown $\theta_{i}$ is modeled as a scale mixture of normals, as,
$\theta_{i}|\lambda_{i},\tau\stackrel{{\scriptstyle ind}}{{\sim}}N(0,\lambda^{2}_{i}\sigma^{2}\tau^{2}),\ \lambda^{2}_{i}\stackrel{{\scriptstyle ind}}{{\sim}}\pi_{1}(\lambda^{2}_{i})=K(\lambda^{2}_{i})^{-a-1}L(\lambda^{2}_{i}),\ \tau\sim\pi_{2}(\tau),$ (3)
for some choices of $\pi_{2}$. In (3) above, $K>0$ is the constant of proportionality, $a$ is a positive real number, and $L\colon(0,\infty)\mapsto(0,\infty)$ is a measurable, continuous, and slowly varying function. In this formulation, $\lambda_{i}$’s are called the local shrinkage parameters (calibrating the shrinkage at the individual level) and $\tau$ is called the global shrinkage parameter (inducing the overall sparsity of the parameter vector). As discussed by Polson and Scott (2010), heavy-tailed priors on the local scales and substantial prior mass for $\tau$ near zero induce marginal priors on the $\theta_{i}$’s that resemble two-group mixtures, with strong concentration near zero and heavy tails for large signals.
Considering the composite loss $\textrm{FDP}+\textrm{FNP}$ and the hamming loss, our aim in this paper is to investigate whether the multiple testing rules (formally stated in (20), (21), and (23)), based on the above class of one-group shrinkage priors, asymptotically achieve the minimax risk under appropriate conditions. One of our theoretical results (Proposition 1) indicates that one-group shrinkage priors (3) with $a>0.5$ cannot achieve the desired minimax risk for either loss function. Therefore, we confine our attention to the subclass of horseshoe-type priors, where $a=0.5$. This subclass of horseshoe-type priors is quite extensive, encompassing various popular prior distributions such as the horseshoe, Strawderman-Berger, standard double Pareto, and inverse-gamma priors (with shape parameter $0.5$), among others. When the sparsity level is known, we show (Theorems 1 and 4) that testing rules based on horseshoe-type priors attain the exact minimax risk under both losses, provided the global shrinkage parameter is chosen appropriately as a function of the proportion of non-zero means. When the sparsity level is unknown, we further establish that both empirical Bayes and fully Bayesian adaptations of the same testing rules (Theorems 2, 3, 5, and 6) continue to achieve exact minimax risk asymptotically as the number of tests grows to infinity. To the best of our knowledge, these findings are the first of their kind in the literature, demonstrating that properly chosen global-local priors can guarantee exact attainment of minimax risk optimality in the context of a simultaneous hypothesis testing problem.
The remainder of the paper is organized as follows. Section 2 reviews existing minimax results in Gaussian sequence models. Section 3 presents our main theoretical contributions, with Subsection 3 focusing on classification loss and Subsection 3.2 on composite loss (FDR+FNR). Section 4 provides concluding remarks and future directions. Proofs are deferred to Section 5.
1.1 Notations
For any two sequences $\{a_{n}\}$ and $\{b_{n}\}$ with $\{b_{n}\}\neq 0$, we write $a_{n}\asymp b_{n}$ to denote that there exist two constants $c_{1}$ and $c_{2}$ such that $0<c_{1}\leq a_{n}/b_{n}\leq c_{2}<\infty$ for sufficiently large $n$. Similarly, $a_{n}\lesssim b_{n}$ denotes that for sufficiently large $n$, there exists a constant $c>0$ such that $a_{n}\leq cb_{n}$. Similarly, $a_{n}\propto b_{n}$ implies there exists some constant $C>0$ such that $a_{n}=Cb_{n}$ for all $n$. Finally, $a_{n}=o(b_{n})$ denotes $\lim_{n\to\infty}\frac{a_{n}}{b_{n}}=0$. Let $\Phi(\cdot)$ and $\phi(\cdot)$ denote the cumulative distribution function (CDF) and the probability density function (PDF) of the standard normal distribution. Also, $\bar{\Phi}(x)=1-\Phi(x)$.
2 Existing Results on Relevant Minimax Risk Theory
Before reviewing the relevant literature on minimax theory in the context of multiple hypothesis testing, we first introduce some notations and concepts that will be used throughout the rest of this paper. We consider the multiple testing problem (2) under the Gaussian sequence model (1).
A multiple testing rule is a measurable function $\boldsymbol{\psi}=(\psi_{1},\ldots,\psi_{n})$ taking values in $\{0,1\}^{n}$, where $\psi_{i}=1$ indicates rejection of $H_{0i}$. Given $\boldsymbol{\theta}\in\mathbb{R}^{n}$ and a testing rule $\boldsymbol{\psi}=(\psi_{1},\ldots,\psi_{n})$, define the numbers of false discoveries and false non-discoveries as
$V(\boldsymbol{\theta},\boldsymbol{\psi})=\sum_{i:\,\theta_{i}=0}\psi_{i},\qquad T(\boldsymbol{\theta},\boldsymbol{\psi})=\sum_{i:\,\theta_{i}\neq 0}(1-\psi_{i}).$
Let $R(\boldsymbol{\psi})=\sum_{i=1}^{n}\psi_{i}$ denote the total number of rejections and $S(\boldsymbol{\theta})=\sum_{i=1}^{n}\mathbf{1}\{\theta_{i}\neq 0\}$ be the number of true signals. Then, the false discovery proportion (FDP) and the false non-discovery proportion (FNP) are defined as
$\mathrm{FDP}(\boldsymbol{\theta},\boldsymbol{\psi})=\frac{V(\boldsymbol{\theta},\boldsymbol{\psi})}{\max\{R(\boldsymbol{\psi}),1\}},\qquad\mathrm{FNP}(\boldsymbol{\theta},\boldsymbol{\psi})=\frac{T(\boldsymbol{\theta},\boldsymbol{\psi})}{\max\{S(\boldsymbol{\theta}),1\}}.$
Given $\boldsymbol{\theta}\in\mathbb{R}^{n}$ and a testing rule $\boldsymbol{\psi}$, the false discovery rate (FDR) is defined as
$\mathrm{FDR}(\boldsymbol{\theta},\boldsymbol{\psi})$ $=\mathbb{E}_{\boldsymbol{\theta}}\!\left[\mathrm{FDP}\right].$ (4)
Given that a multiple testing rule controls the FDR at a desired level, a natural next question is whether the same rule can control a measure of type II error. A commonly used criterion is the false non-discovery rate (FNR), defined as
$\mathrm{FNR}(\boldsymbol{\theta},\boldsymbol{\psi})$ $=\mathbb{E}_{\boldsymbol{\theta}}\!\left[\mathrm{FNP}\right].$ (5)
In sparse multiple testing, sparsity alone does not guarantee reliable detection if nonzero signals are too close to the noise level. In such cases, no procedure can simultaneously control both type I and type II errors, and the minimax risk remains trivial. Therefore, sharp minimax analysis must incorporate lower bounds on signal strength in addition to sparsity, leading naturally to beta-min conditions that characterize the detectability boundary. The most recent and influential contributions on minimax risk for the Gaussian sequence model are due to Abraham, Castillo, and Roquain (2024). They introduced a beta-min type condition designed to separate the null from the alternative, defined as follows. Given $\tilde{a}\in\mathbb{R}$, let
$\boldsymbol{\Theta}(\tilde{a},q_{n})=\bigl\{\boldsymbol{\theta}\in\ell_{0}[q_{n}]:|\theta_{i}|\geq\tilde{a}\text{ for all }i\in S_{\boldsymbol{\theta}},\ |S_{\boldsymbol{\theta}}|=q_{n}\bigr\},$ (6)
where $S_{\boldsymbol{\theta}}=\{i:\theta_{i}\neq 0\}$ denotes the support of $\boldsymbol{\theta}$. Motivated by earlier works of Arias-Castro and Chen (2017) and Rabinovich et al. (2020), they chose $\tilde{a}$ to be close to the universal threshold $\sqrt{2\log(n/q_{n})}$. Specifically, they considered
$\boldsymbol{\Theta}_{b}=\boldsymbol{\Theta}(\tilde{a}_{b},q_{n}),\qquad\tilde{a}_{b}=\sqrt{2\log\!\left(\frac{n}{q_{n}}\right)}+b,$ (7)
where $b\in\mathbb{R}$ is fixed.
Under this framework, Abraham, Castillo, and Roquain (2024) established that for any $b\in\mathbb{R}$,
$\inf_{\boldsymbol{\psi}}\sup_{\boldsymbol{\theta}\in\boldsymbol{\Theta}_{b}}R(\boldsymbol{\theta},\boldsymbol{\psi})=1-\Phi(b)+o(1),$ (8)
where
$R(\boldsymbol{\theta},\boldsymbol{\psi})=\mathrm{FDR}(\boldsymbol{\theta},\boldsymbol{\psi})+\mathrm{FNR}(\boldsymbol{\theta},\boldsymbol{\psi}).$ (9)
This result remains valid when $b=b_{n}\to\infty$ or $b=b_{n}\to$-$\infty$. It follows that the oracle rule
$\psi_{i}=\mathbf{1}\!\left\{|X_{i}|\geq\sqrt{2\log(n/q_{n})}\right\},\qquad i=1,\ldots,n,$
is asymptotically minimax. Moreover, they identified a sharp boundary for nontrivial testing: when $b\to$-$\infty$, no procedure outperforms trivial rules, whereas finite $b$ yields nontrivial minimax risk.
Since the oracle rule depends on $q_{n}$, which is typically unknown, an important question is whether data-driven procedures can achieve the minimax risk in (8). Assuming $q_{n}\lesssim n^{c}$ for some $0<c<1$, Abraham, Castillo, and Roquain (2024) showed that both the BH procedure with $\alpha_{n}\to 0$ and $\ell$-value procedures under spike-and-slab priors with suitable slabs attain the minimax bound.
Specifically, under the two-group prior
$\theta_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}(1-p)\delta_{0}+pF,\ \textrm{ for }i=1,\ldots,n,$ (10)
the marginal distribution of $X_{i}$ is
$X_{i}\sim(1-p)\phi(x)+p\int_{\mathbb{R}}\phi(x-\theta)\,F(\theta)\,d\theta,$ (11)
where $\phi(\cdot)$ denotes the standard normal density. With a quasi-Cauchy slab, they defined the posterior $\ell$-value
$\ell_{i,p}(\mathbf{X})=\Pi_{p}(\theta_{i}=0\mid\mathbf{X})=\frac{(1-p)\varphi(X_{i})}{(1-p)\varphi(X_{i})+p\tilde{f}(X_{i})},$ (12)
and rejected $H_{0i}$ when $\ell_{i,p}(\mathbf{X})<t$. Since $p$ is unknown, they used the MMLE $\hat{p}$ and showed that the resulting procedure achieves minimax risk asymptotically.
Abraham, Castillo, and Roquain (2024) also extended their work to heterogeneous signals. For $\tilde{a}=(\tilde{a}_{1},\ldots,\tilde{a}_{q_{n}})$, define
$\boldsymbol{\Theta}(\tilde{a},q_{n})=\Bigl\{\boldsymbol{\theta}\in\ell_{0}[q_{n}]:|\theta_{i_{j}}|\geq\tilde{a}_{j},\ j=1,\ldots,q_{n}\Bigr\}.$ (13)
They introduced
$\Lambda_{n}(\tilde{a})=\frac{1}{q_{n}}\sum_{j=1}^{q_{n}}F_{\tilde{a}_{j}}(a_{n}^{*}),\qquad a_{n}^{*}=\bigl(\zeta\log(n/q_{n})\bigr)^{1/\zeta},\ \zeta>1,$
and proved that for symmetric location models,
$\inf_{\psi}\sup_{\theta\in\Theta(\tilde{a},q_{n})}R(\boldsymbol{\theta},\boldsymbol{\psi})=\Lambda_{n}(\tilde{a})+o(1).$ (14)
For the Gaussian sequence model, $\Lambda_{n}(\tilde{a})=q_{n}^{-1}\sum_{j=1}^{q_{n}}[1-\Phi(\tilde{a}_{j}-a_{n}^{*})]$ with $a_{n}^{*}=\sqrt{2\log(n/q_{n})}$. Both BH and $\ell$-value procedures attain this bound.
They also studied minimax risk under classification (Hamming) loss defined as
$L(\boldsymbol{\theta},\boldsymbol{\psi})=\sum_{i=1}^{n}\bigl[\mathbf{1}\{\theta_{i}=0,\psi_{i}\neq 0\}+\mathbf{1}\{\theta_{i}\neq 0,\psi_{i}=0\}\bigr],$ (15)
and showed that
$\frac{1}{q_{n}}\inf_{\psi}\sup_{\theta\in\Theta_{b}}\mathbb{E}L(\boldsymbol{\theta},\boldsymbol{\psi})=1-\Phi(b)+o(1),$ (16)
which is also achieved by BH and $\ell$-value rules.
As noted earlier, the only prior work addressing minimax risk using one-group priors in the Gaussian sequence model is Salomond (2017). Modeling
$\theta_{i}\mid\sigma_{i}^{2}\sim N(0,\sigma_{i}^{2}),\qquad\sigma_{i}^{2}\sim\pi(\sigma_{i}^{2}),$ (17)
they proposed thresholding rules based on posterior shrinkage factors and obtained upper bounds on FDR+FNR under separation conditions. However, they did not establish exact minimaxity, did not study misclassification loss, and did not derive sharp minimax boundaries. Thus, the minimax optimality of multiple testing rules based on one-group priors remained open, which we address in the next section.
3 Theoretical Results
In this section, we present the main theoretical results of the paper. As indicated earlier, our interest lies in investigating asymptotic minimaxity properties of multiple testing rules based on one-group priors when the observations arise from the Gaussian sequence model (1). We begin by formally defining the testing rules. Recall that we model the $\theta_{i}$’s through the scale-mixture specification in (3).
In the hierarchical formulation (3), the posterior mean of $\theta_{i}$ admits the representation
$\mathbb{E}(\theta_{i}\mid\mathbf{X},\tau,\sigma)=\bigl(1-\mathbb{E}(\kappa_{i}\mid X_{i},\tau,\sigma)\bigr)X_{i},$ (18)
where $\kappa_{i}=1/(1+\lambda_{i}^{2}\tau^{2})$. Consequently, $\mathbb{E}(1-\kappa_{i}\mid X_{i},\tau,\sigma)$ is the $i$th posterior shrinkage coefficient, quantifying the fraction by which the MLE $X_{i}$ is shrunk toward zero in the Bayes estimator of $\theta_{i}$.
By comparing posterior means under spike-and-slab priors and the horseshoe prior, Carvalho, Polson and Scott (2009) observed that in sparse regimes, the posterior shrinkage coefficient behaves analogously to the posterior inclusion probability $\Pi(\theta_{i}\neq 0\mid\mathbf{X})$ arising from spike-and-slab models with heavy-tailed slabs. This led them to propose the following decision rule (under the horseshoe prior) for the multiple testing problem (2) in model (1):
$\text{Reject }H_{0i}\ \text{ if }\ \mathbb{E}(1-\kappa_{i}\mid X_{i},\tau)>\frac{1}{2},\ \textrm{ for }i=1,\ldots,n.$ (19)
Their simulations yielded promising misclassification performance relative to the Bayes rule under additive 011 loss when the $\theta_{i}$’s are generated from a spike-and-slab model. As discussed earlier, this phenomenon was subsequently investigated theoretically: several authors established formal (near-)optimal asymptotic properties of related testing rules under a broad class of one-group priors of the form (3), when the data are generated from a two-group model. Motivated by these positive results (albeit not with respect to minimax benchmarks), we investigate the minimaxity properties of the following testing rule $\boldsymbol{\psi}=(\psi_{1},\ldots,\psi_{n})$, where for $i=1,\ldots,n$,
$\psi_{i}\equiv\psi_{i}(X_{i})=\mathbf{1}\Bigl\{\mathbb{E}(1-\kappa_{i}\mid X_{i},\tau)>\frac{1}{2}\Bigr\}.$ (20)
Assumption on $L(\cdot)$.
For the theoretical study of the rule (20), we assume that the slowly varying function $L(\cdot)$ in (3) satisfies the following condition.
Assumption 1. For $a\geq\frac{1}{2}$:
  • (A1)
    There exists $c_{0}>0$ such that $L(t)\geq c_{0}$ for all $t\geq t_{0}$, for some $t_{0}>0$ depending on $L$ and $c_{0}$.
  • (A2)
    There exists $M\in(0,\infty)$ such that $\sup_{t\in(0,\infty)}L(t)\leq M$.
Earlier results of Datta and Ghosh (2013), van der Pas, Kleijn and van der Vaart (2014), Ghosh et al. (2016), and Ghosh and Chakrabarti (2017) clearly indicate that the choice of the global shrinkage parameter $\tau$ plays a pivotal role in capturing the underlying sparsity and in determining asymptotic optimality of the rule (20) when $\tau$ is tuned using (known) sparsity. In practice, however, the sparsity level is typically unknown. In that case, we replace $\tau$ by an estimate $\widehat{\tau}$ and employ an empirical Bayes version of (20):
$\psi_{i}\equiv\psi^{\mathrm{EB}}_{i}(\mathbf{X})=\mathbf{1}\Bigl\{\mathbb{E}(1-\kappa_{i}\mid X_{i},\widehat{\tau})>\frac{1}{2}\Bigr\},\ \textrm{ for }i=1,\ldots,n.$ (21)
Here
$\widehat{\tau}=\max\Bigl\{\frac{1}{n},\frac{1}{c_{2}n}\sum_{i=1}^{n}\mathbf{1}\bigl(|X_{i}|>\sqrt{c_{1}\log n}\bigr)\Bigr\},$ (22)
with $c_{1}\geq 2$ and $c_{2}\geq 1$, as proposed by van der Pas, Kleijn and van der Vaart (2014).
An alternative approach under unknown sparsity is to assign an absolutely continuous prior $\pi_{2}(\tau)$ to $\tau$ (full Bayes). Motivated by Paul and Chakrabarti (2025), we consider a class of priors $\pi_{2}(\cdot)$ satisfying:
[…]
In the full Bayes procedure, the decision rule is modified as
$\psi_{i}\equiv\psi^{\mathrm{FB}}_{i}(\mathbf{X})=\mathbf{1}\Bigl\{\mathbb{E}(1-\kappa_{i}\mid\mathbf{X})>\frac{1}{2}\Bigr\},\ \textrm{ for }i=1,\ldots,n.$ (23)
where the full Bayes posterior mean of $1-\kappa_{i}$ is
$\mathbb{E}(1-\kappa_{i}\mid\mathbf{X})$ $=\int_{1/n}^{\alpha_{n}}\mathbb{E}(1-\kappa_{i}\mid\mathbf{X},\tau)\,\pi(\tau\mid\mathbf{X})\,d\tau$
$=\int_{1/n}^{\alpha_{n}}\mathbb{E}(1-\kappa_{i}\mid X_{i},\tau)\,\pi(\tau\mid\mathbf{X})\,d\tau.$
We study minimax optimality of the rules (20), (21), and (23)—depending on whether the sparsity level is known or unknown—under both the classification (Hamming) loss and the composite loss $\mathrm{FDP}+\mathrm{FNP}$. Results under the classification loss are presented in Subsection 3.1, while Subsection 3.2 contains results for the $\mathrm{FDP}+\mathrm{FNP}$ loss.
Choice of the global shrinkage level.
The global parameter $\tau$ controls the overall amount of shrinkage and therefore plays a crucial role in balancing type I and type II errors. If $\tau$ is too small, the procedure becomes overly conservative and fails to detect signals even at the beta-min boundary; if $\tau$ is too large, excessive noise coordinates escape shrinkage, leading to inflated false discoveries. Under sparse asymptotics, the scale $\tau\asymp q_{n}/n$ represents the only regime in which posterior shrinkage adapts to the effective signal proportion while still allowing detection at the critical level $\sqrt{2\log(n/q_{n})}$. This scaling has also appeared in earlier Bayes-optimality results for one-group priors and will be adopted throughout our minimax analysis.
Why $a>1/2$ is suboptimal for minimaxity.
We conclude this preliminary part with a lower bound on the type II error probability for the $i$th hypothesis, which explains why the case $a>1/2$ in (3) is not compatible with minimax optimality for the losses considered here.
Proposition 1.
Consider the Gaussian sequence model (1) and the simultaneous hypothesis testing problem (2) based on the decision rule (20) using the class of one-group shrinkage priors (3). Assume that $\tau\equiv\tau_{n}\to 0$ as $n\to\infty$ such that $\frac{n\tau}{q_{n}}\to C\in(0,\infty)$. Further assume that the slowly varying function $L(\cdot)$ in (3) satisfies Assumption 1. Then for $a>\frac{1}{2}$ in (3), for $i=1,\ldots,q_{n}$, and $b\in\mathbb{R}$,
$\sup_{|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b}t_{2i}\to 1,\quad\text{as }n\to\infty,$ (24)
where $t_{2i}$ denotes the type II error probability corresponding to the $i$th hypothesis.
Observe that the expected misclassification loss $\mathbb{E}L(\boldsymbol{\theta},\boldsymbol{\psi})$ (see (15)) depends on both the type I and type II error probabilities $t_{1i}$ and $t_{2i}$, respectively. As will be seen from the proofs of Theorems 13, the type II error is the dominating term in the asymptotic regime considered here. Hence, the fact that $t_{2i}\to 1$ uniformly over $\boldsymbol{\Theta}_{b}$ for $a>1/2$ implies that the resulting risk cannot attain the sharp minimax benchmark $1-\Phi(b)+o(1)$. A similar phenomenon occurs for the composite loss $\mathrm{FDP}+\mathrm{FNP}$: the proofs of Theorems 46 show that the minimax behavior is primarily driven by $\mathrm{FNP}$ (equivalently, by type II errors). Consequently, the conclusion of Proposition 1 rules out $a>1/2$ for attaining minimax risk under this loss as well. In what follows, we therefore focus on the case $a=1/2$ (the horseshoe-type class).
Why we do not consider the case $a<\tfrac{1}{2}$.
Proposition 1 rules out $a>\tfrac{1}{2}$ due to excessively large type II errors. This leaves open the possibility $a<\tfrac{1}{2}$, which corresponds to priors with even heavier tails than the horseshoe-type class. While such priors are attractive for estimation due to reduced shrinkage of large signals, they tend to induce insufficient shrinkage of noise coordinates. Consequently, posterior shrinkage coefficients $\mathbb{E}(1-\kappa_{i}\mid X_{i},\tau)$ remain non-negligible even for moderate values of $X_{i}$ under the null, leading to elevated type I error rates unless the global parameter $\tau$ is tuned more aggressively than $\tau\asymp q_{n}/n$. In particular, for $a<\tfrac{1}{2}$ the posterior shrinkage coefficient under the null does not decay sufficiently fast at the detection scale $\sqrt{2\log(n/q_{n})}$, so that the rejection event in (20) occurs too often when $\theta_{i}=0$. Heuristically, this yields an inflation of false discoveries (and hence of $\mathrm{FDP}$ and $\mathrm{FDR}$) that cannot be offset while keeping the tuning $\tau\asymp q_{n}/n$ required for sharp power at the beta-min boundary.
Under the beta-min regime, minimax optimality requires a precise balance between type I and type II errors at the detection boundary $\sqrt{2\log(n/q_{n})}$. Priors with $a<\tfrac{1}{2}$ place too much mass on large local scales, which heuristically shifts this balance toward false discoveries and prevents simultaneous control of both error components at the sharp minimax rate using the decision rule (20) with the tuning $\tau\asymp q_{n}/n$. Establishing formal lower bounds for this regime would require different techniques and potentially different tuning of $\tau$, which is beyond the scope of the present work.
Together, these observations suggest that exact minimaxity can only occur at the boundary case $a=\tfrac{1}{2}$, which balances tail robustness with sufficient shrinkage of null coordinates. We therefore restrict attention to this horseshoe-type subclass henceforth.
3.1 Results Based On the Classification Loss
In this subsection, we establish sharp asymptotic minimaxity under the Classification loss for all three testing rules: the fixed-$\tau$ rule, the EB rule, and the FB rule. We first consider the idealized setting in which the sparsity level is known, and the global shrinkage parameter $\tau$ is tuned accordingly, and then turn to adaptive procedures when sparsity is unknown. Motivated by Proposition 1, we restrict attention to the case $a=\tfrac{1}{2}$ in (3), corresponding to the horseshoe-type class of priors.
Theorem 1.
Consider the Gaussian sequence model (1), and the simultaneous hypothesis testing problem (2) based on the decision rule (20) using the class of one-group shrinkage priors (3). Assume that $\tau\equiv\tau_{n}\to 0$ as $n\to\infty$ such that $\frac{n\tau}{q_{n}}\to C\in(0,\infty)$. Further assume that the slowly varying function $L(\cdot)$ in (3) satisfies Assumption 1. Then for the sub-class of horseshoe-type priors (i.e. $a=\frac{1}{2}$),
$\frac{1}{q_{n}}\sup_{\boldsymbol{\theta}\in\boldsymbol{\Theta}_{b}}\mathbb{E}L(\boldsymbol{\theta},\boldsymbol{\psi})=1-\Phi(b)+o(1),$ (25)
where $L(\boldsymbol{\theta},\boldsymbol{\psi})$ is defined in (15) and $\boldsymbol{\Theta}_{b}$ is defined in (7). This result holds for any $b\in\mathbb{R}$ and also for sequences $b=b_{n}\to\infty$.
Remark 1.
Theorem 1 provides a multiple testing rule based on one-group priors that attains the sharp minimax risk under classification (Hamming) loss. To the best of our knowledge, this is the first such result for global–local priors in this setting. The only condition required on $\tau$ is that it is asymptotically of the same order as $q_{n}/n$, i.e. the proportion of non-null signals. The same scaling was used by Ghosh et al. (2016) to establish Bayes risk optimality (ABOS-type) when the data are generated from a two-group spike-and-slab model. Our proof proceeds by obtaining appropriate bounds on the type I and type II error probabilities. For the type I error, Theorem 6 of Ghosh et al. (2016) is useful; the control of the type II error, however, requires additional technical arguments.
Theorem 1 assumes that the sparsity level is known, which is often unrealistic. This motivates the study of the adaptive rules (21) and (23). The next two theorems show that minimax risk under classification loss can still be attained. Their proofs are given in the Section 5.
Theorem 2.
Consider the setup of Theorem 1, and the empirical Bayes rule (21), where $\widehat{\tau}$ is defined in (22). Assume that $q_{n}\propto(\log n)^{\delta_{2}}$ for some $\delta_{2}>1$. Then for the class horseshoe-type priors (i.e. $a=\frac{1}{2}$),
$\frac{1}{q_{n}}\sup_{\boldsymbol{\theta}\in\boldsymbol{\Theta}_{b}}\mathbb{E}L(\boldsymbol{\theta},\boldsymbol{\psi})=1-\Phi(b)+o(1),$ (26)
where $L(\boldsymbol{\theta},\boldsymbol{\psi})$ and $\boldsymbol{\Theta}_{b}$ are as in Theorem 1. This result holds for any $b\in\mathbb{R}$ and also for sequences $b=b_{n}\to\infty$.
Remark 2.
Theorem 2 answers affirmatively the question of whether the minimax risk under classification loss can still be attained by our decision rule when the sparsity level is unknown. Since $\widehat{\tau}$ depends on the entire dataset, the decision rule for the $i$th hypothesis is no longer a function of $X_{i}$ alone, which complicates the analysis. Our proof exploits monotonicity properties of $\mathbb{E}(1-\kappa_{i}\mid X_{i},\tau)$ and $\mathbb{E}(\kappa_{i}\mid X_{i},\tau)$ as functions of $\tau$ for fixed $X_{i}$. For type I error control, we first truncate $\widehat{\tau}$ to a suitable range and apply monotonicity to reduce the problem to bounds derived for fixed $\tau$; outside this range, we use concentration arguments tailored to the structure of $\widehat{\tau}$. For type II error control, we proceed analogously using the monotonicity of $\mathbb{E}(\kappa_{i}\mid X_{i},\tau)$. Although these ideas resemble those used in Theorems 10–11 of Ghosh et al. (2016), the arguments here are necessarily different since we do not assume a two-group data-generating mechanism.
We now present the corresponding optimality result for the full Bayes rule (23).
Theorem 3.
Consider the setup of Theorem 1, and the full Bayes rule (23), where the prior on $\tau$ satisfies (C4) with $\alpha_{n}\propto\frac{(\log n)^{\delta_{3}}}{n}$ for some $0<\delta_{3}<\frac{1}{2}$. Assume that $q_{n}\propto(\log n)^{\delta_{2}}$ for some $\delta_{2}>0$. Then for the sub-class of horseshoe-type priors (i.e. $a=\frac{1}{2}$),
$\frac{1}{q_{n}}\sup_{\boldsymbol{\theta}\in\boldsymbol{\Theta}_{b}}\mathbb{E}L(\boldsymbol{\theta},\boldsymbol{\psi})=1-\Phi(b)+o(1),$ (27)
where $L(\boldsymbol{\theta},\boldsymbol{\psi})$ and $\boldsymbol{\Theta}_{b}$ are as in Theorem 1. This result holds for any $b\in\mathbb{R}$ and also for sequences $b=b_{n}\to\infty$.
3.2 Results Based on the Composite Loss $\mathrm{FDP}+\mathrm{FNP}$
We now establish sharp asymptotic minimaxity under the composite loss $\mathrm{FDP}+\mathrm{FNP}$ for the same three testing rules. As in the previous subsection, we first analyze the fixed-$\tau$ rule under known sparsity and then show that both the EB and FB rules remain minimax optimal when sparsity is unknown.
Theorem 4.
Consider the setup of Theorem 1, and the decision rule (20) based on one-group priors for the simultaneous hypothesis testing problem (2). Assume that $\tau\to 0$ as $n\to\infty$ such that $\frac{n\tau}{q_{n}}\to C\in(0,\infty)$. Also assume that the slowly varying function $L(\cdot)$ satisfies Assumption 1. Then for the subclass of horseshoe-type priors (i.e. $a=\frac{1}{2}$),
$\sup_{\boldsymbol{\theta}\in\boldsymbol{\Theta}_{b}}R(\boldsymbol{\theta},\boldsymbol{\psi})=1-\Phi(b)+o(1),$ (28)
where $\boldsymbol{\Theta}_{b}$ is as defined in Theorem 1 and $R(\boldsymbol{\theta},\boldsymbol{\psi})$ is defined in (9). This result holds for any $b\in\mathbb{R}$ and also for sequences $b=b_{n}\to\infty$.
Remark 3.
Theorem 4 establishes that the rule (20) also attains the sharp minimax risk for the composite loss $\mathrm{FDP}+\mathrm{FNP}$. In contrast, Salomond (2017) obtained only upper bounds on the maximum risk for related rules, without establishing exact minimaxity. An advantage of our result is that the minimax risk can be made arbitrarily small by taking $b=b_{n}\to\infty$. The proof proceeds by separately bounding $\mathrm{FDP}$ and $\mathrm{FNP}$: the truncation used in controlling $\mathrm{FDP}$ is inspired by Salomond (2017), but the subsequent bounds require delicate calculations, including the use of Hoeffding’s inequality. The bound for $\mathrm{FNP}$ relies on arguments developed for Theorem 1 and is independent of Salomond (2017).
Given that (20) attains the minimax risk, it is natural to ask whether the adaptive rules (21) and (23) retain minimax optimality when the sparsity level is unknown. The next results answer this question affirmatively. Proofs are provided in the Section 5.
Theorem 5.
Consider the setup of Theorem 2. Assume that $q_{n}\propto(\log n)^{\delta_{2}}$ for some $\delta_{2}>1$, and that $L(\cdot)$ satisfies Assumption 1. Then for horseshoe-type priors (i.e. $a=\frac{1}{2}$),
$\sup_{\boldsymbol{\theta}\in\boldsymbol{\Theta}_{b}}R(\boldsymbol{\theta},\boldsymbol{\psi})=1-\Phi(b)+o(1),$ (29)
where $\boldsymbol{\Theta}_{b}$ is as defined in Theorem 1 and $R(\boldsymbol{\theta},\boldsymbol{\psi})$ is defined in (9). This result holds for any $b\in\mathbb{R}$ and also for sequences $b=b_{n}\to\infty$.
Remark 4.
Theorem 5 shows that the empirical Bayes rule (21) attains the sharp minimax risk for the normal means model even when the sparsity level is unknown. While Salomond (2017) proved that an upper bound on the maximum risk converges to 0 for their proposed rules, our result establishes exact minimaxity within the horseshoe-type subclass. In particular, the minimax risk can be made arbitrarily small by taking $b=b_{n}\to\infty$.
Our final result concerns minimax optimality of the full Bayes adaptation (23) when the sparsity level is unknown.
Theorem 6.
Consider the setup of Theorem 3. Assume that $q_{n}\propto(\log n)^{\delta_{2}}$ for some $\delta_{2}>0$, and that $L(\cdot)$ satisfies Assumption 1. Then for horseshoe-type priors (i.e. $a=\frac{1}{2}$),
$\sup_{\boldsymbol{\theta}\in\boldsymbol{\Theta}_{b}}R(\boldsymbol{\theta},\boldsymbol{\psi})=1-\Phi(b)+o(1),$ (30)
where $\boldsymbol{\Theta}_{b}$ is as defined in Theorem 1 and $R(\boldsymbol{\theta},\boldsymbol{\psi})$ is defined in (9). This result holds for any $b\in\mathbb{R}$ and also for sequences $b=b_{n}\to\infty$.
4 Concluding Remarks
This paper provides the first rigorous demonstration that one-group global–local shrinkage priors, in particular the subclass of horseshoe-type priors ($a=1/2$), attain the sharp asymptotic minimax risk in the sparse Gaussian sequence model within the beta-min framework of Abraham, Castillo, and Roquain (2024). We show that such priors achieve optimality under both the classification (Hamming) loss and the composite loss $\mathrm{FDP}+\mathrm{FNP}$, thereby matching the sharp minimax benchmark achieved by the Benjamini–Hochberg procedure and spike-and-slab $\ell$-value rules at the precise minimax boundary. Moreover, minimaxity continues to hold when the global parameter is chosen adaptively, either via an empirical Bayes estimate or through a fully Bayesian prior specification, ensuring that these procedures remain effective even when the degree of sparsity is unknown.
The impact of these findings is twofold. First, they show that one-group priors, long valued for computational simplicity and robustness in estimation, can also be relied upon for decision-theoretic optimality in multiple testing. This aligns Bayesian shrinkage priors with prominent frequentist and Bayesian benchmarks, demonstrating that they are capable of achieving the same sharp minimax limits as BH and spike-and-slab approaches. Second, the results provide theoretical justification for the use of global–local priors in modern large-scale applications such as genomics, proteomics, and signal processing, where high-dimensional multiple testing is central.
At the same time, our results suggest several directions for future research. One natural question is whether the marginal maximum likelihood estimator (MMLE) of the global parameter $\tau$ (van der Pas, Szabo and van der Vaart, 2017), which is widely used in estimation, also yields minimax-optimal testing rules in this framework. Another direction is to examine whether broader classes of global–local priors, including Dirichlet–Laplace, horseshoe++, and other heavy-tailed mixtures, enjoy similar optimality properties. A further challenge is to extend the present results beyond the independent Gaussian sequence model, for instance, to dependent noise structures, generalized linear models, or heavy-tailed error distributions. Finally, it would be of interest to develop fully adaptive procedures that remain minimax optimal when both sparsity and variance parameters are unknown, thereby unifying robust estimation and optimal testing within a single framework.
In conclusion, our results establish one-group shrinkage priors as a theoretically grounded alternative to two-group formulations for high-dimensional multiple testing. By showing that these priors can attain the sharp minimax risk, our work connects Bayesian shrinkage modeling with decision-theoretic testing theory and lays the foundation for further advances at the intersection of minimax analysis, adaptivity, and high-dimensional Bayesian inference.
5 Proofs
Proof of Proposition 1.
Let us fix any $i\geq 1$. We consider the cases $a\in(\frac{1}{2},1)$ and $a\geq 1$, separately.
Case 1: Let us first consider the case when $a\in(\frac{1}{2},1)$. Using Theorem 4 of Ghosh et al. (2016), and the fact that the slowly varying function $L(\cdot)$ is bounded above by some constant M(>0)M\ (>0), we have
$E(1-\kappa_{i}\big|X_{i},\tau)\leq K_{1}e^{\frac{X_{i}^{2}}{2}}\tau^{2a}(1+o(1)),$ (31)
where the $o(1)$ term above is independent of both the index $i$ and the data point $X_{i}$, but depends on $\tau$ in such a way that $\lim_{\tau\rightarrow 0}o(1)=0$. Now, applying the definition of type II error coupled with (31), the type II error rate for the $i^{th}$ hypothesis testing problem is given by
$t_{2i}$ $=\mathbb{P}_{\theta_{i}\neq 0}\left(\mathbb{E}(1-\kappa_{i}|X_{i},\tau)\leq\frac{1}{2}\right)$
$\geq\mathbb{P}_{\theta_{i}\neq 0}\left(K_{1}\tau^{2a}\exp(\frac{X^{2}_{i}}{2})(1+o(1))\leq\frac{1}{2}\right)$
$=\mathbb{P}_{\theta_{i}\neq 0}\left(X^{2}_{i}\leq 4a\log(\frac{1}{\tau})(1+o(1))\right)$
$=\Phi\left(\sqrt{4a\log(\frac{1}{\tau})}(1+o(1)-\theta_{i}\right)-\Phi\left(-\sqrt{4a\log(\frac{1}{\tau})}(1+o(1)-\theta_{i}\right).$ (32)
Note that, to provide a lower bound to the type II error rate on the set $\boldsymbol{\Theta}_{b}$, it would be enough to consider only one candidate of $\theta_{i}$ from that set and show that the lower bound of (5) for that choice of $\theta_{i}$ tends to $1$ as $n\to\infty$. Let us fix any $b\in\mathbb{R}$ and choose $\theta_{i}=\sqrt{2\log(\frac{n}{q_{n}})}$. Then under the assumption that $\frac{n\tau}{q_{n}}\to C\in(0,\infty)$ as $n\rightarrow\infty$, we obtain
$\log\left(\frac{1}{\tau}\right)=\log\left(\frac{n}{q_{n}}\right)+\log\left(\frac{1}{C}\right)+\log(1+o(1))=\log\left(\frac{n}{q_{n}}\right)(1+o(1)),$ (33)
where the $o(1)$ term above is independent of both the index $i$ and the data point $X_{i}$, but depends on $\tau$ in such a way that $\lim_{\tau\rightarrow 0}o(1)=0$. This implies that for all sufficiently large $n$, we have
$\sup_{|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b}t_{2i}$ $\geq\Phi\left((\sqrt{4a}-\sqrt{2})\log\left(\frac{n}{q_{n}}\right)\right)(1+o(1))-\Phi\left((-\sqrt{4a}-\sqrt{2})\log\left(\frac{n}{q_{n}}\right)\right)(1+o(1))$
$=\Phi\left((\sqrt{4a}-\sqrt{2})\log\left(\frac{n}{q_{n}}\right)\right)(1+o(1))$
$\to 1,\textrm{ as }n\to\infty.$ (34)
In the above chain of inequalities, observe that for any fixed $a\in(\frac{1}{2},1)$, $\Phi\left((\sqrt{4a}-\sqrt{2})\log(\frac{n}{q_{n}})\right)(1+o(1))\to 1$ as $n\to\infty$ and $\Phi\left((-\sqrt{4a}-\sqrt{2})\log(\frac{n}{q_{n}})\right)(1+o(1))\to 0$ as $n\to\infty$.
Case 2: Next, we consider the case where $a\geq 1$. First, we derive an upper bound to $E(1-\kappa_{i}|X_{i},\tau)$ for any fixed $\tau\in(0,1)$ and $X_{i}\in\mathbb{R}$.
Observe that, the posterior distribution of $\kappa_{i}$ given $X_{i}$ and $\tau$ is,
[…]
Using the transformation $t=\frac{1}{\tau^{2}}(\frac{1}{\kappa}-1)$, $E(1-\kappa_{i}|X_{i},\tau)$ becomes
$E(1-\kappa_{i}|X_{i},\tau)=\frac{\tau^{2}\int_{0}^{\infty}(1+t\tau^{2})^{-\frac{3}{2}}t^{-a}L(t)\exp\bigg({\frac{X^{2}_{i}}{2}\cdot\frac{t\tau^{2}}{1+t\tau^{2}}}\bigg.)dt}{\int_{0}^{\infty}(1+t\tau^{2})^{-\frac{1}{2}}t^{-a-1}L(t)\exp\bigg({\frac{X^{2}_{i}}{2}\cdot\frac{t\tau^{2}}{1+t\tau^{2}}}\bigg.)dt}.$
Note that,
$\int_{0}^{\infty}(1+t\tau^{2})^{-\frac{1}{2}}t^{-a-1}L(t)\exp\bigg({\frac{X^{2}_{i}}{2}\cdot\frac{t\tau^{2}}{1+t\tau^{2}}}\bigg.)dt$ $\geq\int_{0}^{\infty}(1+t\tau^{2})^{-\frac{1}{2}}t^{-a-1}L(t)dt=K^{-1}(1+o(1)),$
where $o(1)$ depends only on $\tau$ and tends to zero as $\tau\to 0$. The equality in the last line follows due to the Dominated Convergence Theorem. Hence
$E(1-\kappa_{i}|X_{i},\tau)\leq K(A_{1}+A_{2})(1+o(1))\hskip 2.84544pt,$ (35)
where
$A_{1}$ $=\int_{0}^{1}\frac{t\tau^{2}}{1+t\tau^{2}}\cdot\frac{1}{\sqrt{1+t\tau^{2}}}t^{-a-1}L(t)\exp\bigg({\frac{X^{2}_{i}}{2}\cdot\frac{t\tau^{2}}{1+t\tau^{2}}}\bigg.)dt.$
and $A_{2}=\int_{1}^{\infty}\frac{t\tau^{2}}{1+t\tau^{2}}\cdot\frac{1}{\sqrt{1+t\tau^{2}}}t^{-a-1}L(t)\exp\bigg({\frac{X^{2}_{i}}{2}\cdot\frac{t\tau^{2}}{1+t\tau^{2}}}\bigg.)dt$. Since, for any $t\leq 1$ and $\tau\leq 1$, $\frac{t\tau^{2}}{1+t\tau^{2}}\leq\frac{1}{2}$, using the fact $\int_{0}^{\infty}t^{-a-1}L(t)dt=K^{-1}$,
[…]
For providing an upper bound on $A_{2}$, we divide our calculations depending on $a=1$ and $a>1$.
Next, note that, $1+t\tau^{2}\geq t\tau^{2}$ for any $t>0$ and $t\tau^{2}\geq\sqrt{t}$ if and only if $t\geq\frac{1}{\tau^{4}}$. As a result, for $a=1$, we have
$\int_{1}^{\frac{1}{\tau^{4}}}\frac{t\tau^{2}}{1+t\tau^{2}}\cdot t^{-2}L(t)dt\leq M\tau^{2}\int_{1}^{\frac{1}{\tau^{4}}}t^{-1}dt=M\tau^{2}\log(\frac{1}{\tau^{4}}),$ (37)
and
$\int_{\frac{1}{\tau^{4}}}^{\infty}\frac{t\tau^{2}}{1+t\tau^{2}}\cdot t^{-2}L(t)dt\leq M\tau^{2}\int_{\frac{1}{\tau^{4}}}^{\infty}t^{-\frac{3}{2}}dt=2M\tau^{4}.$ (38)
Hence, combining (37) and (38), for $a=1$, we obtain
$A_{2}$ $\leq 2M\exp\bigg(\frac{X^{2}_{i}}{2}\bigg.)\tau^{2}[\log(\frac{1}{\tau^{4}})+\tau^{2}]=8M\exp\bigg(\frac{X^{2}_{i}}{2}\bigg.)\tau^{2}\log(\frac{1}{\tau})(1+o(1)).$ (39)
On the other hand, for $a>1$,
$\int_{1}^{\infty}\frac{t\tau^{2}}{1+t\tau^{2}}\cdot t^{-a-1}L(t)dt\leq M\tau^{2}\int_{1}^{\infty}t^{-a}dt=\bigg(\frac{M}{(a-1)}\bigg)\tau^{2}.$ (40)
As a result, for $a>1$,
[…]
Therefore, combining (39) and (41), for $a\geq 1$, we have
$A_{2}$ $\leq K_{1}\exp\bigg(\frac{X^{2}_{i}}{2}\bigg.)\tau^{2}\log(\frac{1}{\tau})(1+o(1)),$ (42)
where $K_{1}$ is a constant depending on $M$ and $a$. Now, combining (35), (36) and (42), we obtain, for $a\geq 1$,
$E(1-\kappa_{i}|X_{i},\tau)$ $\leq K_{2}\exp\bigg(\frac{X^{2}_{i}}{2}\bigg.)\tau^{2}\log(\frac{1}{\tau})(1+o(1)),$ (43)
where $K_{2}$ is a constant depending on $K^{-1}$ and $K_{1}$. Next, again using the same technique used when $a\in(\frac{1}{2},1)$ with (43), a lower bound of $t_{2i}$ is obtained as,
$t_{2i}$ $=\mathbb{P}_{\theta_{i}\neq 0}(\mathbb{E}(\kappa_{i}|X_{i},\tau)>\frac{1}{2})$
$\geq\mathbb{P}_{\theta_{i}\neq 0}\left(K_{2}\exp\bigg(\frac{X^{2}_{i}}{2}\bigg.)\tau^{2}\log(\frac{1}{\tau})(1+o(1))\leq\frac{1}{2}\right)$
$=\mathbb{P}_{\theta_{i}\neq 0}\left(X^{2}_{i}\leq 4\log(\frac{1}{\tau})-4\log\log(\frac{1}{\tau})-o(1)-2\log 2\right)$
$=\mathbb{P}_{\theta_{i}\neq 0}\left(X^{2}_{i}\leq 4\log(\frac{1}{\tau})(1+o(1))\right)$
$=\Phi\left(\sqrt{4\log(\frac{1}{\tau})}(1+o(1)-\theta_{i}\right)-\Phi\left(-\sqrt{4\log(\frac{1}{\tau})}(1+o(1)-\theta_{i}\right).$ (44)
Next, again using the arguments used in the previous case, we have, for all sufficiently large $n$,
$\sup_{|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b}t_{2i}$ $\geq\Phi((\sqrt{4}-\sqrt{2})\log(\frac{n}{q_{n}}))(1+o(1))-\Phi((-\sqrt{4}-\sqrt{2})\log(\frac{n}{q_{n}}))(1+o(1))$
$=\Phi((\sqrt{4}-\sqrt{2})\log(\frac{n}{q_{n}}))(1+o(1))$
$\to 1,\textrm{ as }n\to\infty.$ (45)
Proof of Theorem 1.
The desired lower bound to the left-hand side of (25) follows from Theorem 7 of Abraham, Castillo, and Roquain (2024). Hence, we only need to obtain an upper bound for it. Using arguments similar to Theorem 7 of Abraham, Castillo, and Roquain (2024), note that
$\mathbb{E}L(\theta,\psi)=\sum_{i:\theta_{i}=0}t_{1i}+\sum_{i:\theta_{i}\neq 0}t_{2i},$ (46)
where $t_{1i}$ and $t_{2i}$ denote the type I and type II error rates corresponding to the $i^{th}$ hypothesis testing problem, respectively. Hence, it is enough to show that
$\frac{1}{q_{n}}\sum_{i:\theta_{i}=0}t_{1i}=o(1),$ (47)
and
$\frac{1}{q_{n}}\sum_{i:\theta_{i}\neq 0}\sup_{|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b}t_{2i}\leq 1-\Phi(b)+o(1).$ (48)
First, we prove (47). Using the architecture of the proof of Theorem 6 of Ghosh et al. (2016) with $a=\frac{1}{2}$, we have
$t_{1i}\leq\tilde{K}_{1}\frac{\tau L\left(\frac{1}{\tau^{2}}\right)}{\sqrt{\log\left(\frac{1}{\tau^{2}}\right)}}(1+o(1)),$ (49)
where $\tilde{K}_{1}$ is a positive constant that is independent of $\tau$ and $i$, and the $o(1)$ term is such that $\lim_{n\to\infty}o(1)=0$ and is independent of any $i$. Hence, using (48) and under the assumption Assumption 1 on the slowly varying function $L(\cdot)$, we get
$\frac{1}{q_{n}}\sum_{i:\theta_{i}=0}t_{1i}$ $\leq K_{1}M\frac{(n-q_{n})}{q_{n}}\frac{\tau}{\sqrt{\log(\frac{1}{\tau^{2}})}}(1+o(1))$
$\leq K_{1}M\frac{n\tau}{q_{n}}\frac{1}{\sqrt{\log(\frac{1}{\tau^{2}})}}(1+o(1))$
$\rightarrow 0,\textrm{ as }n\to\infty.$ (50)
In the last line of the above chain of inequalities, we used the facts $\frac{n\tau}{q_{n}}\rightarrow C\in(0,\infty)$ as $n\rightarrow\infty$, $\tau\rightarrow 0$ as $n\rightarrow\infty$, and $\log(\frac{1}{\tau^{2}})\rightarrow\infty$ as $\tau\rightarrow 0$. This completes the proof of (47).
Our next goal is to provide an upper bound to $t_{2i}$ that holds uniformly for each $i=1,\dots,n$. Towards that, we first use Lemma 3 of Ghosh and Chakrabarti (2017) that asserts that, for each fixed $x\in\mathbb{R}$ and any fixed $0<\tau<1$, the posterior shrinkage coefficient $\mathbb{E}(\kappa|x,\tau)$ can be bounded above by a non-negative, continuously decreasing, and measurable function $g(x,\tau,\eta,\delta)$ given by
$g(x,\tau,\eta,\delta)=g_{1}(x,\tau,\eta,\delta)+g_{2}(x,\tau,\eta,\delta),$
where
$g_{1}(x,\tau,\eta,\delta)=C_{1}\left[x^{2}\int_{0}^{\frac{x^{2}}{1+t_{0}}}\exp(-\frac{u}{2})du\right]^{-1}=C_{1}\left[x^{2}\left\{1-\exp(-\frac{x^{2}}{2(1+t_{0})})\right\}^{-1}\right],$
and
$g_{2}(x,\tau,\eta,\delta)=C_{2}\tau^{-1}\exp(-\frac{\eta(1-\delta)}{2}x^{2}),$
for any given $\eta,\delta\in(0,1)$. Here, $C_{1}$ and $C_{2}$ are two finite positive constants that do not depend on both $x$ and $\tau$ (but they do depend on $\eta$ and $\delta$), and $t_{0}$ was previously defined in Assumption 1. In addition, the above nonnegative function $g(x,\tau,\eta,\delta)$ satisfies the following:
For any $\rho>\frac{2}{\eta(1-\delta)}$,
$\lim_{\tau\rightarrow 0}\sup_{|x|>\sqrt{\rho\log(\frac{1}{\tau})}}g(x,\tau,\eta,\delta)=0.$
Let us fix any $\eta\in(0,1)$ and $\delta\in(0,1)$. Choose $\rho=\frac{2}{\eta(1-\delta)}[1+\tilde{g}(\tau)]$ where $\tilde{g}(\tau)>0$ and tends to zero as $\tau\to 0$ such that ${\tau}^{\tilde{g}(\tau)}\to 0$ as $\tau\to 0$. One possible choice of $\tilde{g}(\tau)=(\log(\frac{1}{\tau}))^{-\delta_{1}}$ for some $0<\delta_{1}<1$. Later, we may choose a more specific choice of $\tilde{g}(\tau)$, as required.
Next, we define two events $B_{n}$ and $C_{n}$ as $B_{n}=\{g(X_{i},\tau,\eta,\delta)>\frac{1}{2}\}$ and $C_{n}=\{|X_{i}|>\sqrt{\rho\log(\frac{1}{\tau})}\}$ with $\rho$ as defined before. Then, using the preceding observations, the type II error rate for the $i^{th}$ hypothesis testing problem is given by
$t_{2i}$ $=P_{\theta_{i}\neq 0}(E(\kappa_{i}|X_{i},\tau)>\frac{1}{2})$
$\leq P_{\theta_{i}\neq 0}(B_{n})$
$\leq P_{\theta_{i}\neq 0}(B_{n}\cap C_{n})+P_{\theta_{i}}(C^{c}_{n}).$ (51)
Observe that, for any $\theta_{i}\neq 0$ ,
$P_{\theta_{i}\neq 0}(B_{n}\cap C_{n})$ $=P_{\theta_{i}\neq 0}\left(g(X_{i},\tau,\eta,\delta)>\frac{1}{2},|X_{i}|>\sqrt{\rho\log(\frac{1}{\tau})}\right)$
$\leq P_{\theta_{i}\neq 0}\left(g_{1}(X_{i},\tau,\eta,\delta)>\frac{1}{4},|X_{i}|>\sqrt{\rho\log(\frac{1}{\tau})}\right)$
$+P_{\theta_{i}\neq 0}\left(g_{2}(X_{i},\tau,\eta,\delta)>\frac{1}{4},|X_{i}|>\sqrt{\rho\log(\frac{1}{\tau})}\right).$ (52)
Since, for given $\tau,\eta,\delta$, the function $g_{1}(x,\tau,\eta,\delta)$ steadily decreases in $x^{2}$, for any $|x|>\sqrt{\rho\log(\frac{1}{\tau})}$, we have
$g_{1}(x,\tau,\eta,\delta)\geq\frac{{C}_{1}}{\rho}\left[\log\left(\frac{1}{\tau}\right)\left\{1-\tau^{\frac{\rho}{2(1+t_{0})}}\right\}\right]^{-1}\rightarrow 0\textrm{ as }\tau\rightarrow 0.$
Since $\tau\rightarrow 0$ as $n\rightarrow\infty$, we have, $g_{1}(X_{i},\tau,\eta,\delta)\leq\frac{1}{4}$ with probability $1$ over the set $C_{n}$ for all sufficiently large $n$, and this would be true for any fixed $\eta\in(0,1)$ and $\delta\in(0,1)$, and any $\theta_{i}\neq 0$. Therefore,
$P_{\theta_{i}\neq 0}\left(g_{1}(X_{i},\tau,\eta,\delta)>\frac{1}{4}\mid|X_{i}|>\sqrt{\rho\log(\frac{1}{\tau})}\right)=0,\textrm{ for all sufficiently large }n.$ (53)
Again, we observe that, for given $\tau,\eta,\delta$, the function $g_{2}(x,\tau,\eta,\delta)$ steadily decreases in $x^{2}$. Therefore, using a similar argument, it can be shown easily that,
$P_{\theta_{i}\neq 0}\left(g_{2}(X_{i},\tau,\eta,\delta)>\frac{1}{4}\mid|X_{i}|>\sqrt{\rho\log(\frac{1}{\tau})}\right)=0,\textrm{ for all sufficiently large }n.$ (54)
Combining (5), (53) and (54), it therefore follows that
$P_{\theta_{i}\neq 0}(B_{n}\cap C_{n})=0,\textrm{ for all sufficiently large }n.$ (55)
(5) and (55) together imply that, for all sufficiently large $n$, we have
$t_{2i}$ $\leq P_{\theta_{i}\neq 0}\left(|X_{i}|\leq\sqrt{\rho\log(\frac{1}{\tau})}\right)$
$=\Phi(T_{n}-\theta_{i})-\Phi(-T_{n}-\theta_{i}),$ (56)
where $T_{n}=\sqrt{\rho\log(\frac{1}{\tau})}$. Next, we divide our calculations depending on whether $\theta_{i}$ is positive or not.
Case: Let us assume $\theta_{i}>0$, that is, $\theta_{i}>\sqrt{2\log\left(\frac{n}{q_{n}}\right)}+b$. Then, using the lower bound on $\theta_{i}$ coupled with (5), the type II error rate $t_{2i}$ can be bounded above as
$t_{2i}$ $\leq\Phi(T_{n}-\theta_{i})$
$\leq\Phi\left(\sqrt{\rho\log(\frac{1}{\tau})}-\sqrt{2\log(\frac{n}{q_{n}})}-b\right),$ (57)
where the choice of $\rho$ is same as discussed before. Observe that the decision rule does not depend on how $\eta\in(0,1),\delta\in(0,1)$ and $\rho>\frac{2}{\eta(1-\delta)}$ is chosen. Hence, taking infimum over all such $\rho$’s and subsequently over all possible choices of $(\eta,\delta)\in(0,1)\times(0,1)$, and finally using the continuity of $\Phi(\cdot)$, we have
$t_{2i}$ $\leq\Phi\left(\sqrt{2(1+g(\tau))\log(\frac{1}{\tau})}-\sqrt{2\log(\frac{n}{q_{n}})}-b\right).$ (58)
Next, under the assumption $\frac{n\tau}{q_{n}}\to C\in(0,\infty)$, we obtain
$\log(\frac{1}{\tau})=\log(\frac{n}{q_{n}})+\log(\frac{1}{C})+\log(1+o(1))$ (59)
and
$g(\tau)\log(\frac{1}{\tau})=\left(\log(\frac{n}{q_{n}})\right)^{1-\delta_{1}}(1+o(1)).$ (60)
Combining all these observations, we have
$t_{2i}$ $\leq\Phi(\sqrt{a_{n}+d_{n}}-\sqrt{a_{n}}-b),$ (61)
where $a_{n}=2\log(\frac{n}{q_{n}})$ and $d_{n}=2[\log(\frac{1}{C})+\log(1+o(1))]+2\left(\log(\frac{n}{q_{n}})\right)^{1-\delta_{1}}(1+o(1))$. Now, note that
$\Phi(\sqrt{a_{n}+d_{n}}-\sqrt{a_{n}}-b)-\Phi(-b)$
$=\frac{1}{\sqrt{2\pi}}\int_{-b}^{\sqrt{a_{n}+d_{n}}-\sqrt{a_{n}}-b}\exp(-\frac{x^{2}}{2})dx$
$\leq\sqrt{a_{n}+d_{n}}-\sqrt{a_{n}}$
$=\frac{(\sqrt{a_{n}+d_{n}}-\sqrt{a_{n}})(\sqrt{a_{n}+d_{n}}+\sqrt{a_{n}})}{\sqrt{a_{n}+d_{n}}+\sqrt{a_{n}}}$
$=\frac{d_{n}}{\sqrt{a_{n}+d_{n}}+\sqrt{a_{n}}}$
$\leq\frac{d_{n}}{2\sqrt{a_{n}}}.$ (62)
Note that for $d_{n}=(\log(\frac{n}{q_{n}}))^{1-\delta_{1}}(1+o(1))$, with $\frac{1}{2}<\delta_{1}<1$, we have $d_{n}=o(\sqrt{a_{n}})$, as $n\to\infty$. Using this observation, coupled with (5), we obtain
$t_{2i}$ $\leq\Phi(\sqrt{a_{n}+d_{n}}-\sqrt{a_{n}}-b)$
$=\Phi(-b)+\left[\Phi(\sqrt{a_{n}+d_{n}}-\sqrt{a_{n}}-b)-\Phi(-b)\right]$
$=\Phi(-b)+o(1).$ (63)
Note that, this $o(1)$ is independent of $i$.
Case 2: Let us assume that $\theta_{i}<0$, that is, $\theta_{i}<-\sqrt{2\log\left(\frac{n}{q_{n}}\right)}-b$. Then, using the lower bound on $\theta_{i}$ the upper bound of $t_{2i}$ obtained in (5) modifies to
$t_{2i}$ $\leq 1-\Phi(-T_{n}-\theta_{i})$
$\leq 1-\Phi\left(-\sqrt{\rho\log(\frac{1}{\tau})}+\sqrt{2\log(\frac{n}{q_{n}})}+b\right).$ (64)
Again using arguments similar to those used in establishing (58)-(61), an upper bound to the probability of type II error in this case is obtained as
$t_{2i}$ $\leq 1-\Phi(-\sqrt{a_{n}+d_{n}}+\sqrt{a_{n}}+b),$ (65)
where $a_{n}$ and $d_{n}$ are same as before. Next, using the same arguments used to derive (5), we obtain
$\Phi(-\sqrt{a_{n}+d_{n}}+\sqrt{a_{n}}+b)=\Phi(b)+o(1),$ (66)
where $o(1)$ is independent of $i$. Now, combining (5)-(66),
$t_{2i}$ $\leq 1-\Phi(b)+o(1).$ (67)
Finally (5) and (67) imply, for each $i$ with $|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b$,
$t_{2i}$ $\leq 1-\Phi(b)+o(1).$ (68)
This also ensures the upper bound to the left-hand side of (48) is of the order of $1-\Phi(b)+o(1)$ and completes the proof of Theorem 1. ∎
Proof of Theorem 2.
Let $t^{EB}_{1i}$ and $t^{EB}_{2i}$ denote the probabilities of the type I and type II errors for the empirical Bayes decision rule corresponding to the $i^{th}$ hypothesis, respectively. Then the overall risk associated with the empirical Bayes decision rule is given by
$\mathbb{E}L(\theta,\psi)=\sum_{i:\theta_{i}=0}t^{EB}_{1i}+\sum_{i:\theta_{i}\neq 0}t^{EB}_{2i},$ (69)
Similar to Theorem 1, here too, we need to obtain an upper bound to the left hand side of (26). Hence, it suffices to prove that
$\sum_{i:\theta_{i}=0}t^{EB}_{1i}=o(q_{n})$ (70)
and
$\frac{1}{q_{n}}\sum_{i:\theta_{i}\neq 0}\sup_{|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b}t^{EB}_{2i}\leq 1-\Phi(b)+o(1).$ (71)
For any $\alpha_{n}>0$, we have
$t^{EB}_{1i}$ $=\mathbb{P}_{\theta_{i}=0}(\mathbb{E}(1-\kappa_{i}|X_{i},\widehat{\tau})>\frac{1}{2})$
$=\mathbb{P}_{\theta_{i}=0}(\mathbb{E}(1-\kappa_{i}|X_{i},\widehat{\tau})>\frac{1}{2},\widehat{\tau}>2\alpha_{n})+\mathbb{P}_{\theta_{i}=0}(\mathbb{E}(1-\kappa_{i}|X_{i},\widehat{\tau})>\frac{1}{2},\widehat{\tau}\leq 2\alpha_{n})$
$=u_{1n}+u_{2n},\textrm{ say, }$ (72)
where $u_{1n}=\mathbb{P}_{\theta_{i}=0}(\mathbb{E}(1-\kappa_{i}|X_{i},\widehat{\tau})>\frac{1}{2},\widehat{\tau}>2\alpha_{n})$ and $u_{2n}=\mathbb{P}_{\theta_{i}=0}(\mathbb{E}(1-\kappa_{i}|X_{i},\widehat{\tau})>\frac{1}{2},\widehat{\tau}\leq 2\alpha_{n})$. Next, we observe that, for any given $x\in\mathbb{R}$, $\mathbb{E}(1-\kappa|x,\tau)$ is non-decreasing in $\tau$. Hence, using (49), we have
$u_{2n}$ $=\mathbb{P}_{\theta_{i}=0}(\mathbb{E}(1-\kappa_{i}|X_{i},\widehat{\tau})>\frac{1}{2},\widehat{\tau}\leq 2\alpha_{n})$
$=\mathbb{P}_{\theta_{i}=0}(\mathbb{E}(1-\kappa_{i}|X_{i},\widehat{\tau})>\frac{1}{2}\mid\widehat{\tau}\leq 2\alpha_{n})\mathbb{P}_{\theta_{i}=0}(\widehat{\tau}\leq 2\alpha_{n})$
$\leq\mathbb{P}_{\theta_{i}=0}(\mathbb{E}(1-\kappa_{i}|X_{i},2\alpha_{n})>\frac{1}{2})$
$\leq K_{1}M\frac{2\alpha_{n}}{\sqrt{\log(\frac{1}{4\alpha^{2}_{n}})}}(1+o(1)).$ (73)
Let us define, $\widehat{\tau}_{1}=\frac{1}{n}$ and $\widehat{\tau}_{2}=\frac{1}{c_{2}n}\sum_{i=1}^{n}\mathbf{1}(|X_{i}|>\sqrt{c_{1}\log n})$, with $c_{1}\geq 2$ and $c_{2}\geq 1$. Then, $\widehat{\tau}=\max\{\widehat{\tau}_{1},\widehat{\tau}_{2}\}$. Based on this observation, and using the definition of $u_{1n}$, we obtain
$u_{1n}$ $\leq\mathbb{P}_{\theta_{i}=0}(\widehat{\tau}>2\alpha_{n})$
$\leq\mathbb{P}_{\theta_{i}=0}(\widehat{\tau}_{1}>2\alpha_{n})+\mathbb{P}_{\theta_{i}=0}(\widehat{\tau}_{2}>2\alpha_{n}).$ (74)
Let us choose $\alpha_{n}>0$ such that $\widehat{\tau}_{1}=\dfrac{1}{n}\leq 2\alpha_{n}$ for all sufficiently large $n$, so that $\mathbb{P}_{\theta_{i}=0}(\widehat{\tau}_{1}>2\alpha_{n})=0$, provided $n$ is sufficiently large. Hence, (5) implies, for all sufficiently large $n$
$u_{1n}$ $\leq\mathbb{P}_{\theta_{i}=0}(\widehat{\tau}_{2}>2\alpha_{n})$
$\leq\mathbb{P}_{\theta_{i}=0}(\widehat{\tau}_{3}>\alpha_{n})+\mathbb{P}_{\theta_{i}=0}(\widehat{\tau}_{4}>\alpha_{n}),$ (75)
where $\widehat{\tau}_{3}=\dfrac{1}{c_{2}n}\sum\limits_{i:\theta_{i}\neq 0}\mathbf{1}(|X_{i}|>\sqrt{c_{1}\log n})$ and $\widehat{\tau}_{4}=\dfrac{1}{c_{2}n}\sum\limits_{i:\theta_{i}=0}\mathbf{1}(|X_{i}|>\sqrt{c_{1}\log n})$. Observe that,
$\widehat{\tau}_{3}$ $\leq\frac{1}{c_{2}}\frac{q_{n}}{n}.$ (76)
Let us choose $\alpha_{n}>0$ such that $\alpha_{n}\geq\frac{1}{c_{2}}\frac{q_{n}}{n}$ for all sufficiently large $n$, whence $\mathbb{P}_{\theta_{i}=0}(\widehat{\tau}_{3}>\alpha_{n})=0$.
Therefore, for all sufficiently large $n$, we obtain
$u_{1n}$ $\leq\mathbb{P}_{\theta_{i}=0}(\widehat{\tau}_{4}>\alpha_{n})$
$=\mathbb{P}_{\theta_{i}=0}\left(\frac{1}{n-q_{n}}\sum_{i:\theta_{i}=0}\mathbf{1}(|X_{i}|>\sqrt{c_{1}\log n})>c_{2}\frac{n\alpha_{n}}{n-q_{n}}\right)$
$\leq\mathbb{P}_{\theta_{i}=0}\left(\frac{1}{n-q_{n}}\sum_{i:\theta_{i}=0}\mathbf{1}(|X_{i}|>\sqrt{c_{1}\log n})>{\alpha_{n}}\right)$
$=\mathbb{P}_{\theta_{i}=0}\left(\frac{1}{n-q_{n}}\sum_{i:\theta_{i}=0}\mathbf{1}(|X_{i}|>\sqrt{c_{1}\log n})>p_{n}+\epsilon_{n}\right),$ (77)
where $p_{n}=2\mathbb{P}_{\theta_{i}=0}\left(|X_{i}|>\sqrt{c_{1}\log n}\right)$ and $\epsilon_{n}=\alpha_{n}-p_{n}$.
Using Mill’s ratio, we obtain,
$p_{n}=2\mathbb{P}_{\theta_{i}=0}\left(X_{i}>\sqrt{c_{1}\log n}\right)\propto\frac{n^{-c_{1}/2}}{\sqrt{\log n}}.$
Again, by our choice of $\alpha_{n}$, we have
$\epsilon_{n}$ $\geq$ $\dfrac{1}{c_{2}}\dfrac{q_{n}}{n}\left(1-\dfrac{\widetilde{C}}{\sqrt{\log n}}\cdot\dfrac{n^{1-c_{1}/2}}{q_{n}}\right)>0,$ (78)
for all sufficiently large $n$.
Using the preceding observations and applying Hoeffding’s Inequality, we obtain
$u_{1n}$ $\leq\mathbb{P}_{\theta_{i}=0}\left(\frac{1}{n-q_{n}}\sum_{i:\theta_{i}=0}\mathbf{1}(|X_{i}|>\sqrt{c_{1}\log n})>p_{n}+\epsilon_{n}\right),$
$\leq e^{-(n-q_{n})D\left(p_{n}+\epsilon_{n}||p_{n}\right)}$
$=e^{-(n-q_{n})D\left(\alpha_{n}||p_{n}\right)},$ (79)
where $D\left(p||q\right)=p\log\left(\dfrac{p}{q}\right)+(1-p)\log\left(\dfrac{1-p}{1-q}\right)$ denotes the Kullback-Leibler divergence between two Bernoulli distributed random variables with parameters $p$ and $q$, respectively.
Now, using the facts $\alpha_{n}\geq\frac{1}{c_{2}}\frac{q_{n}}{n}$, and $p_{n}\propto\frac{n^{-c_{1}/2}}{\sqrt{\log n}}$, we obtain
$D\left(\alpha_{n}||p_{n}\right)$ $=\alpha_{n}\log\left(\dfrac{\alpha_{n}}{p_{n}}\right)+(1-\alpha_{n})\log\left(\dfrac{1-\alpha_{n}}{1-p_{n}}\right)$
$=\alpha_{n}\log\left(\dfrac{\alpha_{n}}{p_{n}}\right)-(1-\alpha_{n})\left(\dfrac{\alpha_{n}-p_{n}}{1-p_{n}}\right)(1+o(1))$
$\gtrsim\frac{q_{n}}{n}\log(n)(1+o(1)),$ (80)
whence
$u_{1n}\leq e^{-(n-q_{n})D\left(\alpha_{n}||p_{n}\right)}\lesssim e^{-(n-q_{n})\frac{q_{n}}{n}\log(n)(1+o(1))}\lesssim e^{-q_{n}\log(n)}(1+o(1)).$ (81)
Let us choose $\alpha_{n}=c_{2}\frac{q_{n}}{n}$, so that $\alpha_{n}\to 0$ as $n\rightarrow\infty$. The aforesaid choice of $\alpha_{n}$ also satisfies the condition $\alpha_{n}\geq\frac{1}{c_{2}}\frac{q_{n}}{n}$ for all sufficiently large $n$. Hence, combining (5),(5) and (81), we have
$\frac{1}{q_{n}}\sum_{i\colon\theta_{i}=0}t^{EB}_{1i}$ $\lesssim\frac{(n-q_{n})\alpha_{n}}{q_{n}}\frac{1}{\sqrt{\log(\frac{1}{\alpha_{n}})}}(1+o(1))+\frac{n}{q_{n}}e^{-q_{n}\log(n)}(1+o(1)).$ (82)
Note that the choice of $\alpha_{n}$ satisfies $\frac{n\alpha_{n}}{q_{n}}\to C_{3}$ for some $0<C_{3}<\infty$. Hence, the first term in the right-hand side of (82) goes to zero as $n\to\infty$. On the other hand, since $q_{n}$ satisfies $q_{n}\propto{\log n}^{\delta_{2}}$ for some $\delta_{2}>1$, the second term too tends to zero as $n\to\infty$. This implies (70) holds.
Now we move towards establishing (71). Recall that, by definition, for any given $x\in\mathbb{R}$, $\mathbb{E}(\kappa|x,\tau)$ is non-increasing in $\tau$. Since $\widehat{\tau}\geq\frac{1}{n}$, hence,
$t^{EB}_{2i}$ $=\mathbb{P}_{\theta_{i}\neq 0}(\mathbb{E}(\kappa_{i}|X_{i},\widehat{\tau})>\frac{1}{2})$
$\leq\mathbb{P}_{\theta_{i}\neq 0}(\mathbb{E}(\kappa_{i}|X_{i},\frac{1}{n})>\frac{1}{2})$ (83)
Hence, using arguments similar to (5)-(58) with $\tau=\frac{1}{n}$, when $\theta_{i}>0$ for $i\in S_{\boldsymbol{\theta}}$, we have,
$t^{\text{EB}}_{2i}$ $\leq\Phi(\sqrt{2(1+(\log n)^{-\delta_{1}})\log n}-\sqrt{2\log(\frac{n}{q_{n}})}-b).$ (84)
Hence, in order to obtain the desired upper bound on $t^{\text{EB}}_{2i}$, it is sufficient to show that
$\sqrt{(1+(\log n)^{-\delta_{1}})\log n}-\sqrt{\log(\frac{n}{q_{n}})}\to 0$ as $n\to\infty$. Note that, for $q_{n}\propto(\log n)^{\delta_{2}}$,
$\sqrt{(1+(\log n)^{-\delta_{1}})\log n}-\sqrt{\log(\frac{n}{q_{n}})}$ $=\sqrt{(1+(\log n)^{-\delta_{1}})\log n}-\sqrt{\log n-\delta_{2}\log\log n}>0,$
for all $n$. Hence, the upper bound converging to zero is enough. Observe that
$\sqrt{(1+(\log n)^{-\delta_{1}})\log n}-\sqrt{\log(\frac{n}{q_{n}})}$ $=\sqrt{(1+(\log n)^{-\delta_{1}})\log n}-\sqrt{\log n-\delta_{2}\log\log n}$
$=\frac{(\log n)^{(1-\delta_{1})}+\delta_{2}\log\log n}{\sqrt{(1+(\log n)^{-\delta_{1}})\log n}+\sqrt{\log n-\delta_{2}\log\log n}}$
$\leq(\log n)^{(\frac{1}{2}-\delta_{1})}+\delta_{2}\frac{\log\log n}{\sqrt{\log n}}(1+o(1))$
$\to 0,\textrm{ as }n\to\infty,$ (85)
since $\frac{1}{2}<\delta_{1}<1$. Combining (84) and (5) implies the desired upper bound on $t^{\text{EB}}_{2i}$ is established for $\theta_{i}>0$ when $i\in S_{\boldsymbol{\theta}}$. The proof when $\theta_{i}<0$ holds using similar arguments. This completes the proof of Theorem 2. ∎
Proof of Theorem 3.
Similar to Theorem 2, here too, we need to obtain an upper bound of the left-hand side of (27). Hence, it suffices to prove that
$\sum_{i:\theta_{i}=0}t^{FB}_{1i}=o(q_{n})$ (86)
and
$\frac{1}{q_{n}}\sum_{i:\theta_{i}\neq 0}\sup_{|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b}t^{FB}_{2i}\leq 1-\Phi(b)+o(1),$ (87)
where $t^{FB}_{1i}$ and $t^{FB}_{2i}$ are the type I and type II errors for the full Bayes approach corresponding to $i^{th}$ hypothesis, respectively. Note that,
$E(1-\kappa_{i}|\mathbf{X})$ $=\int_{\frac{1}{n}}^{\alpha_{n}}E(1-\kappa_{i}|\mathbf{X},\tau)\pi(\tau|\mathbf{X})d\tau$
$=\int_{\frac{1}{n}}^{\alpha_{n}}E(1-\kappa_{i}|{X}_{i},\tau)\pi(\tau|\mathbf{X})d\tau\leq E(1-\kappa_{i}|{X}_{i},\alpha_{n})\cdot$ (88)
For proving the inequality above, we first use the fact that given any $X_{i}\in\mathbb{R}$, $E(1-\kappa_{i}|X_{i},\tau)$ is non-decreasing in $\tau$, and next we use (C4). Now, using Theorem 4 of Ghosh et al. (2016), we have for $a=\frac{1}{2}$, as $n\to\infty$
$E(1-\kappa_{i}|{X}_{i},\alpha_{n})\leq{K_{1}}e^{\frac{X^{2}_{i}}{2}}{\alpha_{n}}(1+o(1)).$ (89)
Here $o(1)$ depends only on $n$ such that $\lim_{n\to\infty}o(1)=0$ and is independent of $i$ and $K_{1}$ is a constant depending on $M$. Hence, we have the following :
$t^{\text{FB}}_{1i}$ $=P_{H_{0i}}(E(1-\kappa_{i}|\mathbf{X})>\frac{1}{2})$
$\leq P_{H_{0i}}\bigg(\frac{X^{2}_{i}}{2}>\log(\frac{1}{\alpha_{n}})-\log K_{1}-\log(1+o(1))\bigg)$
$=P_{H_{0i}}\bigg(|X_{i}|>\sqrt{2\log(\frac{1}{\alpha_{n}})}\bigg)(1+o(1)).$ (90)
Note that, as $\alpha_{n}<1$ for all $n\geq 1$, $2\log(\frac{1}{\alpha_{n}})>0$ for all $n\geq 1$. Next, using the fact that under $H_{0i},X_{i}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ with $1-\Phi(t)<\frac{\phi(t)}{t}$ for $t>0$, we get
$t^{\text{FB}}_{1i}$ $\leq 2\frac{\phi(\sqrt{2\log(\frac{1}{\alpha_{n}})})}{\sqrt{2\log(\frac{1}{\alpha_{n}})}}(1+o(1))=\sqrt{\frac{1}{\pi}}\frac{{\alpha}_{n}}{\sqrt{\log(\frac{1}{\alpha^{2}_{n}})}}(1+o(1)),$ (91)
where $o(1)$ depends only on $n$ such that $\lim_{n\to\infty}o(1)=0$. This implies, for sufficiently large $n$
$\frac{1}{q_{n}}\sum_{i:\theta_{i}=0}t^{\text{FB}}_{1i}$ $\lesssim\frac{n\alpha_{n}}{q_{n}\sqrt{\log(\frac{1}{\alpha_{n}})}}(1+o(1))$
$=\frac{(\log n)^{\delta_{3}}}{(\log n)^{\delta_{2}}\sqrt{\log n-\delta_{2}\log\log n}}(1+o(1))$
$=(\log n)^{\delta_{3}-\frac{1}{2}-\delta_{2}}(1+o(1))\to 0,\textrm{ as }n\to\infty,$ (92)
since $\delta_{3}<\frac{1}{2}$ and $\delta_{2}>0$. This implies (86) holds. Next, in order to establish (87) holds, it is enough to show that, for each $i=1,\cdots,q_{n}$, $\sup_{|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b}t^{\text{FB}}_{2i}\leq 1-\Phi(b)+o(1)$, where $o(1)$ is independent of $i$.
In order to provide an upper bound on the probability of type-II error induced by the decision rule (23), $t^{\text{FB}}_{2i}=P_{H_{1i}}(E(\kappa_{i}|\mathbf{X})>\frac{1}{2})$, we first note that,
$E(\kappa_{i}|\mathbf{X})$ $=\int_{\frac{1}{n}}^{\alpha_{n}}E(\kappa_{i}|\mathbf{X},\tau)\pi(\tau|\mathbf{X})d\tau$
$=\int_{\frac{1}{n}}^{\alpha_{n}}E(\kappa_{i}|{X}_{i},\tau)\pi(\tau|\mathbf{X})d\tau\leq E(\kappa_{i}|{X}_{i},\frac{1}{n}),$ (93)
where the inequality in the last line follows due to the fact that given any $X_{i}\in\mathbb{R}$, $E(\kappa_{i}|X_{i},\tau)$ is non-increasing in $\tau$. Hence, following the arguments of (84)-(5) with $\tau=\frac{1}{n}$, when $\theta_{i}>0$ for $i=1,\cdots,q_{n}$, we have,
$\sup_{|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b}t^{\text{FB}}_{2i}$ $\leq 1-\Phi(b)+o(1),$ (94)
where the $o(1)$ term is independent of any $i$. This completes the proof of Theorem 3. ∎
Proof of Theorem 4.
The lower bound to the left hand side of (28) follows from Theorem 1 of Abraham, Castillo, and Roquain (2024). Hence, it is enough to establish that the upper bound to the left hand side of (28) is of the order $1-\Phi(b)+o(1)$.
Using the definition of $FDR(\theta,\psi)$, we have for any $\lambda\in(0,1)$,
$FDR(\theta,\psi)$ $=\mathbb{E}_{\theta}\bigg[\frac{\sum_{i:\theta_{i}=0}\psi_{i}}{\max(\sum_{i=1}^{n}\psi_{i},1)}\bigg]$
$=\mathbb{E}_{\theta}\bigg[\frac{\sum_{i:\theta_{i}=0}\psi_{i}}{\max(\sum_{i=1}^{n}\psi_{i},1)}\mathbf{1}_{\sum_{i=1}^{n}\psi_{i}>\lambda q_{n}}\bigg]+\mathbb{E}_{\theta}\bigg[\frac{\sum_{i:\theta_{i}=0}\psi_{i}}{\max(\sum_{i=1}^{n}\psi_{i},1)}\mathbf{1}_{\sum_{i=1}^{n}\psi_{i}\leq\lambda q_{n}}\bigg]$
$=u_{3n}+u_{4n},\textrm{ say}.$ (95)
Observe that, for the set $\mathbf{1}_{\sum_{i=1}^{n}\psi_{i}>\lambda q_{n}}$, $\max(\sum_{i=1}^{n}\psi_{i},1)>\lambda q_{n}$. Hence, $u_{3n}$ can be bounded as
$u_{3n}$ $\leq\frac{\sum_{i:\theta_{i}=0}t_{1i}}{\lambda q_{n}}=\frac{(n-q_{n})t_{1}}{\lambda q_{n}}$
$\leq\frac{M}{\lambda}\frac{n\tau}{q_{n}}\frac{(\tau^{2})^{a-\frac{1}{2}}}{\sqrt{\log(\frac{1}{\tau^{2}})}}(1+o(1))=o(1),\textrm{ as }n\to\infty.$ (96)
The equality in the second line holds since by definition the type I error for $i^{th}$ hypothesis $t_{1i}$ is the same for all $i$, and is denoted as $t_{1}$. The second inequality follows due to the use of (49) and the assumption on $L(\cdot)$. Now, note that
$\{\sum_{i=1}^{n}\psi_{i}\leq\lambda q_{n}\}\subseteq\{\sum_{i:\theta_{i}\neq 0}\psi_{i}\leq\lambda q_{n}\}.$ (97)
This implies
$u_{4n}$ $\leq\mathbb{P}_{\theta}(\sum_{i:\theta_{i}\neq 0}\psi_{i}\leq\lambda q_{n}).$ (98)
Since $\psi_{i}$’s are independent Bernoulli random variables, we use Hoeffding’s inequality. Choose $0<\lambda<\Phi(b)$. As a result, for all sufficiently large $n$, the constant $t$ defined in Hoeffding’s inequality is obtained as
$t$ $=q_{n}-\sum_{i:\theta_{i}\neq 0}t_{2i}-\lambda q_{n}$
$=(1-\lambda)q_{n}-\sum_{i:\theta_{i}\neq 0}t_{2i}$
$\geq[\Phi(b)-\lambda+o(1)]q_{n}.$ (99)
Note that the choice of $\lambda$ implies $t>0$ for sufficiently large $n$. In order to obtain (5), we use a uniform upper bound to all those $t_{2i}$ for which $\theta_{i}\neq 0$ as derived in (68). Using (5), we obtain, for all $|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b$, $i\in S_{\boldsymbol{\theta}}$
$\mathbb{P}_{\theta}(\sum_{i:\theta_{i}\neq 0}\psi_{i}\leq\lambda q_{n})$ $\leq\exp(-\frac{2t^{2}}{q_{n}})$
$\leq\exp(-2[\Phi(b)-\lambda+o(1)]^{2}q_{n}).$ (100)
Next, combining (98)-(5), for sufficiently large $n$
$\sup_{\theta\in\Theta_{b}}u_{4n}=o(1),\textrm{ as }n\to\infty.$ (101)
Now, (5), (5) and (101) combine to prove that
$\sup_{\theta\in\Theta_{b}}FDR(\theta,\psi)=o(1),\textrm{ as }n\to\infty.$ (102)
On the other hand, for $FNR(\theta,\psi)$, from definition, we obtain
$FNR(\theta,\psi)$ $=\frac{\mathbb{E}_{\theta}(\sum_{i:\theta_{i}\neq 0}(1-\psi_{i}))}{q_{n}}$
$=\frac{1}{q_{n}}\sum_{i:\theta_{i}\neq 0}t_{2i}.$ (103)
Recall that, in (68), we have already derived that, for each $|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b$, $i\in S_{\boldsymbol{\theta}}$,
$t_{2i}$ $\leq 1-\Phi(b)+o(1),$ (104)
where the $o(1)$ term is independent of any $i$. This, in combination with (5), implies that, for all sufficiently large $n$,
$\sup_{\theta\in\Theta_{b}}FNR(\theta,\psi)$ $\leq 1-\Phi(b)+o(1).$ (105)
Finally, (102) and (105) together imply the upper bound to the left-hand side of (28) is of the order $1-\Phi(b)+o(1)$ and completes the proof. ∎
Proof of Theorem 5.
Here we follow the basic architecture of the proof of Theorem 4. Hence, we only need to show that the upper bound to the left-hand side of (29) is of the order $1-\Phi(b)+o(1)$.
Next, using the argument of (5), we have
$u_{3n}$ $\leq\frac{\sum_{i:\theta_{i}=0}t^{EB}_{1i}}{\lambda q_{n}}$
$\lesssim\frac{n\alpha_{n}}{q_{n}}\frac{(\alpha_{n})^{2a-1}}{\sqrt{\log(\frac{1}{\alpha_{n}})}}(1+o(1))+\frac{n}{q_{n}}e^{-\frac{q_{n}}{c_{2}}(1+o(1))}.$ (106)
Here inequality in the second line holds due to (82). Note that the choice of $\alpha_{n}$ satisfies $\frac{n\alpha_{n}}{q_{n}}\to C_{3}$ for some $0<C_{3}<\infty$. Hence, the first term to the right hand side of (5) goes to zero as $n\to\infty$. On the other hand, since $q_{n}$ satisfies $q_{n}\propto\left({\log n}\right)^{\delta_{2}}$ for some $\delta_{2}>1$, the second term too tends to zero as $n\to\infty$. This implies
$u_{3n}=o(1),\textrm{ as }n\to\infty.$ (107)
Again employing (97) and (98), we obtain for sufficiently large $n$
$t$ $=q_{n}-\sum_{i:\theta_{i}\neq 0}t^{EB}_{2i}-\lambda q_{n}$
$\geq[\Phi(b)-\lambda+o(1)]q_{n},$ (108)
where the upper bound on $t^{EB}_{2i}$ as obtained in (84) is used. This lower bound with Hoeffding’s inequality confirms, for sufficiently large $n$
$\sup_{\theta\in\Theta_{b}}u_{4n}=o(1),\textrm{ as }n\to\infty.$ (109)
Combining (107) and (109) ensures
$\sup_{\theta\in\Theta_{b}}FDR(\theta,\psi)=o(1),\textrm{ as }n\to\infty.$ (110)
For $FNR(\theta,\psi)$, from definition, we obtain
$FNR(\theta,\psi)$ $=\frac{\mathbb{E}_{\theta}(\sum_{i:\theta_{i}\neq 0}(1-\psi_{i}))}{q_{n}}$
$=\frac{1}{q_{n}}\sum_{i:\theta_{i}\neq 0}t^{EB}_{2i}.$ (111)
Recall that, in (84), we have already derived that, for each $|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b$, $i\in S_{\boldsymbol{\theta}}$,
$t^{EB}_{2i}$ $\leq 1-\Phi(b)+o(1),$
where the $o(1)$ term is independent of any $i$. This combining with (5) implies for all sufficiently large $n$
$\sup_{\theta\in\Theta_{b}}FNR(\theta,\psi)$ $\leq 1-\Phi(b)+o(1).$ (112)
Finally, (110) and (112) imply the upper bound to the left hand side of (29) is of the order $1-\Phi(b)+o(1)$ and completes the proof. ∎
Proof of Theorem 6.
Here, we also follow the basic architecture of the proof of Theorem 4. Hence, in this case, too, it is sufficient to show that the upper bound to the left-hand side of (30) is of the order $1-\Phi(b)+o(1)$.
Note that, using the argument of (5) and (5), we have
$u_{3n}$ $\leq\frac{\sum_{i:\theta_{i}=0}t^{FB}_{1i}}{\lambda q_{n}}$
$\lesssim\frac{n\alpha_{n}}{q_{n}\sqrt{\log(\frac{1}{\alpha_{n}})}}(1+o(1))$
$=o(1),\textrm{ as }n\to\infty.$ (113)
Again employing arguments of (97) and (98), we obtain for sufficiently large $n$
$t$ $=q_{n}-\sum_{i:\theta_{i}\neq 0}t^{FB}_{2i}-\lambda q_{n}$
$\geq[\Phi(b)-\lambda+o(1)]q_{n},$ (114)
where the upper bound on $t^{FB}_{2i}$ as obtained in Theorem 3 is used. This lower bound with Hoeffding’s inequality confirms, for sufficiently large $n$
$\sup_{\theta\in\Theta_{b}}u_{4n}=o(1),\textrm{ as }n\to\infty.$ (115)
Combining (5) and (115) ensures
$\sup_{\theta\in\Theta_{b}}FDR(\theta,\psi)=o(1),\textrm{ as }n\to\infty.$ (116)
For $FNR(\theta,\psi)$, from definition, we obtain
$FNR(\theta,\psi)$ $=\frac{\mathbb{E}_{\theta}(\sum_{i:\theta_{i}\neq 0}(1-\psi_{i}))}{q_{n}}$
$=\frac{1}{q_{n}}\sum_{i:\theta_{i}\neq 0}t^{FB}_{2i}.$ (117)
Recall that, in Theorem 3, we have already derived that, for each $|\theta_{i}|\geq\sqrt{2\log(\frac{n}{q_{n}})}+b$, $i\in S_{\boldsymbol{\theta}}$,
$t^{FB}_{2i}$ $\leq 1-\Phi(b)+o(1),$ (118)
where the $o(1)$ term is independent of any $i$. This combining with (5) implies for all sufficiently large $n$
$\sup_{\theta\in\Theta_{b}}FNR(\theta,\psi)$ $\leq 1-\Phi(b)+o(1).$ (119)
Finally, (116) and (119) imply the upper bound to the left hand side of (30) is of the order $1-\Phi(b)+o(1)$ and completes the proof. ∎